### v1

# On dense subalgebras of the singular ideal in groupoid C*-algebras

Abstract.
We prove that ideals in amenable second-countable non-Hausdorff étale groupoid $C^{*}$-algebras are determined by their isotropy fibres. As an application, we characterise when the singular functions in Connes’ algebra are dense in the singular ideal in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras.
Key words and phrases:
Étale groupoid, non-Hausdorff, $C^{*}$-algebra, singular ideal, isotropy fibres
2020 Mathematics Subject Classification:
22A22, 46L05 (Primary); 37A55, 22D25 (Secondary)
Introduction
In recent years non-Hausdorff groupoids and their $C^{*}$-algebras have gained increased attention. Many groupoids arising from dynamics and geometry are non-Hausdorff, and it is therefore important to develop a robust theory. Indeed, there are natural examples of groupoids of germs, groupoids arising from self-similar groups, and groupoids arising from foliations which are non-Hausdorff ([10], [21], [22], [23]).
In contrast with the Hausdorff case, the functions in the reduced $C^{*}$-algebra $C^{*}_{r}(G)$ of a non-Hausdorff groupoid $G$ are not necessarily continuous. Indeed, there can exist (non-zero) functions in $C^{*}_{r}(G)$ whose set of non-zero values has empty interior in $G$. The set of all such functions is known as the singular ideal $J$ in $C^{*}_{r}(G)$. Historically, the singular ideal has been an obstruction to understanding simplicity for reduced $C^{*}$-algebras of non-Hausdorff groupoids ([9]). Characterisations of simplicity and the ideal intersection property have been obtained for the quotient by this ideal ([9], [18]), which is known as the essential groupoid $C^{*}$-algebra ([19], [13]). Therefore, the only obstacle to understanding these properties for the reduced $C^{*}$-algebra is the singular ideal.
Much work has been done to find conditions that ensure the singular ideal $J$ vanishes (see [9], [19], [24], [15]). Recently in [16] vanishing of the singular ideal was characterised in terms of a property of certain collections of subgroups in isotropy groups of $G$. In [4] it is characterised when the singular ideal $J$ has trivial intersection with the underlying groupoid -algebra, $\mathscr{C}_{c}(G)$ known as Connes’ algebra (first defined in [10] for groupoids arising from foliations).
When the singular ideal is non-zero, less is known about its structure. For example, it is not known in general whether $J\neq\{0\}$ implies $J\cap\mathscr{C}_{c}(G)\neq\{0\}$. This was shown for groupoids with finite “non-Hausdorfness” in [4] and in [16] the problem was shown to be equivalent to whether group $C^{*}$-algebras satisfy the Intersection Property. This reduction of the problem led to its solution for groupoids with isotropy groups that are direct limits of virtually torsion free solvable groups [16].
A related question asks for which groupoids is $J\cap\mathscr{C}_{c}(G)$ dense in $J$ ([4, Question 4.11(III)]). Progress could be made towards understanding the structure of $J$ if it has a dense subalgebra of functions in $\mathscr{C}_{c}(G)$ (since functions in $\mathscr{C}_{c}(G)$ are easier to understand). In recent work ([20]), it is shown that there exist non-amenable groupoids for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$. The following question remains open however.
Question.
Is $J\cap\mathscr{C}_{c}(G)$ dense in $J$ for any amenable étale groupoid $G$?
The goal of our paper is to study the question highlighted above. Let us outline the main achievements of the paper. All results listed here are for amenable second-countable étale groupoids.
  • We show that ideals in the groupoid $C^{*}$-algebra are determined by their isotropy fibres. This generalises [8, Theorem 2.10] to the setting of non-Hausdorff groupoids.
  • Following the approach in [16], we characterise when $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras.
  • We describe some large classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$. In particular, we show density holds for groupoids with abelian isotropy groups, and groupoids arising from contracting self-similar groups.
We now state our main results and describe them in more detail. Our first result concerns ideals in groupoid $C^{*}$-algebras. In [8] it is shown that ideals $I\trianglelefteq C^{*}(G)$ in a groupoid $C^{*}$-algebra are mapped to ideals in the isotropy group $C^{*}$-algebras by the restriction maps $C^{*}(G)\to C^{*}(G^{x}_{x})$. These images are known as the isotropy fibres of the ideal, and are denoted $(I_{x})_{x\in G^{(0)}}$. It is shown in [8, Theorem 2.10] that for amenable second-countable Hausdorff étale groupoids, ideals in the groupoid $C^{*}$-algebra are determined by their isotropy fibres. We generalise this result to non-Hausdorff groupoids.
Theorem A (See Theorem 2.5).
Let $G$ be an amenable and second-countable (non-Hausdorff) étale groupoid and suppose $I$ and $K$ are ideals in $C^{*}(G)$. Then, $I=K$ if and only if $I_{x}=K_{x}$ for all $x\in G^{(0)}$.
We apply this result to the singular ideal $J$ in the reduced $C^{*}$-algebra of a non-Hausdorff groupoid. The following Theorem reduces the density question for $J\cap\mathscr{C}_{c}(G)$ in $J$ to one about density of ideals in isotropy group $C^{*}$-algebras.
Theorem B (See Proposition 2.6).
Let $G$ be an amenable and second-countable étale groupoid. Then, $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ if and only if $J_{x}\cap\mathbb{C}[G^{x}_{x}]$ is dense in $J_{x}$ for all $x\in G^{(0)}$.
The isotropy fibres $J_{x}$ are calculated explicitly in [16, Theorem 5.5]. Moreover, when $G$ is covered by countably many bisections and $G^{x}_{x}$ is amenable, it is shown $J_{x}$ equals the ideal $J_{G^{x}_{x},\mathcal{X}(x)}$ defined as the intersection of the kernels of quasi-regular representations associated to subgroups of $G^{x}_{x}$ in a certain a collection $\mathcal{X}(x)$. We refer the reader to Subsection 2.1 for the definition of $\mathcal{X}(x)$ and $J_{G^{x}_{x},\mathcal{X}(x)}$. The following is then a rephrasing of Theorem B.
Theorem C (See Theorem 2.9).
Let $G$ be an amenable and second-countable étale groupoid. Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ if and only if $J_{G^{x}_{x},\mathcal{X}(x)}\cap\mathbb{C}[G^{x}_{x}]$ is dense in $J_{G^{x}_{x},\mathcal{X}(x)}$ for all $x\in G^{(0)}$. In particular, if all the isotropy groups $G^{x}_{x}$ satisfy the Density Property, then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
In the above theorem, the Density Property is a property of discrete amenable groups we introduce in Definition 2.8. Finite groups satisfy the Density Property automatically, and in Proposition 3.11 we show that all abelian groups have the Density Property. Currently, we do not know of any discrete amenable group that fails the Density Property. We remark that it is a corollary of Theorem C and a non-Hausdorff groupoid construction in [16, Section 6] that if there is such a group, then there is an amenable groupoid for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$ (see Corollary 2.12).
By the definition of the ideals $J_{G^{x}_{x},\mathcal{X}(x)}$, Theorem C asserts that density of $J\cap\mathscr{C}_{c}(G)$ in $J$ can be understood entirely in terms of quasi-regular representations of isotropy groups. In light of this, we are able to describe some explicit classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Theorem D (See Corollary 3.12).
Let $G$ be an amenable and second-countable étale groupoid. Assume for every $x\in G^{(0)}$ one of the following holds.
  1. (I)
    The isotropy group $G_{x}^{x}$ is abelian.
  2. (II)
    The subgroups $X\in\mathcal{X}(x)$ are finite.
Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Theorem D covers many important classes of groupoids. In particular, any groupoid of germs associated with the action of a countable abelian group will satisfy (I). All groupoids with finite isotropy groups will satisfy (II). More generally, condition (II) is satisfied whenever the unit space of $G$ has finite source and range fibres.
In the setting of non-Hausdorff ample groupoids, one can intersect $J$ with the complex Steinberg algebra $\mathbb{C}G$ (see [26]) to obtain what is known as the algebraic singular ideal $J_{\mathbb{C}}$. A characterisation for the vanishing of $J_{\mathbb{C}}$ is given in [4], and together with [9] and [27] this leads to a complete characterisation of simplicity for Steinberg algebras (see also [28] and [15]). In this setting, one can ask the following density question: for which ample groupoids is $J_{\mathbb{C}}$ dense in $J$? Indeed, this is presented as an open problem in [15]. In order to answer this question, we show it suffices to study density of $J\cap\mathscr{C}_{c}(G)$ in $J$.
Theorem E (See Theorem 4.4).
Let $G$ be an ample groupoid. Then $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
An important class of ample groupoids are those arising from self-similar groups (see [22]). These groupoids can be non-Hausdorff, and can also have non-vanishing singular ideal. Using our results, we answer the density question for groupoids arising from contracting self-similar groups.
Corollary F (See Corollary 4.5).
Let $G$ be the groupoid arising from the action of a contracting self-similar group on a finite alphabet. Then $J_{\mathbb{C}}$ is dense in $J$.
In [20], examples of groupoids arising from self-similar actions on infinite alphabets are described for which $J_{\mathbb{C}}$ is not dense in $J$.
Acknowledgments
The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Topological groupoids and their $C^{*}$-algebras, where work on this paper was undertaken. This work was supported by EPSRC grant EP/V521929/1. Hume was supported by NSERC Discovery Grant RGPIN-2021-03834. Gonzales would like to thank Xin Li for helpful discussions on ideas related to this paper.
1. Preliminaries
1.1. Étale groupoids
[…]
A topological groupoid $G$ is a groupoid equipped with a topology such that the product and inverse maps are continuous. By an étale groupoid $G$, we mean a locally compact topological groupoid for which the unit space $G^{(0)}$ is Hausdorff, and the range $r\colon G\to G^{(0)}$ and source $s\colon G\to G^{(0)}$ maps are local homeomorphisms. This paper will only consider étale groupoids, and primarily concerns non-Hausdorff groupoids (all results also hold for Hausdorff groupoids, but often for trivial reasons i.e. because the singular ideal vanishes for Hausdorff groupoids).
An open subset $U\subseteq G$ will be called an open bisection if the restrictions $r|_{U}\colon U\to r(U)$ and $s|_{U}\colon U\to s(U)$ are homeomorphisms. Note that any étale groupoid has a basis of open bisections. We define
$\mathscr{C}_{c}(G)\coloneq\text{span}\{f\colon G\to\mathbb{C}\colon f|_{U}\in C_{c}(U)\text{ and }f|_{G\setminus U}=0\text{ for some open bisection }U\subseteq G\}.$
If the groupoid $G$ is Hausdorff, then $\mathscr{C}_{c}(G)=C_{c}(G)$, the set of all compactly supported continuous functions on $G$. However if $G$ is not Hausdorff, then functions in $\mathscr{C}_{c}(G)$ need not be continuous. We equip $\mathscr{C}_{c}(G)$ with the structure of a -algebra as follows. For $f,f_{1},f_{2}\in\mathscr{C}_{c}(G)$, and $g\in G$ define
$f^{*}(g)\coloneq\overline{f(g^{-1})}\quad\text{and}\quad f_{1}*f_{2}(g)\coloneq\sum_{h\in G_{s(g)}}f_{1}(gh^{-1})f_{2}(h).$
We now describe the reduced and full groupoid $C^{*}$-algebras. Given $x\in G^{(0)}$, let $\lambda_{x}\colon\mathscr{C}_{c}(G)\to B(\ell^{2}(G_{x}))$ be the representation given by $\lambda_{x}(f)(\xi)=f*\xi$ for $f\in\mathscr{C}_{c}(G)$ and $\xi\in\ell^{2}(G_{x})$, where $(f*\xi)(g)=\sum_{h\in G_{x}}f(gh^{-1})\xi(h)$ for $g\in G_{x}$. Then the reduced groupoid $C^{*}$-algebra $C^{*}_{r}(G)$ of $G$ is defined to be the completion of $\mathscr{C}_{c}(G)$ with respect to the norm $\|f\|_{C^{*}_{r}(G)}=\sup_{x\in G^{(0)}}\|\lambda_{x}(f)\|$. The full groupoid $C^{*}$-algebra $C^{*}(G)$ of $G$ is defined to be the completion of $\mathscr{C}_{c}(G)$ with respect to the norm $\|f\|_{C^{*}(G)}=\sup_{\pi}\|\pi(f)\|$, where the supremum is taken over all -algebra representations $\pi\colon\mathscr{C}_{c}(G)\to B(H)$, for some Hilbert space $H$.
Elements of the reduced groupoid $C^{*}$-alegbra $C^{*}_{r}(G)$ can be viewed as bounded Borel functions on $G$ in a canonical way (see [6, Lemma 3.27], [25, Proposition II.4.2]). From now on, we will always treat elements of $C^{*}_{r}(G)$ as functions on $G$ in this way. Given $a\in C^{*}_{r}(G)$, we define
${\rm supp}^{\circ}(a)\coloneq\{g\in G\colon a(g)\neq 0\}.$
Definition 1.1 ([9], [19], [13]).
The singular ideal $J$ in $C^{*}_{r}(G)$ is defined as
$J\coloneq\{a\in C^{*}_{r}(G)\colon{\rm supp}^{\circ}(a)\text{ has empty interior}\}.$
The singular ideal $J$ is a (closed) ideal in $C^{*}_{r}(G)$. For $a\in C^{*}_{r}(G)$, ${\rm supp}^{\circ}(a)$ has empty interior in $G$ if and only if $s\big({\rm supp}^{\circ}(a)\big)$ has empty interior in $G^{(0)}$ (see [1, Proposition 4.6], [4, Lemma 4.1(iii)]).
1.2. Hausdorff cover
We now introduce the Hausdorff cover groupoid. See [4] for an extensive study of the Hausdorff cover for non-Hausdorff étale groupoids. Let $G$ be a (non-Hausdorff) étale groupoid, and let $\mathfrak{C}(G)$ be its space of closed subsets. Singleton sets in $G$ are closed, hence there is a canonical inclusion $\iota\colon G\to\mathfrak{C}(G)$ given by $\iota(g)\coloneq\{g\}$. When equipped with the Fell topology, $\mathfrak{C}(G)$ has the structure of a compact Hausdorff space (see [14]).
Definition 1.2 ([29]).
Let $G$ be an étale groupoid. The Hausdorff cover $\tilde{G}$ is defined to be the closure of $\iota(G)$ in $\mathfrak{C}(G)\setminus\{\emptyset\}$ with respect to the Fell topology.
Elements of $\tilde{G}$ can be described explicitly as follows. A non-empty subset $\bm{g}\subseteq G$ lies in $\tilde{G}$ if and only if there exist a net $(g_{i})$ in $G$ such that $\bm{g}$ is the set of limit points of $(g_{i})$, and every limit point of a subnet of $(g_{i})$ (cluster point) is a limit point of $(g_{i})$. Moreover, given a net $(g_{i})$ in $G$, the image net $\iota(g_{i})$ converges to $\bm{g}\in\tilde{G}$ in the topology of $\tilde{G}$ if and only if the conditions above are satisfied.
Define $\tilde{G}^{(0)}\coloneq\{\bm{g}\in\tilde{G}\colon\bm{g}\cap G^{(0)}\neq\emptyset\}$. This set will be the unit space of $\tilde{G}$ with the structure described below. The canonical inclusion map $\iota\colon G\hookrightarrow\tilde{G}$ restricts to an inclusion $G^{(0)}\hookrightarrow\tilde{G}^{(0)}$, and the image $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$. For each $\bm{x}\in\tilde{G}^{(0)}$ there is a unique element $\pi(\bm{x})\in\bm{x}\cap G^{(0)}$, and the map $\pi\colon\tilde{G}^{(0)}\to G^{(0)}$ is a continuous surjection. Since $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$, continuity of the groupoid operations on $G$ implies that any element $\bm{x}\in\tilde{G}^{(0)}$ is a subgroup of the isotropy group $G_{\pi(\bm{x})}^{\pi(\bm{x})}$. We will often write $\bm{x}=X$ to distinguish elements $x\in G^{(0)}$ from $\bm{x}\in\tilde{G}^{(0)}$.
We now describe the groupoid operations on $\tilde{G}$. Note that $r(g_{1})=r(g_{2})$ and $s(g_{1})=s(g_{2})$ whenever $g_{1},g_{2}\in\bm{g}$ and $\bm{g}\in\tilde{G}$. The fact that elements of $\tilde{G}^{(0)}$ are subgroups implies, for every $\bm{g}\in\tilde{G}$, there are unique elements $\bm{x},\bm{y}\in\tilde{G}^{(0)}$ such that $\bm{g}=g\cdot\bm{x}=\bm{y}\cdot g$ for any $g\in\bm{g}$. Let $s(\bm{g})=\bm{x}$ and $r(\bm{g})=\bm{y}$. Define the inverse $\bm{g}^{-1}\coloneq\{g^{-1}\colon g\in\bm{g}\}$ and the product $\bm{g}\bm{h}\coloneq\{gh\colon g\in\bm{g},h\in\bm{h}\}$ whenever $s(\bm{g})=r(\bm{h})$. Note also that
$\bm{g}\bm{h}=\{gh\colon h\in\bm{h}\}\text{ for any }g\in\bm{g}.$ (1.1)
With this structure, $\tilde{G}$ becomes a Hausdorff étale groupoid.
[…]
$\mathfrak{i}(f)(\bm{g})=\sum_{g\in\bm{g}}f(g)$ (1.2)
for all $\bm{g}\in\tilde{G}$ ([4, Lemma 3.2]). The embedding $Cc(G~)\mathfrak{i}\colon\mathscr$$\hookrightarrow C_{c}(\tilde{G})$ is a -homomorphism, and induces canonical embeddings $Cr(G~)\mathfrak{i}_{r}\coloneq C^{*}_{r}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}_{r}(\tilde{G})$ and $C(G~)\mathfrak{i}\coloneq C^{*}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}(\tilde{G})$ ([4, Lemma 3.8 & Corollary 6.8]).
Now assume that the groupoid $G$ is covered by countably many open bisections. We say a unit $x\in G^{(0)}$ is Hausdorff if $\pi^{-1}(x)=\{\iota(x)\}$, and we let $C\subseteq G^{(0)}$ denote the set of Hausdorff units. By [19, Lemma 7.15], $C$ is dense in $G^{(0)}$ (the Hausdorff points complement the so-called dangerous points of [19]). Define $\tilde{G}^{(0)}_{\rm{ess}}\coloneq\overline{\iota(C)}$. This is an invariant subset of the unit space $\tilde{G}^{(0)}$. We define the essential Hausdorff cover $\tilde{G}_{\rm{ess}}\coloneq\tilde{G}|_{\tilde{G}^{(0)}_{\rm{ess}}}=\{\bm{g}\in\tilde{G}\colon s(\bm{g})\in\tilde{G}^{(0)}_{\rm{ess}}\}$ ([4, Definition 4.13]).
1.3. Isotropy fibres of ideals
Our attention now turns to restriction maps. Let $G$ be an étale groupoid. Given a unit $x\in G^{(0)}$, consider the restriction map $\eta_{x}\colon\mathscr{C}_{c}(G)\to\mathbb{C}[G^{x}_{x}]$, where $\mathbb{C}[G^{x}_{x}]$ denotes the group algebra of the isotropy group $G^{x}_{x}$ (thought of as finitely supported functions $G^{x}_{x}\to\mathbb{C}$). The map $\eta_{x}$ extends to a completely positive contraction $C^{*}(G)\to C^{*}(G^{x}_{x})$, which we will also denote by $\eta_{x}$. This can be seen using induced representations of the isotropy group (see [7, Lemma 1.2] for example). In fact, any $C^{*}$-norm on $\mathscr{C}_{c}(G)$ dominating the reduced norm will define a $C^{*}$-norm on the group algebra $\mathbb{C}[G^{x}_{x}]$ dominating the reduced norm, and the restriction map $\mathscr{C}_{c}(G)\to\mathbb{C}[G^{x}_{x}]$ will always extend to a completely positive contraction between the associated $C^{*}$-algebras. See [8, Section 1.3] for a discussion in this direction. For a (closed) ideal $I\trianglelefteq C^{*}(G)$, the image $\eta_{x}(I)$ is also a (closed) ideal in $C^{*}(G^{x}_{x})$ ([8, Lemma 2.1]). We write $I_{x}\coloneq\eta_{x}(I)$ and, following [8], call $I_{x}$ the isotropy fibre of $I$ at $x$.
For $g\in G^{x}_{x}$, denote by $\delta_{g}\in\mathbb{C}[G^{x}_{x}]$ the function equal to one at $g$ and zero everywhere else. Each $g\in G$ induces an isomorphism of group algebras $\text{Ad}_{g}\colon\mathbb{C}\big[G^{s(g)}_{s(g)}\big]\to\mathbb{C}\big[G^{r(g)}_{r(g)}\big]$, $\delta_{h}\mapsto\delta_{ghg^{-1}}$ and this maps extends to an isomorphism of $C^{*}$-algebras $\Psi_{g}\colon C^{*}\big(G^{s(g)}_{s(g)}\big)\to C^{*}\big(G^{r(g)}_{r(g)}\big)$. A family $(K_{x})_{x\in G^{(0)}}$ of ideals satisfying $K_{x}\trianglelefteq C^{*}(G_{x}^{x})$ for each $x\in G^{(0)}$ is said to be invariant if $\Psi_{g}(K_{s(g)})=K_{r(g)}$ for all $g\in G$. Given an ideal $I\trianglelefteq C^{*}(G)$, its isotropy fibres form an invariant family of ideals. Going in the other direction, if $\mathcal{K}=(K_{x})_{x\in G^{(0)}}$ is an invariant family of ideals, we define
[…]
By [8, Corollary 2.9], $I(\mathcal{K})$ is an ideal in $C^{*}(G)$. We isolate the following observation from [8].
Lemma 1.3.
Let $\mathcal{K}=(K_{x})_{x\in G^{(0)}}$ be an invariant family of ideals, and let $A\subseteq C^{*}(G)$ be a sub-$C^{*}$-algebra. Then $A\subseteq I(\mathcal{K})$ if and only if $\eta_{x}(A)\subseteq K_{x}$ for all $x\in G^{(0)}$.
Proof.
Assume that $A$ satisfies $A\subseteq I(\mathcal{K})$. Fix $x\in G^{(0)}$. We have that $\eta_{x}(a)\in K_{x}$ for all positive elements $a\in A^{+}$. Since $A$ is spanned by its positive elements, it follows that $\eta_{x}(a)\in K_{x}$ for all $a\in A$. Conversely, assume that $\eta_{x}(A)\subseteq K_{x}$ for all $x\in G^{(0)}$. For any $a\in A$ we have $\eta_{x}(a^{*}a)\in K_{x}$ for all $x\in G^{(0)}$, implying that $a\in I(\mathcal{K})$ as desired. ∎
2. Ideals are determined by their isotropy fibres
[…]
Remark 2.1.
[…]
Our first observation is that the assignment $I\mapsto\tilde{I}$ is injective.
Proposition 2.2.
For any ideal $I$ in $C^{*}(G)$, we have $\mathfrak{i}^{-1}(\tilde{I})=I$.
Proof.
[…]
Let $X\in\tilde{G}^{(0)}$ and write $x\coloneq\pi(X)\in G^{(0)}$. By definition, $X$ is equal to a subgroup of the isotropy group $G^{x}_{x}$. It is easy to see that the isotropy group $\tilde{G}^{X}_{X}\subseteq\tilde{G}$ is equal to the quotient $N_{X}/X$, where $N_{X}=\{g\in G^{x}_{x}:gXg^{-1}=X\}$ is the normaliser of $X$ in $G^{x}_{x}$. Denote by $E_{N_{X}}\colon C^{*}(G^{x}_{x})\to C^{*}(N_{X})$ the canonical conditional expectation and $Q_{X}\colon C^{*}(N_{X})\to C^{*}(N_{X}/X)$ the -homomorphism induced by the quotient map $N_{X}\to N_{X}/X$.
Lemma 2.3.
Let $X\in\tilde{G}^{(0)}$ and write $x\coloneq\pi(X)\in G^{(0)}$. The following diagram commutes.
$C^{*}(G)$${C^{*}(\tilde{G})}$$C^{*}(G^{x}_{x})$$C^{*}(\tilde{G}^{X}_{X})$$\scriptstyle{\mathfrak{i}}$$\scriptstyle{\eta_{x}}$$\scriptstyle{\eta_{X}}$$\scriptstyle{Q_{X}\circ E_{N_{X}}}$
Proof.
Equality is verified readily on the dense *-subalgebra $\mathscr{C}_{c}(G)$, and the lemma follows immediately. ∎
We can now describe the isotropy fibres of the ideals $\tilde{I}$ $C^{*}(\tilde{G}) in$ terms of $I$.
Proposition 2.4.
Let $I$ be an ideal in $C^{*}(G)$. Let $X\in\tilde{G}^{(0)}$ and write $x\coloneq\pi(X)\in G^{(0)}$. Then,
$\tilde{I}_{X}=\overline{Q_{X}\circ E_{N_{X}}(I_{x})}\trianglelefteq C^{*}(\tilde{G}^{X}_{X}).$
Proof.
For each $X\in\tilde{G}^{(0)}$ define the ideal $K_{X}\coloneq\overline{Q_{X}\circ E_{N_{X}}(I_{x})}\trianglelefteq C^{*}(\tilde{G}^{X}_{X})$. Let $\bm{g}\in\tilde{G}$, and observe that $\Psi_{\bm{g}}\circ Q_{s(\bm{g})}\circ E_{N_{s(\bm{g})}}=Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}\circ\Psi_{g_{0}}$ for any $g_{0}\in\bm{g}$, where $\Psi_{\bm{g}}$ and $\Psi_{g_{0}}$ are as defined in subsection 1.3 (this equality can be seen using (1.1)). Since $\Psi_{\bm{g}}$ is an isomorphism, and since $(I_{x})_{x\in G^{(0)}}$ is an invariant family of ideals, it follows that
$\Psi_{\bm{g}}(K_{s(\bm{g})})$ $=\overline{\Psi_{\bm{g}}\big(Q_{s(\bm{g})}\circ E_{N_{s(\bm{g})}}(I_{s(g_{0})})\big)}=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}\big(\Psi_{g_{0}}(I_{s(g_{0})})\big)}$
$=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}(I_{r(g_{0})})}=K_{r(\bm{g})}.$
Hence, $\mathcal{K}\coloneq(K_{X})_{X\in\tilde{G}^{(0)}}$ is an invariant family of ideals.
By Lemma 2.3, we have $K_{X}=\overline{\eta_{X}\big(\mathfrak{i}(I)\big)}$, and therefore $K_{X}\subseteq\tilde{I}_{X}$ for all $X\in\tilde{G}^{(0)}$. By Lemma 1.3, this equality also implies that $\mathfrak{i}(I)\subseteq I(\mathcal{K})$. Hence $\tilde{I}\subseteq I(\mathcal{K})$, and another application of Lemma 1.3 gives $\tilde{I}_{X}\subseteq K_{X}$ for all $X\in\tilde{G}^{(0)}$, as desired. ∎
We now prove that ideals in $C^{*}(G)$ are determined by their isotropy fibres whenever $G$ is an amenable second-countable étale groupoid. This generalises [8, Theorem 2.10] to the setting of non-Hausdorff groupoids. Our proof utilises this theorem, which in turn uses known cases of the Effros-Hahn conjecture [17]. Therefore it is necessary for us to only consider amenable second-countable groupoids. The theorem does not hold in the non-amenable setting - counterexamples are provided in [8, Examples 2.11 & 2.12].
Theorem 2.5.
Let $G$ be an amenable second-countable (non-Hausdorff) étale groupoid and suppose $I$ and $K$ are ideals in $C^{*}(G)$. Then, $I=K$ if and only if $I_{x}=K_{x}$ for all $x\in G^{(0)}$.
Proof.
The forwards implication is trivial. Assume $I_{x}=K_{x}$ for all $x\in G^{(0)}$. Proposition 2.4 implies that $\tilde{I}_{X}=\tilde{K}_{X}$ for all $X\in\tilde{G}^{(0)}$. Since $G$ is amenable and second-countable, so is $\tilde{G}$ by [4, Theorem 6.5]. Moreover, the groupoid $\tilde{G}$ is Hausdorff, and therefore [8, Theorem 2.10] ensures that $\tilde{I}=\tilde{K}$. Finally, Proposition 2.2 gives $I=\mathfrak{i}^{-1}(\tilde{I})=\mathfrak{i}^{-1}(\tilde{K})=K$. ∎
2.1. Characterisation of density
In this subsection, we assume that $G$ is an amenable second-countable (not necessarily Hausdorff) étale groupoid, and therefore $C^{*}(G)=C^{*}_{r}(G)$ by [4, Corollary 6.10]. We denote by $J$ the singular ideal in $C^{*}_{r}(G)$. Theorem 2.5 yields the following characterisation for the density of $J\cap\mathscr{C}_{c}(G)$ in $J$.
Proposition 2.6.
Let $G$ be an amenable and second-countable étale groupoid. Then, $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ if and only if $J_{x}\cap\mathbb{C}[G^{x}_{x}]$ is dense in $J_{x}$ for all $x\in G^{(0)}$.
Proof.
Let $I=\overline{J\cap\mathscr{C}_{c}(G)}$. Since $\mathscr{C}_{c}(G)$ is dense in $C^{*}_{r}(G)$, $I$ is an ideal and $I\subseteq J$. By [16, Theorem 5.5 & Proposition 5.19], we have $\eta_{x}(J\cap\mathscr{C}_{c}(G))=J_{x}\cap\mathbb{C}[G^{x}_{x}]$. Continuity of $\eta_{x}$ then implies $J_{x}\cap\mathbb{C}[G^{x}_{x}]\subseteq\eta_{x}(I)\subseteq\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}$. Since $\eta_{x}(I)=I_{x}$ is closed, it follows that $I_{x}=\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}$. Therefore, Theorem 2.5 shows that $I=J$ if and only if $\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}=J_{x}$ for all $x\in G^{(0)}$, as desired. ∎
Remark 2.7.
The “only if” direction of Proposition 2.6 holds for all étale groupoids $G$ because $\eta_{x}(J\cap\mathscr{C}_{c}(G))\subseteq J_{x}\cap\mathbb{C}[G^{x}_{x}]$ holds trivially ($J_{x}$ here means the isotropy fibre from the reduced groupoid $C^{*}$-algebra). It can be shown using the calculation of isotropy fibres of the singular ideal in [16, Theorem 5.5] that the first example in [20] of a non-Hausdorff groupoid $G$ with $J\cap\mathscr{C}_{c}(G)$ not dense in $J$ has an isotropy fibre $J_{\epsilon}\neq\{0\}$ with $J_{\epsilon}\cap\mathbb{C}[\Gamma]=\{0\}$. Therefore, this remark may be applied to this example to provide an alternate proof of non-density.
Following [16], we now explain how the density question can be reduced to one about explicit ideals in group $C^{*}$-algebras. Let $\Gamma$ be an amenable discrete group. Let $\text{Sub}(\Gamma)$ denote the set of all subgroups in $\Gamma$. Consider the map
$\text{Sub}(\Gamma)$ $\longrightarrow\{0,1\}^{\Gamma}$
$X$ $\longmapsto\mathbbm{1}_{X}$
where $\mathbbm{1}_{X}\colon\Gamma\to\{0,1\}$ is the characteristic function, and equip $\{0,1\}^{\Gamma}$ with the product topology. We equip $\text{Sub}(\Gamma)$ with the topology induced by this map, known as the Chabauty topology, making $\text{Sub}(\Gamma)$ a Stone space (this topology agrees with the Fell topology of [14]).
For a subgroup $X\in\text{Sub}(\Gamma)$, write $\Gamma/X$ for the set of left-cosets. Let $\lambda_{\Gamma/X}\colon\Gamma\to B\big(\ell^{2}(\Gamma/X)\big)$ denote the associated quasi-regular representation as described in [3]. Concretely, we have
$\lambda_{\Gamma/X}(g)\delta_{hX}\coloneq\delta_{ghX}$ for $g\in\Gamma$ and $hX\in\Gamma/X$, where $\delta_{hX}\in\ell^{2}(\Gamma/X)$ denotes a standard basis vector. We also write $\lambda_{\Gamma/X}$ for the associated -representation on $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$.
Let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ be closed in the Chabauty topology and invariant under conjugation by elements of $\Gamma$. Following [16], we let the ideal $J_{\Gamma,\mathcal{X}}$ denote the $C^{*}$-kernel of $\lambda_{\Gamma/\mathcal{X}}\coloneq\oplus_{X\in\mathcal{X}}\lambda_{\Gamma/X}$ inside $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$. Concretely, we have $J_{\Gamma,\mathcal{X}}=\bigcap_{X\in\mathcal{X}}\text{ker}(\lambda_{\Gamma/X})$.
Definition 2.8.
Let $\Gamma$ be an amenable discrete group. Define $\mathcal{D}_{\Gamma}$ to be the set of all closed conjugation-invariant sets of subgroups $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ for which $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$. We will say $\Gamma$ satisfies the Density Property if $\mathcal{D}_{\Gamma}$ contains all closed conjugation-invariant sets of subgroups.
Let $G$ be an amenable second-countable étale groupoid. For $x\in G^{(0)}$ write $\mathcal{X}(x)\coloneq\pi^{-1}(x)\cap\tilde{G}^{(0)}_{{\rm{ess}}}$, where $\pi\colon\tilde{G}^{(0)}\to G^{(0)}$ is the canonical surjection and $\tilde{G}^{(0)}_{\rm{ess}}$ is the unit space of the essential Hausdorff cover (see Subsection 1.2). The isotropy groups of $G$ are amenable, and in this case it was shown in [16, Definition 5.1 & Theorem 5.5] that $J_{x}=J_{G^{x}_{x},\mathcal{X}(x)}$ for every $x\in G^{(0)}$. The following Theorem is then immediate.
Theorem 2.9.
Let $G$ be an amenable and second-countable étale groupoid. Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ if and only if $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ for all $x\in G^{(0)}$. In particular, if all the isotropy groups $G^{x}_{x}$ satisfy the Density Property, then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Remark 2.10.
Since $J_{\Gamma,\mathcal{X}}=\{0\}$ whenever $\{e\}\in\mathcal{X}$, in the above characterisation one only needs to check whether $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ whenever $\{x\}\notin\mathcal{X}(x)$. These units $x$ are shown in [16, Corollary 5.6] to be precisely the extremely dangerous points described in [24, Proposition 1.12]. Note also that the set of units $x\in G^{(0)}$ for which $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ is a $G$-invariant set.
It is clear that finite groups satisfy the Density Property, and hence $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for any amenable second-countable étale groupoid with finite isotropy groups.
Question 2.11.
Which discrete amenable groups satisfy the Density Property?
Assume that there exists a (discrete) countable amenable group $\Gamma$ which fails to have the Density Property, and let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ be a closed conjugation-invariant set of subgroups for which $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is not dense in $J_{\Gamma,\mathcal{X}}$, i.e. $\mathcal{X}\notin\mathcal{D}_{\Gamma}$. Note that the trivial subgroup, $\{e\}$, is not in $\mathcal{X}$ (amenability of $\Gamma$ ensures that the quasi-regular representation $\lambda_{\Gamma/\{e\}}$ has trivial kernel). By [16, Section 6], there exists an amenable second-countable étale groupoid $G$ and a unit element $x_{0}\in G^{(0)}$ for which $G^{x_{0}}_{x_{0}}=\Gamma$ and $\mathcal{X}(x_{0})=\mathcal{X}$. It follows from Theorem 2.9 that $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$. Therefore, we have established the following.
Corollary 2.12.
The intersection $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for all amenable second-countable étale groupoids if and only if all countable amenable groups satisfy the Density Property.
3. Groups with the Density Property
In this section, we study the set $\mathcal{D}_{\Gamma}$ for amenable discrete groups $\Gamma$, and show that certain classes of groups satisfy the Density Property (see Definition 2.8). Using Theorem 2.9, we are then able to deduce that $J\cap\mathscr{C}_{c}(G)$ is dense in the singular ideal $J$ for certain classes of groupoids.
Throughout this section we work with the reduced group $C^{*}$-algebra $C^{*}_{r}(\Gamma)$ since this is canonically isomorphic to the full $C^{*}$-algebra. The set of subgroups $\text{Sub}(\Gamma)$ will always be equipped with the Chabauty topology, and given a closed and conjugation-invariant subset $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ we write $J_{\Gamma,\mathcal{X}}\coloneq\bigcap_{X\in\mathcal{X}}\text{ker}\lambda_{\Gamma/X}\trianglelefteq C^{*}_{r}(\Gamma)$ (see Subsection 2.1).
Given a subgroup $\Lambda\subseteq\Gamma$, we write $E_{\Lambda}\colon C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Lambda)$ for the canonical conditional expectation and $I_{\Lambda}\colon C^{*}_{r}(\Lambda)\to C^{*}_{r}(\Gamma)$ for the canonical inclusion. If $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ is closed and conjugation-invariant, then $\Lambda\cap\mathcal{X}\coloneq\{\Lambda\cap X:X\in\mathcal{X}\}$ is a closed and conjugation-invariant set of subgroups of $\Lambda$ by [16, Proposition 7.4]. We isolate the following observations from [16, Propositions 7.4 & 7.7].
Lemma 3.1.
Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ be a closed and conjugation-invariant set of subgroups.
  1. (i)
    For any subgroup $\Lambda\subseteq\Gamma$ we have
    $I_{\Lambda}\big(J_{\Lambda,\Lambda\cap\mathcal{X}}\big)=J_{\Gamma,\mathcal{X}}\cap I_{\Lambda}\big(C^{*}_{r}(\Lambda)\big)\quad\text{and}\quad I_{\Lambda}\big(J_{\Lambda,\Lambda\cap\mathcal{X}}\cap\mathbb{C}[\Lambda]\big)=J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Lambda]$
  2. (ii)
    Let $N$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. Then,
    $E_{N}(J_{\Gamma,\mathcal{X}})=J_{N,\mathcal{X}}\quad\text{and}\quad E_{N}(J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma])=J_{N,\mathcal{X}}\cap\mathbb{C}[N].$
Proof.
For a proof of (i) see [16, Proposition 7.4]. For a proof of (ii), see the proof of [16, Proposition 7.7] (where $\Phi$ denotes $E_{N}$). ∎
We begin our study into the Density Property by showing it is preserved under countable increasing unions.
Proposition 3.2.
Suppose that an amenable discrete group $\Gamma$ is the countable increasing union of subgroups $\Lambda_{1}\subseteq\Lambda_{2}\subseteq\dots\subseteq\Gamma$. If $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ is closed, conjugation-invariant, and such that $\Lambda_{n}\cap\mathcal{X}\in\mathcal{D}_{\Lambda_{n}}$ for all $n\in\mathbb{N}$, then $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
In particular, if all $\Lambda_{n}$ have the Density Property, then so does $\Gamma$.
Proof.
For each $n\in\mathbb{N}$, write $\mathcal{X}_{n}\coloneq\Lambda_{n}\cap\mathcal{X}$. Since $\mathcal{X}_{n}\in\mathcal{D}_{\Lambda_{n}}$ for every $n\in\mathbb{N}$, Lemma 3.1 (i) implies that $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Lambda_{n}]$ is dense in $J_{\Gamma,\mathcal{X}}\cap I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$ for every $n\in\mathbb{N}$. Hence, $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]=\bigcup_{n}J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Lambda_{n}]$ is dense in $\bigcup_{n}J_{\Gamma,\mathcal{X}}\cap I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$. The countable increasing union $\bigcup_{n=1}^{\infty}I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$ is dense in $C^{*}_{r}(\Gamma)$, so $\bigcup_{n}J_{\Gamma,\mathcal{X}}\cap I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$ is dense in $J_{\Gamma,\mathcal{X}}$ by [11, Lemma III.4.1], completing the proof. ∎
Next, we prove a general lemma about ideals and dense -subalgebras.
Lemma 3.3.
Let A be a $C^{*}$-algebra, $\mathcal{A}\subseteq A$ a dense -subalgebra, and $I\trianglelefteq A$ an ideal. Then $I\cap\mathcal{A}$ is dense in $I$ if and only if $I\cap\mathcal{A}$ contains a net $(u_{\beta})$ of contractions such that $\lim_{\beta}\|a-au_{\beta}\|=0$ for all $a\in I$. Moreover, $u_{\beta}$ can be chosen to be positive.
We will need the following consequence of this lemma: For any closed and conjugation-invariant set $\mathcal{X}\subseteq\text{Sub}(\Gamma)$, we have $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ contains a net of positive contractions converging strongly to the identity in $J_{\Gamma,\mathcal{X}}$.
Proof.
Assume that $I\cap\mathcal{A}$ is dense in $I$. Since $I$ is a $C^{*}$-algebra, it contains an approximate unit $(u_{\beta})$. Then $I\cap\mathcal{A}$ is a dense -subalgebra, so each $u_{\beta}$ can be approximated by a sequence of positive elements in the unit ball of $I\cap\mathcal{A}$. By a diagonalisation argument, $I\cap\mathcal{A}$ contains a net of positive elements $I$ converging strongly to the identity in $I$. Conversely, assume that $I\cap\mathcal{A}$ contains a net of contractions converging strongly to the identity in $I$. Let $a\in I$ and $\varepsilon>0$. There exist $u\in I\cap\mathcal{A}$ and $b\in\mathcal{A}$ such that $\left\|u\right\|\leq 1$, $\left\|a-au\right\|<\frac{\varepsilon}{2}$ and $\left\|a-b\right\|<\frac{\varepsilon}{2}$. Then $bu\in I\cap\mathcal{A}$ and
$\left\|a-bu\right\|\leq\left\|a-au\right\|+\left\|(a-b)u\right\|<\varepsilon.$
We now prove our first application of this lemma.
Proposition 3.4.
Let $\Gamma$ be an amenable discrete group and suppose $\mathcal{X}_{1},\dots,\mathcal{X}_{n}\subseteq\mathrm{Sub}(\Gamma)$ are closed and conjugation-invariant sets of subgroups satisfying $\mathcal{X}_{i}\in\mathcal{D}_{\Gamma}$ for all $i=1,\dots,n$. Then, $\mathcal{X}\coloneq\bigcup_{i=1}^{n}\mathcal{X}_{i}$ satisfies $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
Proof.
By definition, we have $J_{\Gamma,\mathcal{X}}=\bigcap^{n}_{i=1}J_{\Gamma,\mathcal{X}_{i}}$. For each $i=1,\dots,n$ there exists, by Lemma 3.3, a net $(u_{i,\beta})$ of positive contractions in $J_{\Gamma,\mathcal{X}_{i}}\cap\mathbb{C}[\Gamma]$ converging strongly to the identity in $J_{\Gamma,\mathcal{X}_{i}}$ and hence in $J_{\Gamma,\mathcal{X}}\subseteq J_{\Gamma,\mathcal{X}_{i}}$. Therefore, by a diagonalisation argument, the net $(u_{\beta_{1},\beta_{2}..,\beta_{n}}):=(u_{1,\beta_{1}}\cdot u_{2,\beta_{n}}\cdot...\cdot u_{n,\beta_{n}})$ (equipped with the lexicographic partial ordering on $(\beta_{1},...,\beta_{n})$) has a subnet $(u_{\gamma})$ converging strongly to the identity in $J_{\Gamma,\mathcal{X}}$. The net $(u_{\gamma})$ lies in $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$, so $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Lemma 3.3. ∎
Next, we show the Density Property is preserved under taking quotients of groups by normal subgroups. For an amenable discrete group $\Gamma$ and a normal subgroup $N$, let $Q_{N}:C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Gamma/N)$ be the *-homomorphism induced by the quotient map $q_{N}:\Gamma\to\Gamma/N$. The map
$q^{*}_{N}:\text{Sub}(\Gamma/N)$ $\rightarrow\text{Sub}(\Gamma),\text{ }X\mapsto q_{N}^{-1}(X),$
is a continuous injection satisfying $\text{Conj}_{g}\circ q^{*}_{N}=q^{*}_{N}\circ\text{Conj}_{q_{N}(g)}$ for all $g\in\Gamma$, where $\text{Conj}_{g}$ denotes conjugation by $g$. Therefore, if $\mathcal{X}$ is a closed and conjugation-invariant set of subgroups of $\Gamma/N$, then $q^{*}_{N}(\mathcal{X})$ is a closed and conjugation-invariant set of subgroups of $\Gamma$.
Theorem 3.5.
Suppose $\Gamma$ is an amenable discrete group and $N$ a normal subgroup. If $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma/N)$ is a closed and conjugation-invariant set of subgroups of $\Gamma/N$, then
$J_{\Gamma,q_{N}^{*}(\mathcal{X})}=Q_{N}^{-1}(J_{\Gamma/N,\mathcal{X}}).$
Moreover, if $q^{*}_{N}(\mathcal{X})\in\mathcal{D}_{\Gamma}$, then $\mathcal{X}\in\mathcal{D}_{\Gamma/N}$. Consequently, if $\Gamma$ satisfies the Density Property, then so does $\Gamma/N$.
Proof.
For $X\in\mathcal{X}$, it is easy to see that $\lambda_{\Gamma/q_{N}^{-1}(X)}=\lambda_{(\Gamma/N)/\mathcal{X}}\circ Q_{N}$ and therefore $\ker(\lambda_{\Gamma/q_{N}^{-1}(X)})=Q_{N}^{-1}(\ker(\lambda_{(\Gamma/N)/\mathcal{X}}))$. Hence,
$J_{\Gamma,q^{*}_{N}(\mathcal{X})}=\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/q_{N}^{-1}(X)})=Q^{-1}_{N}\big(\bigcap_{X\in\mathcal{X}}\ker(\lambda_{(\Gamma/N)/X})\big)=Q^{-1}_{N}(J_{\Gamma/N,\mathcal{X}}).$
Now, suppose $q_{N}^{*}(\mathcal{X})\in\mathcal{D}_{\Gamma}$. By Lemma 3.3, we may choose a net of positive contractions $(u_{\lambda})\subseteq J_{\Gamma,q^{*}_{N}(\mathcal{X})}\cap\mathbb{C}[\Gamma]$ converging strongly to the identity in $J_{\Gamma,q^{*}_{N}(\mathcal{X})}$. Since $J_{\Gamma,q_{N}^{*}(\mathcal{X})}=Q_{N}^{-1}(J_{\Gamma/N,\mathcal{X}})$, $Q_{N}$ surjects $J_{\Gamma,q_{N}^{*}(\mathcal{X})}$ onto $J_{\Gamma/N,\mathcal{X}}$ and therefore $(Q_{N}(u_{\lambda}))\subseteq J_{\Gamma/N,\mathcal{X}}\cap\mathbb{C}[\Gamma/N]$ is a net of positive contractions converging strongly to the identity in $J_{\Gamma/N,\mathcal{X}}$. By Lemma 3.3, it follows that $\mathcal{X}\in\mathcal{D}_{\Gamma/N}$. ∎
We will now begin to prove the following result.
Theorem 3.6.
Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. Let $\Lambda\subseteq\Gamma$ be a subgroup such that $X\subseteq\Lambda$ for all $X\in\mathcal{X}$. Then, $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $\mathcal{X}\in\mathcal{D}_{\Lambda}$.
The following lemma is probably well-known to experts in group theory - it allows us to approximate an element $a\in C^{*}_{r}(\Gamma)$ by its restriction to finitely many cosets of a normal subgroup $N$. If $\Gamma$ is an amenable discrete group and $N$ a normal subgroup, for $gN\in\Gamma/N$, and $a\in C^{*}_{r}(\Gamma)$, define $a_{gN}\coloneq\delta_{g}\cdot E_{N}(\delta_{g^{-1}}\cdot a)\in C^{*}_{r}(\Gamma)$. Whenever $a_{gN}=0$ for all but finitely many cosets $gN$, the (finite) sum $\sum_{gN}a_{gN}=a$.
Lemma 3.7.
Let $\Gamma$ be an amenable discrete group and $N$ a normal subgroup. Choose a Følner net $(F_{i})$ in $\Gamma/N$ and define $\varphi_{i}:\Gamma/N\to\mathbb{C}$ via $gN\mapsto\frac{|gF_{i}\cap F_{i}|}{|F_{i}|}$. For any $a\in C^{*}_{r}(\Gamma)$ define
$a_{i}\coloneq\sum_{gN\in\Gamma/N}\varphi_{i}(gN)a_{gN}.$
Then, $\|a_{i}\|\leq\|a\|$ for all $i$ and the net $(a_{i})$ converges to $a$.
Proof.
[…]
The fact that $(F_{i})$ is a Følner net implies $(\tilde{\varphi}_{i}(g))$ converges to $1$ for all $g\in\Gamma$. Therefore, $(m_{i}(a))$ converges to $a$ whenever $a\in\mathbb{C}[\Gamma]$, and hence for all $a\in C^{*}_{r}(\Gamma)$ (by boundedness of $m_{i})$. ∎
Proof of Theorem 3.6.
Let $N$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. It suffices to prove the theorem in the case $\Lambda=N$, since $N$ is independent of the ambient group.
If $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$, then continuity of $E_{N}$ and Lemma 3.1 (ii) imply that $J_{N,\mathcal{X}}\cap\mathbb{C}[N]=E_{N}(J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma])$ is dense in $J_{N,\mathcal{X}}=E_{N}(J_{\Gamma,\mathcal{X}})$. This proves the “only if” direction.
Conversely, assume that $J_{N,\mathcal{X}}\cap\mathbb{C}[N]$ is dense in $J_{N,\mathcal{X}}$. By Lemma 3.3 there exists a net of contractions $(u_{beta})\subseteq J_{N,\mathcal{X}}\cap\mathbb{C}[N]$ converging strongly to the identity in $J_{N,\mathcal{X}}$. Since $J_{\Gamma,\mathcal{X}}\cap I_{N}\big(C^{*}_{r}(N)\big)=I_{N}\big(J_{N,\mathcal{X}}\big)$ by Lemma 3.1 (i), we can regard $(u_{\beta})$ as a net in $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$.
Let $a\in J_{\Gamma,\mathcal{X}}$ and fix a Følner net $(F_{i})$ for $\Gamma/N$, with $(a_{i})$ as in Lemma 3.7. By Lemma 3.1, $a_{gN}\in J_{\Gamma,\mathcal{X}}$ for all $gN\in\Gamma/N$. Since $a_{i}$ is a finite linear combination of the $a_{gN}$, it follows that $a_{i}\in J_{\Gamma,\mathcal{X}}$ for all $i$. Moreover, since $\delta_{g^{-1}}\cdot a_{gN}\in J_{N,\mathcal{X}}$, we have $\lim_{\beta}\|a_{gN}\cdot u_{\beta}-a_{gN}\|=0$ for all $gN\in\Gamma/N$ and therefore $\lim_{\beta}\|a_{i}\cdot u_{\beta}-a_{i}\|=0$ for all $i$. Since $(a_{i})$ converges to $a$, $\|a_{i}\|\leq\|a\|$ and $\|u_{\beta}\|\leq 1$ for all $i,\beta$, it follows from the above that $\lim_{\beta}\|a\cdot u_{\beta}-a\|=0$. Lemma 3.3 then implies $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$, completing the proof. ∎
Corollary 3.8.
If $\Gamma$ is an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ is a finite collection of normal subgroups, then $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
Proof.
By Proposition 3.4, we only need to establish the case where $\mathcal{X}=\{N\}$. By Theorem 3.6, it then suffices to establish the case where $\Gamma=N$. In this case, $J_{\Gamma,\{\Gamma\}}$ is the kernel of the trivial representation $\epsilon:C^{*}_{r}(\Gamma)\to\mathbb{C}$. Let $(F_{i})$ be a Følner net for $\Gamma$ and set $v_{i}=\frac{1}{|F_{i}|}\sum_{g\in F_{i}}\delta_{g}$. For any $g\in\Gamma$, we have $\lim_{i}\|\delta_{g}\cdot v_{i}-v_{i}\|=0$, hence (by linearity and density) $\lim_{i}\|a\cdot v_{i}-\epsilon(a)v_{i}\|=0$ for all $a\in C^{*}_{r}(\Gamma)$. By Kesten’s criteria ([5, Theorem 2.6.8(8)]) we have $\|v_{i}\|=1$ for all $i$, so the above limit implies $\lim_{i}\|a\cdot v_{i}\|=|\epsilon(a)|$ for all $a\in C^{*}_{r}(\Gamma)$. Set $u_{i}=\frac{1}{2}(\delta_{e}-v_{i})$ for all $i$. Then $(u_{i})$ is a net of contractions in $J_{\Gamma,\{\Gamma\}}\cap\mathbb{C}[\Gamma]$, and $\lim_{i}\|a\cdot u_{i}-a\|=\lim_{i}\frac{1}{2}\|a\cdot v_{i}\|=\frac{1}{2}|\epsilon(a)|=0$ for any $a\in J_{\Gamma,\{\Gamma\}}$. It follows from Lemma 3.3 that $J_{\Gamma,\{\Gamma\}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\{\Gamma\}}$, as desired. ∎
Theorem 3.6 shows that in order to determine whether a closed and conjugation-invariant set $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ belongs to $\mathcal{D}_{\Gamma}$, it suffices to assume that $\Gamma$ is generated by the subgroups $X\in\mathcal{X}$. The next result shows that we can also assume $\mathcal{X}$ is minimal in a certain sense.
For any discrete group $\Gamma$ and $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ closed conjugation-invariant, let $\mathcal{X}_{\text{min}}$ be the collection of $X\in\mathcal{X}$ such that $X^{\prime}\subseteq X$ and $X^{\prime}\in\mathcal{X}$ implies $X=X^{\prime}$. Then, $\mathcal{X}_{\text{min}}$ is conjugation-invariant but not necessarily closed.
Theorem 3.9.
Let $\Gamma$ be a discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. Then, for every $X\in\mathcal{X}$, there is $Y\in\mathcal{X}_{\mathrm{min}}$ such that $Y\subseteq X$. Consequently, for $\Gamma$ amenable, we have $J_{\Gamma,\mathcal{X}}=J_{\Gamma,\overline{\mathcal{X}_{\mathrm{min}}}}$ and $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $\overline{\mathcal{X}_{\mathrm{min}}}\in\mathcal{D}_{\Gamma}$.
Proof.
Fix $X\in\mathcal{X}$. Let $\mathcal{Z}_{X}=\{Y\in\mathcal{X}:Y\subseteq X\}$ and (partial) order $\mathcal{Z}_{X}$ by inclusion. Let $\mathcal{C}\subseteq\mathcal{Z}_{X}$ be a chain in $\mathcal{Z}_{X}$. Then, $\mathcal{C}$ is directed, so we can view it as a net, ordered by inclusion. By compactness of $\mathcal{X}$ the net $\mathcal{C}$ has a convergent subnet $(Y_{\beta})$, and let $Y_{*}$ be its limit point. As $(Y_{\beta})$ is a subnet, for every $Y^{\prime}\in\mathcal{C}$, there is $\beta_{0}$ such that $Y_{\beta_{0}}\subseteq Y^{\prime}$. Therefore, for all $\beta\geq\beta_{0}$, we have $Y_{\beta}\subseteq Y^{\prime}$ and hence $Y_{*}\subseteq Y^{\prime}$. Hence, $Y_{*}\subseteq Y^{\prime}\subseteq X$ for all $Y^{\prime}\in\mathcal{C}$. Therefore, every chain in $\mathcal{Z}_{X}$ has a minimal element. Zorn’s lemma implies $\mathcal{Z}_{X}$ contains a minimal element $Y\in\mathcal{\mathcal{X}}$.
If $X_{1}\subseteq X_{2}$ are subgroups of $\Gamma$, then $\ker(\lambda_{\Gamma/X_{1}})\subseteq\ker(\lambda_{\Gamma/X_{2}})$ (see [2, Appendix E and F]). Therefore, $J_{\Gamma,\mathcal{X}}=\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/X})=\bigcap_{X\in\mathcal{X}_{\text{min}}}\ker(\lambda_{\Gamma/X})$. By [3, Proposition 3.3], we have
$\bigcap_{X\in\mathcal{X}_{\text{min}}}\ker(\lambda_{\Gamma/X})=\bigcap_{X\in\overline{\mathcal{X}_{\text{min}}}}\ker(\lambda_{\Gamma/X})$, which completes the proof. ∎
We will now show that abelian groups satisfy the Density Property, and that collections of finite subgroups always belong to $\mathcal{D}_{\Gamma}$. First, we establish a key lemma that unifies the two cases.
Lemma 3.10.
Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. If every $X\in\mathcal{X}_{\mathrm{min}}$ is finitely generated, then $\mathcal{X}_{\mathrm{min}}$ is finite.
Proof.
For each $X\in\mathcal{X}_{\text{min}}$, the set $\mathcal{Z}(X)=\{X^{\prime}\in\mathcal{X}:X\subseteq X^{\prime}\}=\bigcap^{n}_{i=1}\{X^{\prime}\in\mathcal{X}:x_{i}\in X^{\prime}\}$, where $\{x_{1},...,x_{n}\}$ is a finite generating set for $X$. Therefore, $\mathcal{Z}(X)$ is open in $\mathcal{X}$ with respect to the Chabauty topology. Since $\mathcal{Z}(X)\cap\mathcal{X}_{\text{min}}=\{X\}$ for all $X\in\mathcal{X}_{\text{min}}$, the set $\mathcal{X}_{\text{min}}$ is discrete in the compact space $\mathcal{X}$, and is therefore finite. ∎
Now, we see our result for abelian groups and collections of finite subgroups as a simple consequence of Theorems 3.6, Theorem 3.9, Corollary 3.8 and Lemma 3.10.
Theorem 3.11.
Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. If either
  1. (I)
    $\Gamma$ is countable and abelian, or
  2. (II)
    every $X\in\mathcal{X}$ is finite,
then $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
Proof.
Let $\Gamma$ be a countable abelian group. By Proposition 3.2, it suffices to assume that $\Gamma$ is finitely generated. Then, every subgroup of $\Gamma$ is also finitely generated, so Lemma 3.10 implies that $\mathcal{X}_{\text{min}}$ is finite. Since $\Gamma$ is abelian, each $X\in\mathcal{X}_{\text{min}}$ is normal, and hence Corollary 3.8 implies $\mathcal{X}_{\text{min}}\in\mathcal{D}_{\Gamma}$. Theorem 3.9 then implies $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
In the case where every $X\in\mathcal{X}$ is finite, we have that $\mathcal{X}_{\text{min}}$ is finite (by Lemma 3.10) and each $X\in\mathcal{X}_{\text{min}}$ is finite. It follows from [16, Lemma 7.9] (see also the proof of [4, Theorem 4.7]) that the normal subgroup $N$ generated by all $X\in\mathcal{X}_{\text{min}}$ is finite. Since $C^{*}_{r}(N)=\mathbb{C}[N]$, it holds trivially that $\mathcal{X}_{\text{min}}\in\mathcal{D}_{N}$, so Theorems 3.6 and 3.9 imply $\mathcal{X}\in\mathcal{D}_{\Gamma}.$
The following Corollary is immediate by Theorems 2.9 and 3.11.
Corollary 3.12.
Let $G$ be an amenable and second-countable étale groupoid, and let $\tilde{G}$ denote the Hausdorff cover groupoid. Assume for every $x\in G^{(0)}$ one of the following holds.
  1. (I)
    The isotropy group $G_{x}^{x}$ is abelian.
  2. (II)
    The subgroups $X\in\mathcal{X}(x)$ are finite.
Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Note that condition (II) holds whenever all elements of $\tilde{G}^{(0)}_{\rm{ess}}$ are finite. The class of groupoids described in Corollary 3.12 (II) contains all amenable and second-countable étale groupoids with finite isotropy groups, and all groupoids of contracting self-similar groups (see [4, Corollary 7.13]). More generally, condition (II) is satisfied whenever the unit space of $G$ has finite source and range fibres. Therefore, within the class of amenable and second-countable étale groupoids, the class described in Corollary 3.12 (II) is larger than the class described in [4, Theorem 4.7].
4. Ample Groupoids
The main result of this section is that the algebraic singular ideal $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ for any ample groupoid. This reduces the question of whether $J_{\mathbb{C}}$ is dense in $J$ to the corresponding question for $J\cap\mathscr{C}_{c}(G)$. In order to prove this result we will need two preliminary lemmas.
Lemma 4.1.
Let $G$ be an étale groupoid, and let $f\in\mathscr{C}_{c}(G)$. Let $x\in G^{(0)}$, and assume that $f(g)=0$ for all $g\in G_{x}$. Then $f$ is continuous at each $g\in G_{x}$.
Proof.
Let $g\in G_{x}$ and let $(g_{\beta})$ be a net in $G$ converging to $g$. In order to prove that $\lim_{\beta}f(g_{\beta})=0$, it suffices to find a subnet $(g_{\gamma})$ satisfying $\lim_{\gamma}f(g_{\gamma})=0$ (since $(g_{\beta})$ is arbitrary). There exists a subnet $(g_{\gamma})$ for which the net $\iota(g_{\gamma})$ converges in $\tilde{G}$ (i.e. in the Fell topology). Let $\bm{g}\in\tilde{G}$ denote the limit. By (1.2) we have $\mathfrak{i}(f)(\bm{g})=\sum_{g\in\bm{g}}f(g)=0$ since $\bm{g}\subseteq G_{x}$. Moreover, $f(g_{\gamma})=\mathfrak{i}(f)\big(\iota(g_{\gamma})\big)\to\mathfrak{i}(f)(\bm{g})$ by continuity of $\mathfrak{i}(f)$. This completes the proof. ∎
Observations similar to the following lemma have appeared in both [4, Theorem 4.7] and [16, Proposition 5.19]. Our proof uses the same techniques as those developed in the aforementioned papers.
Lemma 4.2.
Let $G$ be an étale groupoid. Fix $f\in J\cap\mathscr{C}_{c}(G)$ and $x_{0}\in G^{(0)}$. Let $g_{1},\dotsc,g_{n}\in G_{x_{0}}$ be all points in $G_{x_{0}}$ on which $f$ is non-zero, and let $U_{1},\dotsc,U_{n}$ be open bisections such that $U_{i}$ contains $g_{i}$ for each $1\leq i\leq n$. Set $V\coloneq\bigcap_{i=1}^{n}s(U_{i})$. Then there exists an open subset $W\subseteq V$ containing $x_{0}$ such that
$f^{\psi}\coloneq\sum_{i=1}^{n}f(g_{i})(\psi\circ s|_{U_{i}})$ (4.1)
satisfies $f^{\psi}\in J$ for any $\psi\in C_{c}(W)$.
Proof.
Without loss of generality, we assume $s(U_{i})=V$ for all $1\leq i\leq n$. For each open set $W\subseteq V$, let $S^{W}_{i}\coloneq U_{i}\cap s^{-1}(W)$ for $1\leq i\leq n$. For $I\subseteq\{1,\dotsc,n\}$, set $\check{S}^{W}_{I}\coloneq(\bigcap_{i\in I}S^{W}_{i})\setminus(\bigcup_{j\notin I}S^{W}_{j})$ as well as $W_{I}\coloneq\text{int}\Big(\overline{s(\check{S}^{W}_{I})}^{G^{(0)}}\Big)$, where $\overline{(\,\cdot\,)}^{G^{(0)}}$ denotes closure in $G^{(0)}$. Our first claim is that there exists an open set $W\subseteq V$ containing $x_{0}$ such that
$W_{I}\neq\emptyset\text{ implies that }x_{0}\in\overline{W_{I}}^{G^{(0)}}\text{ for all }I\subseteq\{1,\dotsc,n\}.$ (4.2)
If $W=V$ satisfies (4.2), then we are done. Otherwise, choose a regular open subset $W$ of $G^{(0)}$ (i.e., $W=\text{int}\big(\overline{W}^{G^{(0)}}\big)$) such that $x_{0}\in W\subseteq V$ and $W\cap\overline{V_{I}}^{G^{(0)}}=\emptyset$ for all $I\subseteq\{1,\dotsc,n\}$ with $V_{I}\neq\emptyset$ and $x_{0}\notin\overline{V_{I}}^{G^{(0)}}$. Note that $\check{S}^{W}_{I}=\check{S}^{V}_{I}\cap s^{-1}(W)$, and therefore $s(\check{S}^{W}_{I})=s(\check{S}^{V}_{I})\cap W$. Since $W$ is open, we have $W_{I}=V_{I}\cap W$, and therefore $\overline{W_{I}}^{G^{(0)}}\cap W=\overline{V_{I}}^{G^{(0)}}\cap W$. Also $W_{I}=\text{int}\Big(\overline{s(\check{S}^{W}_{I})}^{G^{(0)}}\Big)\subseteq\text{int}\Big(\overline{s(\check{S}^{V}_{I})}^{G^{(0)}}\cap\overline{W}^{G^{(0)}}\Big)\subseteq V_{I}\cap\text{int}\big(\overline{W}^{G^{(0)}}\big)=V_{I}\cap W=\emptyset$ whenever $x_{0}\notin\overline{V_{I}}^{G^{(0)}}$. This completes the proof of the claim.
Define $\mathcal{I}\coloneq\{I\subseteq\{1,\dotsc,n\}\colon W_{I}\neq\emptyset\}$. Our second claim is that
$\sum_{i\in I}f(g_{i})=0\text{ for all }I\in\mathcal{I}.$ (4.3)
By [4, Lemma 4.1(ii)], $\overline{s({\rm supp}^{\circ}(f))}^{G^{(0)}}$ has empty interior. Let $I\in\mathcal{I}$, and set $\check{W}_{I}\coloneq W_{I}\setminus\overline{s({\rm supp}^{\circ}(f))}^{G^{(0)}}$. Then, $\check{W}_{I}$ is open and dense in $W_{I}$. Therefore $x_{0}$ lies in the closure of $\check{W}_{I}$ in $G^{(0)}$ by (4.2). As $s(\check{S}_{I})$ is dense in $W_{I}$ and because $\check{W}_{I}$ is open, we conclude that there is a net $(x_{\beta})$ in $\check{W}_{I}\cap s(\check{S}_{I})$ converging to $x_{0}$ in $G^{(0)}$. Let $g_{\beta}$ be the unique element of $\check{S}_{I}$ such that $s(g_{\beta})=x_{\beta}$. By passing to a subnet if necessary, we may assume that $(g_{\beta})$ converges to some $\bm{g}$ in $\tilde{G}$ (i.e., in the Fell topology). By construction, we have that $\bm{g}\cap\{g_{1},\dotsc,g_{n}\}=\{g_{i}\colon i\in I\}$. Therefore,
$\lim_{\beta}f(g_{\beta})=\lim_{\beta}\mathfrak{i}(f)(\iota(g_{\beta}))=\mathfrak{i}(f)(\bm{g})=\sum_{i\in I}f(g_{i}),$
where the third equality used (1.2). Now $s(g_{\beta})\in\check{W}_{I}$ implies that $g_{\beta}\notin{\rm supp}^{\circ}(f)$, so that $f(g_{\beta})=0$. We conclude that $\sum_{i\in I}f(g_{i})=0$, proving (4.3).
Take arbitrary $\psi\in C_{c}(W)$, and let $f^{\psi}$ be as in (4.1). We show that $f^{\psi}\in J$. For $I\in\mathcal{I}$ and $g\in\check{S}^{W}_{I}$ we have $f^{\psi}(g)=\sum_{i\in I}f(g_{i})\psi(s(g))=0$ using (4.3). It follows that ${\rm supp}^{\circ}(f^{\psi})\subseteq\bigcup_{I\,\in\,\mathcal{P}(\{1,\dotsc,n\})\setminus\mathcal{I}}\check{S}^{W}_{I}$ and therefore $s({\rm supp}^{\circ}(f^{\psi}))\subseteq\bigcup_{I\,\in\,\mathcal{P}(\{1,\dotsc,n\})\setminus\mathcal{I}}s(\check{S}^{W}_{I})$, where $\mathcal{P}$ denotes power set. Hence $s({\rm supp}^{\circ}(f^{\psi}))$ has empty interior because it is contained in a finite union of nowhere dense sets. This shows that $f^{\psi}\in J$, as desired. ∎
Remark 4.3.
Let $f\in J$, $x_{0}\in G^{(0)}$ and assume that there are only finitely many points $g_{1},\dotsc,g_{n}$ in the isotropy group $G^{x_{0}}_{x_{0}}$ on which $f$ non-zero. Let $U_{1},\dotsc,U_{n},V$ be as in the statement of Lemma 4.2. If in addition we assume that the isotropy group $G^{x_{0}}_{x_{0}}$ is amenable, then the conclusion of Lemma 4.2 still holds. That is, there exists an open set $W\subseteq V$ containing $x_{0}$ such that $f^{\psi}\in J$ for any $\psi\in C_{c}(W)$, where $f^{\psi}$ is defined as in (4.1). If the isotropy group $G^{x_{0}}_{x_{0}}$ is non-amenable, then there may exist no such $W$.
We now prove the main result of this section. Recall that an ample groupoid is an étale groupoid with a basis of compact open bisections. If $G$ is an ample groupoid, the Steinberg algebra $\mathbb{C}G$ over the complex numbers is defined in [26] as
$\mathbb{C}G\coloneq\text{span}\left\{\mathbbm{1}_{U}\colon U\subseteq G\text{ a compact open bisection}\right\}$
The algebraic singular ideal is the intersection $J_{\mathbb{C}}\coloneq J\cap\mathbb{C}G$. Since $\mathbb{C}G\subseteq\mathscr{C}_{c}(G)$ we have $J_{\mathbb{C}}\subseteq J\cap\mathscr{C}_{c}(G)$. For $f\in\mathscr{C}_{c}(G)$, define
$\left\|f\right\|_{I}\coloneq\max\left\{\sup_{x\in G^{(0)}}\sum_{g\in G_{x}}\left|f(g)\right|,\sup_{x\in G^{(0)}}\sum_{g\in G^{x}}\left|f(g)\right|\right\}.$
Theorem 4.4.
Let $G$ be an ample groupoid. Then $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to the norm $\left\|\cdot\right\|_{I}$.
The norm $\left\|\cdot\right\|_{I}$ dominates the full norm $\left\|\cdot\right\|_{C^{*}(G)}$ and hence any $C^{*}$-norm. Therefore, $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
Proof.
Let $f\in J\cap\mathscr{C}_{c}(G)$. Take a compact set $K\subseteq G$ with ${\rm supp}^{\circ}(f)$ contained in its interior. Whenever $\dot{f}\in\mathscr{C}_{c}(G)$ satisfies ${\rm supp}^{\circ}(\dot{f})\subseteq K$ we have $\left\|\dot{f}\right\|_{I}\leq C_{K}\left\|\dot{f}\right\|_{\infty}$ where $\left\|\cdot\right\|_{\infty}$ denotes the supremum norm, and $C_{K}$ is a constant depending only on $K$. Therefore, it suffices to find, for each $\varepsilon>0$, a function $f^{\prime}\in J_{\mathbb{C}}$ satisfying ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$.
Fix $\varepsilon>0$. Take $x\in s(K)$, and let $g_{1},\dotsc,g_{m}\in G_{x}$ be all points in $G_{x}$ on which $f$ is non-zero. By Lemma 4.2, there exists a compact open neighbourhood $U_{x}\subseteq G^{(0)}$ of $x$, and compact open bisections $U_{1},\dotsc,U_{m}\subseteq K$ satisfying the following: $g_{i}\in U_{i}$ and $s(U_{i})=U_{x}$ for all $i=1,\dotsc,m$, and the function $f^{x}$ defined by
$f^{x}\coloneq\sum_{i=1}^{m}f(g_{i})\mathbbm{1}_{U_{i}}$
lies in $J_{\mathbb{C}}$. Clearly ${\rm supp}^{\circ}(f^{x})\subseteq K$. We have $f|_{G_{x}}=f^{x}|_{G_{x}}$ and ${\rm supp}^{\circ}(f-f^{x})\subseteq K$, a compact set. Therefore, by Lemma 4.1, there exists a compact open neighbourhood $W_{x}\subseteq U_{x}$ of $x$ such that $\left|f(g)-f^{x}(g)\right|\leq\varepsilon$ whenever $s(g)\in W_{x}$. The sets $\{W_{x}\}_{x\in s(K)}$ form an open cover for the compact set $s(K)$, so select a finite subcover $W_{x_{1}},\dotsc,W_{x_{n}}$. Removing intersections if necessary, we may assume that the $W_{x_{j}}$ are pairwise disjoint (this is possible since the $W_{x_{i}}$ are contained in the Hausdorff space $G^{(0)}$, and are thus clopen in $G^{(0)}$). Define
$f^{\prime}\coloneq\sum_{j=1}^{n}(\mathbbm{1}_{W_{x_{j}}}\circ s)f^{x_{j}}.$
Then $f^{\prime}\in J_{\mathbb{C}}$ satisfies ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$, as required. ∎
We apply the previous theorem to the class of groupoids arising from contracting self-similar groups.
Corollary 4.5.
Let $G$ be the groupoid arising from the action of a contracting self-similar group on a finite alphabet. Then $J_{\mathbb{C}}$ is dense in $J$.
Proof.
By Theorem 4.4 it suffices to check that $J\cap\mathscr{C}_{c}(G)$ is dense in $J$. By [23] and [15] the groupoid $G$ is second-countable and amenable. Moreover, [4, Corollary 7.13] implies that every element of $\tilde{G}^{(0)}$ is finite. The result follows by Corollary 3.12 (II). ∎

### v2

@@
 As an application, we characterise when the singular functions in Connes’ algebra are dense in the singular ideal in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras.
+We then show this density property holds for all $C^{*}$-algebras of groupoids with finite-by-nilpotent isotropy groups.
 Key words and phrases:
@@
 Many groupoids arising from dynamics and geometry are non-Hausdorff, and it is therefore important to develop a robust theory.
-Indeed, there are natural examples of groupoids of germs, groupoids arising from self-similar groups, and groupoids arising from foliations which are non-Hausdorff ([10], [21], [22], [23]).
+Indeed, there are natural examples of groupoids of germs, groupoids arising from self-similar groups, and groupoids arising from foliations that are non-Hausdorff ([11], [27], [28], [29]).
 In contrast with the Hausdorff case, the functions in the reduced $C^{*}$-algebra $C^{*}_{r}(G)$ of a non-Hausdorff groupoid $G$ are not necessarily continuous.
-Indeed, there can exist (non-zero) functions in $C^{*}_{r}(G)$ whose set of non-zero values has empty interior in $G$.
+In fact, there can exist (non-zero) functions in $C^{*}_{r}(G)$ whose set of non-zero values has empty interior in $G$.
 The set of all such functions is known as the singular ideal $J$ in $C^{*}_{r}(G)$.
@@
 Characterisations of simplicity and the ideal intersection property have been obtained for the quotient by this ideal ([10], [21]), which is known as the essential groupoid $C^{*}$-algebra ([22], [13]).
-Therefore, the only obstacle to understanding these properties for the reduced $C^{*}$-algebra is the singular ideal.
-Much work has been done to find conditions that ensure the singular ideal $J$ vanishes (see [9], [19], [24], [15]).
-Recently in [16] vanishing of the singular ideal was characterised in terms of a property of certain collections of subgroups in isotropy groups of $G$.
-In [4] it is characterised when the singular ideal $J$ has trivial intersection with the underlying groupoid -algebra, $\mathscr{C}_{c}(G)$ known as Connes’ algebra (first defined in [10] for groupoids arising from foliations).
-When the singular ideal is non-zero, less is known about its structure.
+Therefore, the singular ideal is the only obstacle to understanding these properties for $C^{*}_{r}(G)$.
+Much work has been done to find conditions that ensure that the singular ideal $J$ vanishes (see [10], [22], [30], [15]).
+In [5, Theorem A] (see also [19, Theorem F]) it is characterised when the singular ideal has trivial intersection with the underlying groupoid -algebra $\mathscr{C}_{c}(G)$, known as Connes’ algebra (first defined in [11] for groupoids arising from foliations).
+Recently, [19, Theorem B] characterised vanishing of the singular ideal in terms of vanishing of explicit ideals in the isotropy group $C^{*}$-algebras.
+When the singular ideal is non-zero, little is known about its structure.
 For example, it is not known in general whether $J\neq\{0\}$ implies $J\cap\mathscr{C}_{c}(G)\neq\{0\}$.
-This was shown for groupoids with finite “non-Hausdorfness” in [4] and in [16] the problem was shown to be equivalent to whether group $C^{*}$-algebras satisfy the Intersection Property.
-This reduction of the problem led to its solution for groupoids with isotropy groups that are direct limits of virtually torsion free solvable groups [16].
-A related question asks for which groupoids is $J\cap\mathscr{C}_{c}(G)$ dense in $J$ ([4, Question 4.11(III)]).
-Progress could be made towards understanding the structure of $J$ if it has a dense subalgebra of functions in $\mathscr{C}_{c}(G)$ (since functions in $\mathscr{C}_{c}(G)$ are easier to understand).
-In recent work ([20]), it is shown that there exist non-amenable groupoids for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$.
-The following question remains open however.
-Question.
+This implication was shown to hold for groupoids that are “finitely non-Hausdorff” in [5, Theorem C].
+In [19], the problem was reduced to one about group $C^{*}$-algebras, and this led to a positive answer in many cases ([19, Theorem H]).
+In order to understand the structure of $J$, it is necessary to ask the following deeper question (see [5, Question 4.11(III)]): for which groupoids is $J\cap\mathscr{C}_{c}(G)$ dense in $J$?
+In recent work ([25]), it was shown that there exist non-amenable étale groupoids for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$.
+The following remains open however.
+Question 1.
 Is $J\cap\mathscr{C}_{c}(G)$ dense in $J$ for any amenable étale groupoid $G$?
-The goal of our paper is to study the question highlighted above. Let us outline the main achievements of the paper. All results listed here are for amenable second-countable étale groupoids.
-  • We show that ideals in the groupoid $C^{*}$-algebra are determined by their isotropy fibres. This generalises [8, Theorem 2.10] to the setting of non-Hausdorff groupoids.
-  • Following the approach in [16], we characterise when $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras.
-  • We describe some large classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
-In particular, we show density holds for groupoids with abelian isotropy groups, and groupoids arising from contracting self-similar groups.
-We now state our main results and describe them in more detail.
-Our first result concerns ideals in groupoid $C^{*}$-algebras.
-In [8] it is shown that ideals $I\trianglelefteq C^{*}(G)$ in a groupoid $C^{*}$-algebra are mapped to ideals in the isotropy group $C^{*}$-algebras by the restriction maps $C^{*}(G)\to C^{*}(G^{x}_{x})$.
+A positive answer to this question would allow progress to be made towards understanding the structure of $J$ (since the functions in $\mathscr{C}_{c}(G)$ are easier to understand).
+It is therefore natural to ask a follow up question: can we explicitly describe elements of $J\cap\mathscr{C}_{c}(G)$?
+The goal of our paper is to study these questions following the approach in [19] of studying the singular ideal through its isotropy fibres.
+Let us outline the main achievements and methods of the paper.
+The first four points listed here are for amenable and second-countable étale groupoids.
+  • We show that ideals in a (non-Hausdorff) groupoid $C^{*}$-algebra are determined by their isotropy fibres.
+We derive this result from the Hausdorff case (proven in [9]) through a novel application of the Hausdorff cover (see [5]).
+  • We use this result to characterise when $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras.
+This reduces Question 1 to a property of discrete amenable groups, which we call the Density Property.
+  • We show the Density Property is preserved under various operations on the group.
+We pair this with structural properties of the space of subgroups (equipped with the Chabauty topology) to prove that the Density Property holds for all finite-by-nilpotent groups.
+  • This reduction allows us to prove that $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for groupoids with finite-by-nilpotent isotropy groups, groupoids arising from contracting self-similar group actions, and for amenable étale group bundles.
+  • We describe an explicit family of functions whose span is always dense in $J\cap\mathscr{C}_{c}(G)$.
+  • For ample groupoids, we completely describe the structure of functions in the algebraic singular ideal $J_{\mathbb{C}}$, and prove $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$.
+$\bullet$
+Let us state our main results in more detail.
+The first four results describe the progress made towards answering Question 1.
+In [9] it was shown that ideals $I\trianglelefteq C^{*}(G)$ in a groupoid $C^{*}$-algebra are mapped to ideals in the isotropy group $C^{*}$-algebras by the restriction maps $C^{*}(G)\to C^{*}(G^{x}_{x})$.
 These images are known as the isotropy fibres of the ideal, and are denoted $(I_{x})_{x\in G^{(0)}}$.
-It is shown in [8, Theorem 2.10] that for amenable second-countable Hausdorff étale groupoids, ideals in the groupoid $C^{*}$-algebra are determined by their isotropy fibres.
+For amenable second-countable Hausdorff étale groupoids, ideals in the groupoid $C^{*}$-algebra are determined by their isotropy fibres (see [9, Theorem 2.10]).
 We generalise this result to non-Hausdorff groupoids.
@@
 Let $G$ be an amenable and second-countable (non-Hausdorff) étale groupoid and suppose $I$ and $K$ are ideals in $C^{*}(G)$. Then, $I=K$ if and only if $I_{x}=K_{x}$ for all $x\in G^{(0)}$.
-We apply this result to the singular ideal $J$ in the reduced $C^{*}$-algebra of a non-Hausdorff groupoid.
-The following Theorem reduces the density question for $J\cap\mathscr{C}_{c}(G)$ in $J$ to one about density of ideals in isotropy group $C^{*}$-algebras.
-Theorem B (See Proposition 2.6).
-Let $G$ be an amenable and second-countable étale groupoid. Then, $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ if and only if $J_{x}\cap\mathbb{C}[G^{x}_{x}]$ is dense in $J_{x}$ for all $x\in G^{(0)}$.
-The isotropy fibres $J_{x}$ are calculated explicitly in [16, Theorem 5.5].
-Moreover, when $G$ is covered by countably many bisections and $G^{x}_{x}$ is amenable, it is shown $J_{x}$ equals the ideal $J_{G^{x}_{x},\mathcal{X}(x)}$ defined as the intersection of the kernels of quasi-regular representations associated to subgroups of $G^{x}_{x}$ in a certain a collection $\mathcal{X}(x)$.
-We refer the reader to Subsection 2.1 for the definition of $\mathcal{X}(x)$ and $J_{G^{x}_{x},\mathcal{X}(x)}$.
-The following is then a rephrasing of Theorem B.
-Theorem C (See Theorem 2.9).
+We apply this result to the singular ideal $J$ in the reduced $C^{*}$-algebra of a non-Hausdorff groupoid (see Proposition 3.1).
+The isotropy fibres $J_{x}$ are calculated explicitly in [19, Theorem 5.5], where it is shown that $J_{x}$ is equal to an intersection of kernels of quasi-regular representations associated with subgroups of $G^{x}_{x}$ in a certain collection $\mathcal{X}(x)$.
+This ideal is denoted $J_{G^{x}_{x},\mathcal{X}(x)}\trianglelefteq C^{*}(G^{x}_{x})$.
+We refer the reader to Definition 1.3 and Section 3 for the definitions of $\mathcal{X}(x)$ and $J_{G^{x}_{x},\mathcal{X}(x)}$.
+Theorem B (See Theorem 3.4).
 Let $G$ be an amenable and second-countable étale groupoid.
@@
 In particular, if all the isotropy groups $G^{x}_{x}$ satisfy the Density Property, then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
-In the above theorem, the Density Property is a property of discrete amenable groups we introduce in Definition 2.8.
-Finite groups satisfy the Density Property automatically, and in Proposition 3.11 we show that all abelian groups have the Density Property.
-Currently, we do not know of any discrete amenable group that fails the Density Property.
-We remark that it is a corollary of Theorem C and a non-Hausdorff groupoid construction in [16, Section 6] that if there is such a group, then there is an amenable groupoid for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$ (see Corollary 2.12).
-By the definition of the ideals $J_{G^{x}_{x},\mathcal{X}(x)}$, Theorem C asserts that density of $J\cap\mathscr{C}_{c}(G)$ in $J$ can be understood entirely in terms of quasi-regular representations of isotropy groups.
+In Theorem B, the Density Property is a property of discrete amenable groups introduced in Definition 3.3.
+Finite groups trivially satisfy the Density Property, and in Theorem 4.16 we show that all finite-by-nilpotent groups have the Density Property.
+Currently, we do not know of any discrete amenable group that fails to have the Density Property.
+We remark that it is a consequence of Theorem B and a non-Hausdorff groupoid construction in [19, Section 6] that if there exists such a group, then there also exists an amenable and second-countable étale groupoid for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$ (see Corollary 3.7).
+By the definition of the ideals $J_{G^{x}_{x},\mathcal{X}(x)}$, Theorem B asserts that density of $J\cap\mathscr{C}_{c}(G)$ in $J$ can be understood entirely in terms of quasi-regular representations of isotropy groups.
 In light of this, we are able to describe some explicit classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
-Theorem D (See Corollary 3.12).
-Let $G$ be an amenable and second-countable étale groupoid. Assume for every $x\in G^{(0)}$ one of the following holds.
+Theorem C (See Corollary 4.17).
+Let $G$ be an amenable and second-countable étale groupoid. For each $x\in G^{(0)}$ let $\mathcal{X}(x)$ be as in Definition 1.3, and assume that one of the following holds.
   1. (I)
-    The isotropy group $G_{x}^{x}$ is abelian.
+    The isotropy group $G_{x}^{x}$ is finite-by-nilpotent.
   2. (II)
-    The subgroups $X\in\mathcal{X}(x)$ are finite.
+    The subgroups $X\in\mathcal{X}(x)$ are all finite.
+  3. (III)
+    The subgroups $X\in\mathcal{X}(x)$ are all co-finite in $G_{x}^{x}$.
+  4. (IV)
+    The subgroups $X\in\mathcal{X}(x)$ are all normal in $G^{x}_{x}$.
 Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
-Theorem D covers many important classes of groupoids.
-In particular, any groupoid of germs associated with the action of a countable abelian group will satisfy (I).
-All groupoids with finite isotropy groups will satisfy (II).
-More generally, condition (II) is satisfied whenever the unit space of $G$ has finite source and range fibres.
-In the setting of non-Hausdorff ample groupoids, one can intersect $J$ with the complex Steinberg algebra $\mathbb{C}G$ (see [26]) to obtain what is known as the algebraic singular ideal $J_{\mathbb{C}}$.
+Here, a countable group $\Gamma$ is finite-by-nilpotent if there exists a finite normal subgroup $N\trianglelefteq\Gamma$ for which the quotient $\Gamma/N$ is nilpotent.
+Theorem C covers many important classes of groupoids.
+In particular, any groupoid of germs associated with the action of a countable finite-by-nilpotent group will satisfy (I) for all $x\in G^{(0)}$.
+Condition (II) is satisfied for all $x\in G^{(0)}$ whenever the closure of the unit space of $G$ has finite source and range fibres.
+This is the hypothesis assumed in [5, Theorem C] when proving the weaker property that $J\neq\{0\}$ implies $J\cap\mathscr{C}_{c}(G)\neq\{0\}.$
+The next result considers étale group bundles – that is, étale groupoids whose range and source maps are equal.
+Historically, étale group bundles have proven to be a good resource for providing counterexamples to conjectures (see [18], [37] for example).
+It was shown in [25] that there exist non-amenable étale group bundles for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$.
+A direct corollary of Theorem C (IV) is that density always holds for amenable étale group bundles.
+Corollary D (See Corollary 4.18).
+If $G$ is an amenable and second-countable étale group bundle, then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
+$\bullet$
+The remaining results describe the structure of elements in $J\cap\mathscr{C}_{c}(G)$.
+In Section 5 we describe an explicit set whose span is dense in the subalgebra $J\cap\mathscr{C}_{c}(G)$.
+Elements in this spanning set are weighted sums of functions $\psi\in C_{c}\big(G^{(0)}\big)$ lifted by the source map, and are built using purely topological and algebraic data coming from the groupoid and its isotropy groups.
+This provides an explicit dense subspace of the singular ideal $J$ for groupoids as described in Theorem C.
+Theorem E (See Theorem 5.5).
+Let $G$ be an étale groupoid that is covered by countably many open bisections. For $x\in G^{(0)}$ let $b\colon G^{x}_{x}\to\mathbb{C}$ be a function with finite support $g_{1},\dots,g_{n}$ such that
+(0.1) $\sum_{g\in kX}b(g)=0\text{ for all }k\in G^{x}_{x}\text{ and }X\in\mathcal{X}(x).$
+Then, given any $h\in G_{x}$ and any open bisections $V_{1},\dots,V_{n}$ containing $hg_{1},\dots,hg_{n}$, there exists an (explicitly defined) open neighbourhood $W\subseteq\bigcap_{i=1}^{n}s(V_{i})$ of $x$ such that
+(0.2) $f^{\psi}\coloneq\sum_{i=1}^{n}b(g_{i})(\psi\circ s|_{V_{i}})$
+belongs to $J\cap\mathscr{C}_{c}(G)$ for all $\psi\in C_{c}(W)$. Furthermore, the linear span of such functions $f^{\psi}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
+We remark that an element $b\in\mathbb{C}[G^{x}_{x}]$ satisfies (0.1) if and only if, for every subgroup $X\in\mathcal{X}(x)$, $b$ belongs to the kernel of the associated quasi-regular representation $\lambda_{G^{x}_{x}/X}$.
+$\bullet$
+We now turn our attention to non-Hausdorff ample groupoids, where one can intersect $J$ with the complex Steinberg algebra $\mathbb{C}G$ (see [33]) to obtain what is known as the algebraic singular ideal $J_{\mathbb{C}}$.
 A characterisation for the vanishing of $J_{\mathbb{C}}$ is given in [5], and together with [10] and [34] this leads to a complete characterisation of simplicity for Steinberg algebras (see also [35] and [15]).
-In this setting, one can ask the following density question: for which ample groupoids is $J_{\mathbb{C}}$ dense in $J$?
+In this paper we look beyond the vanishing of $J_{\mathbb{C}}$ and describe an explicit spanning set analogous to that in Theorem E.
+We stress the following: the structure of functions in the algebraic singular ideal $J_{\mathbb{C}}$ is now fully understood.
+Theorem F (See Theorem 6.1).
+Let $G$ be an ample groupoid that is covered by countably many open bisections.
+Let $x,b,h,V_{1},\dots,V_{n}$ be as in Theorem E.
+There exists a compact open neighbourhood $W\subseteq\bigcap_{i=1}^{n}s(V_{i})$ of $x$ such that the function
+(0.3) $f^{W}\coloneq\sum_{i=1}^{n}b(g_{i})\mathbbm{1}_{U_{i}}$
+belongs to $J_{\mathbb{C}}$, where $U_{i}\coloneq V_{i}\cap s^{-1}(W)$. Furthermore, $J_{\mathbb{C}}$ is equal to the linear span of functions of the form $f^{W}$.
+For ample groupoids, one can ask the following density question: for which ample groupoids is $J_{\mathbb{C}}$ dense in $J$?
 Indeed, this is presented as an open problem in [15].
-In order to answer this question, we show it suffices to study density of $J\cap\mathscr{C}_{c}(G)$ in $J$.
-Theorem E (See Theorem 4.4).
+The following theorem implies that $J_{\mathbb{C}}$ is dense in $J$ if and only if $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
+Theorem G (See Theorem 6.3).
 Let $G$ be an ample groupoid. Then $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
@@
 These groupoids can be non-Hausdorff, and can also have non-vanishing singular ideal.
-Using our results, we answer the density question for groupoids arising from contracting self-similar groups.
-Corollary F (See Corollary 4.5).
-Let $G$ be the groupoid arising from the action of a contracting self-similar group on a finite alphabet. Then $J_{\mathbb{C}}$ is dense in $J$.
+For a groupoid arising from a contracting self-similar group, it was shown in [15] that $J_{\mathbb{C}}=\{0\}$ if and only if $J=\{0\}$.
+Using our results, we show that $J_{\mathbb{C}}$ is in fact dense in $J$ for all groupoids in this class.
+Corollary H (See Corollary 6.4).
+Let $G$ be the groupoid arising from a contracting self-similar group action. Then $J_{\mathbb{C}}$ is dense in $J$.
 In [25], examples of groupoids arising from self-similar actions on infinite alphabets are described for which $J_{\mathbb{C}}$ is not dense in $J$.
+$\bullet$
+The paper is structured as follows.
+In Section 1, we give the necessary background on the Hausdorff cover groupoid and on isotropy fibres of ideals.
+In Section 2 we prove Theorem A, and then apply this result to the singular ideal $J$ in Section 3.
+Section 4 studies the Density Property for groups and ends by describing some classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
+Sections 5 studies the structure of functions in $J\cap\mathscr{C}_{c}(G)$.
+Section 6 is exclusively concerned with ample groupoids and describes the structure of functions in the algebraic singular ideal $J_{\mathbb{C}}$.
 Acknowledgments
@@
 Hume was supported by NSERC Discovery Grant RGPIN-2021-03834.
-Gonzales would like to thank Xin Li for helpful discussions on ideas related to this paper.
+Gonzales was supported by EPSRC grant EP/T517896/1.
+Gonzales would like to thank Xin Li for helpful discussions on topics related to this paper.
 1. Preliminaries
@@
 By an étale groupoid $G$, we mean a locally compact topological groupoid for which the unit space $G^{(0)}$ is Hausdorff, and the range $r\colon G\to G^{(0)}$ and source $s\colon G\to G^{(0)}$ maps are local homeomorphisms.
+Note that this ensures that the unit space $G^{(0)}$ is open in $G$.
 This paper will only consider étale groupoids, and primarily concerns non-Hausdorff groupoids (all results also hold for Hausdorff groupoids, but often for trivial reasons i.e.
@@
 An open subset $U\subseteq G$ will be called an open bisection if the restrictions $r|_{U}\colon U\to r(U)$ and $s|_{U}\colon U\to s(U)$ are homeomorphisms.
-Note that any étale groupoid has a basis of open bisections.
-We define
+Note that the open bisections form a basis for the topology of any étale groupoid.
+We define Connes’ algebra as follows:
 $\mathscr{C}_{c}(G)\coloneq\text{span}\{f\colon G\to\mathbb{C}\colon f|_{U}\in C_{c}(U)\text{ and }f|_{G\setminus U}=0\text{ for some open bisection }U\subseteq G\}.$
 If the groupoid $G$ is Hausdorff, then $\mathscr{C}_{c}(G)=C_{c}(G)$, the set of all compactly supported continuous functions on $G$.
-However if $G$ is not Hausdorff, then functions in $\mathscr{C}_{c}(G)$ need not be continuous.
+If $G$ is not Hausdorff, then the functions in $\mathscr{C}_{c}(G)$ need not be continuous.
 We equip $\mathscr{C}_{c}(G)$ with the structure of a -algebra as follows.
@@
 The full groupoid $C^{*}$-algebra $C^{*}(G)$ of $G$ is defined to be the completion of $\mathscr{C}_{c}(G)$ with respect to the norm $\|f\|_{C^{*}(G)}=\sup_{\pi}\|\pi(f)\|$, where the supremum is taken over all -algebra representations $\pi\colon\mathscr{C}_{c}(G)\to B(H)$, for some Hilbert space $H$.
-Elements of the reduced groupoid $C^{*}$-alegbra $C^{*}_{r}(G)$ can be viewed as bounded Borel functions on $G$ in a canonical way (see [6, Lemma 3.27], [25, Proposition II.4.2]).
+Elements of the reduced groupoid $C^{*}$-alegbra $C^{*}_{r}(G)$ can be viewed as bounded Borel functions on $G$ in a canonical way (see [7, Lemma 3.15], [31, Proposition II.4.2]).
 From now on, we will always treat elements of $C^{*}_{r}(G)$ as functions on $G$ in this way.
@@
 The singular ideal $J$ is a (closed) ideal in $C^{*}_{r}(G)$.
-For $a\in C^{*}_{r}(G)$, ${\rm supp}^{\circ}(a)$ has empty interior in $G$ if and only if $s\big({\rm supp}^{\circ}(a)\big)$ has empty interior in $G^{(0)}$ (see [1, Proposition 4.6], [4, Lemma 4.1(iii)]).
+For $a\in C^{*}_{r}(G)$, we have $a\in J$ if and only if $s\big({\rm supp}^{\circ}(a)\big)$ has empty interior in $G^{(0)}$ (see [2, Proposition 4.6], [5, Lemma 4.1(iii)]).
+Now assume that the groupoid $G$ is covered by countably many open bisections.
+We say a unit $x\in G^{(0)}$ is Hausdorff if for any $g\in G_{x}^{x}\setminus\{x\}$ there exists an open neighbourhood $U$ of $g$ such that $U\cap G^{(0)}=\emptyset$.
+It is clear that $G$ is Hausdorff if and only if all units are Hausdorff.
+We let $C\subseteq G^{(0)}$ denote the set of Hausdorff units.
+By [22, Lemma 7.15], $C$ is dense in $G^{(0)}$ (the Hausdorff points complement the so-called dangerous points of [22]).
+An element $a\in C^{*}_{r}(G)$ belongs to the singular ideal $J$ if and only if $s\big({\rm supp}^{\circ}(a)\big)\subseteq G^{(0)}\setminus C$ (see [22, Proposition 7.18]).
+$\bullet$
 1.2. Hausdorff cover
@@
 Elements of $\tilde{G}$ can be described explicitly as follows.
-A non-empty subset $\bm{g}\subseteq G$ lies in $\tilde{G}$ if and only if there exist a net $(g_{i})$ in $G$ such that $\bm{g}$ is the set of limit points of $(g_{i})$, and every limit point of a subnet of $(g_{i})$ (cluster point) is a limit point of $(g_{i})$.
-Moreover, given a net $(g_{i})$ in $G$, the image net $\iota(g_{i})$ converges to $\bm{g}\in\tilde{G}$ in the topology of $\tilde{G}$ if and only if the conditions above are satisfied.
+A non-empty subset $\bm{g}\subseteq G$ belongs to $\tilde{G}$ if and only if there exists a net $(g_{\alpha})$ in $G$ whose set of limit points is precisely $\bm{g}$, and such that every subnet of $(g_{\alpha})$ has its limit points contained in (and hence equal to) $\bm{g}$.
+Moreover, given a net $(g_{\alpha})$ in $G$, the image net $\iota(g_{\alpha})$ converges to $\bm{g}\in\tilde{G}$ in the Fell topology if and only if the conditions above are satisfied.
+We will use the following fact several times: any convergent net $(g_{\alpha})$ in $G$ has a subnet $(g_{\beta})$ such that $\iota(g_{\beta})$ is convergent in $\tilde{G}$.
 Define $\tilde{G}^{(0)}\coloneq\{\bm{g}\in\tilde{G}\colon\bm{g}\cap G^{(0)}\neq\emptyset\}$.
-This set will be the unit space of $\tilde{G}$ with the structure described below.
+This set will be the unit space of $\tilde{G}$ with the groupoid structure described below.
 The canonical inclusion map $\iota\colon G\hookrightarrow\tilde{G}$ restricts to an inclusion $G^{(0)}\hookrightarrow\tilde{G}^{(0)}$, and the image $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$.
-For each $\bm{x}\in\tilde{G}^{(0)}$ there is a unique element $\pi(\bm{x})\in\bm{x}\cap G^{(0)}$, and the map $\pi\colon\tilde{G}^{(0)}\to G^{(0)}$ is a continuous surjection.
-Since $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$, continuity of the groupoid operations on $G$ implies that any element $\bm{x}\in\tilde{G}^{(0)}$ is a subgroup of the isotropy group $G_{\pi(\bm{x})}^{\pi(\bm{x})}$.
+In particular, any element $\bm{x}\in\tilde{G}^{(0)}$ is contained in the closure $\overline{G^{(0)}}$.
+For each $\bm{x}\in\tilde{G}^{(0)}$ there is a unique element $\pi(\bm{x})\in\bm{x}\cap G^{(0)}$, and the map $\pi\colon\tilde{G}^{(0)}\to G^{(0)}$ is a proper continuous surjection (since $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$).
+Continuity of the groupoid operations on $G$ implies that any element $\bm{x}\in\tilde{G}^{(0)}$ is a subgroup of the isotropy group $G_{\pi(\bm{x})}^{\pi(\bm{x})}$.
 We will often write $\bm{x}=X$ to distinguish elements $x\in G^{(0)}$ from $\bm{x}\in\tilde{G}^{(0)}$.
 We now describe the groupoid operations on $\tilde{G}$.
+With these operations, $\tilde{G}$ becomes a Hausdorff étale groupoid.
 Note that $r(g_{1})=r(g_{2})$ and $s(g_{1})=s(g_{2})$ whenever $g_{1},g_{2}\in\bm{g}$ and $\bm{g}\in\tilde{G}$.
-The fact that elements of $\tilde{G}^{(0)}$ are subgroups implies, for every $\bm{g}\in\tilde{G}$, there are unique elements $\bm{x},\bm{y}\in\tilde{G}^{(0)}$ such that $\bm{g}=g\cdot\bm{x}=\bm{y}\cdot g$ for any $g\in\bm{g}$.
-Let $s(\bm{g})=\bm{x}$ and $r(\bm{g})=\bm{y}$.
-Define the inverse $\bm{g}^{-1}\coloneq\{g^{-1}\colon g\in\bm{g}\}$ and the product $\bm{g}\bm{h}\coloneq\{gh\colon g\in\bm{g},h\in\bm{h}\}$ whenever $s(\bm{g})=r(\bm{h})$.
-Note also that
-$\bm{g}\bm{h}=\{gh\colon h\in\bm{h}\}\text{ for any }g\in\bm{g}.$ (1.1)
-With this structure, $\tilde{G}$ becomes a Hausdorff étale groupoid.
+Since elements of $\tilde{G}^{(0)}$ are subgroups of isotropy groups, we have that for any $\bm{g}\in\tilde{G}$, there are unique elements $X,Y\in\tilde{G}^{(0)}$ such that $\bm{g}=g\cdot X=Y\cdot g$ for all $g\in\bm{g}$.
+Let $s(\bm{g})\coloneq X$ and $r(\bm{g})\coloneq Y$.
+Explicitly, we have
+(1.1) $s(\bm{g})=g^{-1}\cdot\bm{g}\quad\quad\text{and}\quad\quad r(\bm{g})=\bm{g}\cdot g^{-1}\quad\quad\text{for any }g\in\bm{g}.$
+Define the inverse $\bm{g}^{-1}\coloneq\{g^{-1}\colon g\in\bm{g}\}$ and the product $\bm{g}\bm{h}\coloneq\{gh\colon g\in\bm{g},h\in\bm{h}\}$ whenever $s(\bm{g})=r(\bm{h})$. Note that
 […]
-$\mathfrak{i}(f)(\bm{g})=\sum_{g\in\bm{g}}f(g)$ (1.2)
-for all $\bm{g}\in\tilde{G}$ ([4, Lemma 3.2]).
-The embedding $Cc(G~)\mathfrak{i}\colon\mathscr$$\hookrightarrow C_{c}(\tilde{G})$ is a -homomorphism, and induces canonical embeddings $Cr(G~)\mathfrak{i}_{r}\coloneq C^{*}_{r}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}_{r}(\tilde{G})$ and $C(G~)\mathfrak{i}\coloneq C^{*}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}(\tilde{G})$ ([4, Lemma 3.8 & Corollary 6.8]).
+[…]
+(1.3) $\mathfrak{i}(f)(\bm{g})=\sum_{g\in\bm{g}}f(g)$
+for all $\bm{g}\in\tilde{G}$ ([5, Lemma 3.2]).
+The embedding $Cc(G~)\mathfrak{i}\colon\mathscr$$\hookrightarrow C_{c}(\tilde{G})$ is a -homomorphism, and induces canonical embeddings $Cr(G~)\mathfrak{i}_{r}\coloneq C^{*}_{r}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}_{r}(\tilde{G})$ and $C(G~)\mathfrak{i}\coloneq C^{*}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}(\tilde{G})$ ([5, Lemma 3.8 & Corollary 6.8]).
 Now assume that the groupoid $G$ is covered by countably many open bisections.
-We say a unit $x\in G^{(0)}$ is Hausdorff if $\pi^{-1}(x)=\{\iota(x)\}$, and we let $C\subseteq G^{(0)}$ denote the set of Hausdorff units.
-By [19, Lemma 7.15], $C$ is dense in $G^{(0)}$ (the Hausdorff points complement the so-called dangerous points of [19]).
-Define $\tilde{G}^{(0)}_{\rm{ess}}\coloneq\overline{\iota(C)}$.
-This is an invariant subset of the unit space $\tilde{G}^{(0)}$.
-We define the essential Hausdorff cover $\tilde{G}_{\rm{ess}}\coloneq\tilde{G}|_{\tilde{G}^{(0)}_{\rm{ess}}}=\{\bm{g}\in\tilde{G}\colon s(\bm{g})\in\tilde{G}^{(0)}_{\rm{ess}}\}$ ([4, Definition 4.13]).
+Observe that a unit $x\in G^{(0)}$ is Hausdorff if and only if the inclusion $\iota\colon G^{(0)}\to\tilde{G}^{(0)}$ is continuous at $x$, if and only if $\pi^{-1}\big(\{x\}\big)=\{\iota(x)\}$.
+Define $\tilde{G}^{(0)}_{\rm{ess}}\coloneq\overline{\iota(C)}$, where $C\subseteq G^{(0)}$ is the set of Hausdorff units.
+This is an invariant subset of the unit space $\tilde{G}^{(0)}$ of the Hausdorff cover.
+We define the essential Hausdorff cover $\tilde{G}_{\rm{ess}}\coloneq\tilde{G}|_{\tilde{G}^{(0)}_{\rm{ess}}}=\{\bm{g}\in\tilde{G}\colon s(\bm{g})\in\tilde{G}^{(0)}_{\rm{ess}}\}$ ([5, Definition 4.13]).
+Definition 1.3 ([19]).
+Let $G$ be an étale groupoid that is covered by countably many open bisections. For each $x\in G^{(0)}$ define $\mathcal{X}(x)\coloneq\pi^{-1}\big(\{x\}\big)\cap\tilde{G}^{(0)}_{{\rm{ess}}}$.
+By [19], $\mathcal{X}(x)$ is a conjugation-invariant collection of subgroups of $G^{x}_{x}$ that is closed in the Chabauty topology (see Section 3).
+A subgroup $X\subseteq G^{x}_{x}$ belongs to $\mathcal{X}(x)$ if and only if there exists a net $(x_{\alpha})\subset C$ of (Hausdorff) units whose set of limit points is precisely $X$, and such that every subnet of $(x_{\alpha})$ has its limit points contained in $X$.
+$\bullet$
 1.3. Isotropy fibres of ideals
@@
 The map $\eta_{x}$ extends to a completely positive contraction $C^{*}(G)\to C^{*}(G^{x}_{x})$, which we will also denote by $\eta_{x}$.
-This can be seen using induced representations of the isotropy group (see [7, Lemma 1.2] for example).
-In fact, any $C^{*}$-norm on $\mathscr{C}_{c}(G)$ dominating the reduced norm will define a $C^{*}$-norm on the group algebra $\mathbb{C}[G^{x}_{x}]$ dominating the reduced norm, and the restriction map $\mathscr{C}_{c}(G)\to\mathbb{C}[G^{x}_{x}]$ will always extend to a completely positive contraction between the associated $C^{*}$-algebras.
-See [8, Section 1.3] for a discussion in this direction.
-For a (closed) ideal $I\trianglelefteq C^{*}(G)$, the image $\eta_{x}(I)$ is also a (closed) ideal in $C^{*}(G^{x}_{x})$ ([8, Lemma 2.1]).
+This can be seen using induced representations of the isotropy group (see [8, Lemma 1.2] for example).
+For a (closed) ideal $I\trianglelefteq C^{*}(G)$, the image $\eta_{x}(I)$ is also a (closed) ideal in $C^{*}(G^{x}_{x})$ ([9, Lemma 2.1]).
 We write $I_{x}\coloneq\eta_{x}(I)$ and, following [9], call $I_{x}$ the isotropy fibre of $I$ at $x$.
-For $g\in G^{x}_{x}$, denote by $\delta_{g}\in\mathbb{C}[G^{x}_{x}]$ the function equal to one at $g$ and zero everywhere else.
-Each $g\in G$ induces an isomorphism of group algebras $\text{Ad}_{g}\colon\mathbb{C}\big[G^{s(g)}_{s(g)}\big]\to\mathbb{C}\big[G^{r(g)}_{r(g)}\big]$, $\delta_{h}\mapsto\delta_{ghg^{-1}}$ and this maps extends to an isomorphism of $C^{*}$-algebras $\Psi_{g}\colon C^{*}\big(G^{s(g)}_{s(g)}\big)\to C^{*}\big(G^{r(g)}_{r(g)}\big)$.
+For $g\in G^{x}_{x}$, denote by $\delta_{g}\in\mathbb{C}[G^{x}_{x}]$ the function equal to 1 at $g$ and zero everywhere else.
+Each $g\in G$ induces an isomorphism of group algebras $\text{Ad}_{g}\colon\mathbb{C}\big[G^{s(g)}_{s(g)}\big]\to\mathbb{C}\big[G^{r(g)}_{r(g)}\big]$; $\delta_{h}\mapsto\delta_{ghg^{-1}}$ and this map extends to an isomorphism of $C^{*}$-algebras $\Psi_{g}\colon C^{*}\big(G^{s(g)}_{s(g)}\big)\to C^{*}\big(G^{r(g)}_{r(g)}\big)$.
 A family $(K_{x})_{x\in G^{(0)}}$ of ideals satisfying $K_{x}\trianglelefteq C^{*}(G_{x}^{x})$ for each $x\in G^{(0)}$ is said to be invariant if $\Psi_{g}(K_{s(g)})=K_{r(g)}$ for all $g\in G$.
@@
 […]
-By [8, Corollary 2.9], $I(\mathcal{K})$ is an ideal in $C^{*}(G)$. We isolate the following observation from [8].
+By [9, Corollary 2.9], $I(\mathcal{K})$ is an ideal in $C^{*}(G)$. We isolate the following observation from [9].
 Lemma 1.4.
-Let $\mathcal{K}=(K_{x})_{x\in G^{(0)}}$ be an invariant family of ideals, and let $A\subseteq C^{*}(G)$ be a sub-$C^{*}$-algebra.
+Let $\mathcal{K}=(K_{x})_{x\in G^{(0)}}$ be an invariant family of ideals, and let $A\subseteq C^{*}(G)$ be a $C^{*}$-subalgebra.
 Then $A\subseteq I(\mathcal{K})$ if and only if $\eta_{x}(A)\subseteq K_{x}$ for all $x\in G^{(0)}$.
@@
 Fix $x\in G^{(0)}$.
-We have that $\eta_{x}(a)\in K_{x}$ for all positive elements $a\in A^{+}$.
+We have $\eta_{x}(a)\in K_{x}$ for all positive elements $a\in A^{+}$.
 Since $A$ is spanned by its positive elements, it follows that $\eta_{x}(a)\in K_{x}$ for all $a\in A$.
 Conversely, assume that $\eta_{x}(A)\subseteq K_{x}$ for all $x\in G^{(0)}$.
-For any $a\in A$ we have $\eta_{x}(a^{*}a)\in K_{x}$ for all $x\in G^{(0)}$, implying that $a\in I(\mathcal{K})$ as desired.
+For any $a\in A$ we have $\eta_{x}(a^{*}a)\in K_{x}$ for all $x\in G^{(0)}$, and this implies that $a\in I(\mathcal{K})$ as desired.
 ∎
@@
 By definition, $X$ is equal to a subgroup of the isotropy group $G^{x}_{x}$.
-It is easy to see that the isotropy group $\tilde{G}^{X}_{X}\subseteq\tilde{G}$ is equal to the quotient $N_{X}/X$, where $N_{X}=\{g\in G^{x}_{x}:gXg^{-1}=X\}$ is the normaliser of $X$ in $G^{x}_{x}$.
-Denote by $E_{N_{X}}\colon C^{*}(G^{x}_{x})\to C^{*}(N_{X})$ the canonical conditional expectation and $Q_{X}\colon C^{*}(N_{X})\to C^{*}(N_{X}/X)$ the -homomorphism induced by the quotient map $N_{X}\to N_{X}/X$.
+By (1.1) it is easily seen that the isotropy group $\tilde{G}^{X}_{X}\subseteq\tilde{G}$ is equal to the quotient $N_{X}/X$, where $N_{X}=\{g\in G^{x}_{x}\colon gXg^{-1}=X\}$ is the normaliser of $X$ in $G^{x}_{x}$.
+Denote by $E_{N_{X}}\colon C^{*}(G^{x}_{x})\to C^{*}(N_{X})$ the canonical conditional expectation and by $Q_{X}\colon C^{*}(N_{X})\to C^{*}(N_{X}/X)$ the -homomorphism induced by the quotient map $N_{X}\to N_{X}/X$.
 Lemma 2.3.
@@
 Proof.
-Equality is verified readily on the dense *-subalgebra $\mathscr{C}_{c}(G)$, and the lemma follows immediately. ∎
+Equality is readily verified on the dense -subalgebra $\mathscr{C}_{c}(G)$. The lemma then follows by continuity. ∎
 We can now describe the isotropy fibres of the ideals $\tilde{I}$ $C^{*}(\tilde{G}) in$ terms of $I$.
@@
 Since $\Psi_{\bm{g}}$ is an isomorphism, and since $(I_{x})_{x\in G^{(0)}}$ is an invariant family of ideals, it follows that
-$\Psi_{\bm{g}}(K_{s(\bm{g})})$ $=\overline{\Psi_{\bm{g}}\big(Q_{s(\bm{g})}\circ E_{N_{s(\bm{g})}}(I_{s(g_{0})})\big)}=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}\big(\Psi_{g_{0}}(I_{s(g_{0})})\big)}$
-$=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}(I_{r(g_{0})})}=K_{r(\bm{g})}.$
+$\Psi_{\bm{g}}\big(K_{s(\bm{g})}\big)$ $=\overline{\Psi_{\bm{g}}\Big(Q_{s(\bm{g})}\circ E_{N_{s(\bm{g})}}\big(I_{s(g_{0})}\big)\Big)}=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}\Big(\Psi_{g_{0}}\big(I_{s(g_{0})}\big)\Big)}$
+$=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}\big(I_{r(g_{0})}\big)}=K_{r(\bm{g})}.$
 Hence, $\mathcal{K}\coloneq(K_{X})_{X\in\tilde{G}^{(0)}}$ is an invariant family of ideals.
 By Lemma 2.3, we have $K_{X}=\overline{\eta_{X}\big(\mathfrak{i}(I)\big)}$, and therefore $K_{X}\subseteq\tilde{I}_{X}$ for all $X\in\tilde{G}^{(0)}$.
-By Lemma 1.3, this equality also implies that $\mathfrak{i}(I)\subseteq I(\mathcal{K})$.
-Hence $\tilde{I}\subseteq I(\mathcal{K})$, and another application of Lemma 1.3 gives $\tilde{I}_{X}\subseteq K_{X}$ for all $X\in\tilde{G}^{(0)}$, as desired.
+By Lemma 1.4, this equality also implies that $\mathfrak{i}(I)\subseteq I(\mathcal{K})$, and hence $\tilde{I}\subseteq I(\mathcal{K})$.
+Another application of Lemma 1.4 gives $\tilde{I}_{X}\subseteq K_{X}$ for all $X\in\tilde{G}^{(0)}$, as desired.
 ∎
 We now prove that ideals in $C^{*}(G)$ are determined by their isotropy fibres whenever $G$ is an amenable second-countable étale groupoid.
-This generalises [8, Theorem 2.10] to the setting of non-Hausdorff groupoids.
+This generalises [9, Theorem 2.10] to the setting of non-Hausdorff groupoids.
 Our proof utilises this theorem, which in turn uses known cases of the Effros-Hahn conjecture [20].
-Therefore it is necessary for us to only consider amenable second-countable groupoids.
-The theorem does not hold in the non-amenable setting - counterexamples are provided in [8, Examples 2.11 & 2.12].
+Therefore, it is necessary for us to only consider amenable second-countable groupoids.
+The theorem does not hold in the non-amenable setting – counterexamples are provided in [9, Examples 2.11 & 2.12].
 Theorem 2.5.
@@
 Proposition 2.4 implies that $\tilde{I}_{X}=\tilde{K}_{X}$ for all $X\in\tilde{G}^{(0)}$.
-Since $G$ is amenable and second-countable, so is $\tilde{G}$ by [4, Theorem 6.5].
-Moreover, the groupoid $\tilde{G}$ is Hausdorff, and therefore [8, Theorem 2.10] ensures that $\tilde{I}=\tilde{K}$.
+Since $G$ is amenable and second-countable, so is $\tilde{G}$ by [5, Theorem 6.5].
+Moreover, the groupoid $\tilde{G}$ is Hausdorff, and therefore [9, Theorem 2.10] ensures that $\tilde{I}=\tilde{K}$.
 Finally, Proposition 2.2 gives $I=\mathfrak{i}^{-1}(\tilde{I})=\mathfrak{i}^{-1}(\tilde{K})=K$.
 ∎
-2.1. Characterisation of density
-In this subsection, we assume that $G$ is an amenable second-countable (not necessarily Hausdorff) étale groupoid, and therefore $C^{*}(G)=C^{*}_{r}(G)$ by [4, Corollary 6.10].
+3. Characterisation of density
+In this section, we assume that $G$ is an amenable second-countable (not necessarily Hausdorff) étale groupoid, and therefore $C^{*}(G)=C^{*}_{r}(G)$ by [5, Corollary 6.10].
 We denote by $J$ the singular ideal in $C^{*}_{r}(G)$.
@@
 Proof.
-Let $I=\overline{J\cap\mathscr{C}_{c}(G)}$.
+Let $I\coloneq\overline{J\cap\mathscr{C}_{c}(G)}$.
 Since $\mathscr{C}_{c}(G)$ is dense in $C^{*}_{r}(G)$, $I$ is an ideal and $I\subseteq J$.
-By [16, Theorem 5.5 & Proposition 5.19], we have $\eta_{x}(J\cap\mathscr{C}_{c}(G))=J_{x}\cap\mathbb{C}[G^{x}_{x}]$.
+By [19, Theorem 5.5 & Proposition 5.19] we have $\eta_{x}(J\cap\mathscr{C}_{c}(G))=J_{x}\cap\mathbb{C}[G^{x}_{x}]$.
 Continuity of $\eta_{x}$ then implies $J_{x}\cap\mathbb{C}[G^{x}_{x}]\subseteq\eta_{x}(I)\subseteq\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}$.
 Since $\eta_{x}(I)=I_{x}$ is closed, it follows that $I_{x}=\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}$.
-Therefore, Theorem 2.5 shows that $I=J$ if and only if $\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}=J_{x}$ for all $x\in G^{(0)}$, as desired.
+Therefore, Theorem 2.5 shows that $I=J$ if and only if $\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}=J_{x}$ for all $x\in G^{(0)}$.
 ∎
@@
 The “only if” direction of Proposition 3.1 holds for all étale groupoids $G$ because $\eta_{x}(J\cap\mathscr{C}_{c}(G))\subseteq J_{x}\cap\mathbb{C}[G^{x}_{x}]$ holds trivially ($J_{x}$ here means the isotropy fibre from the reduced groupoid $C^{*}$-algebra).
-It can be shown using the calculation of isotropy fibres of the singular ideal in [16, Theorem 5.5] that the first example in [20] of a non-Hausdorff groupoid $G$ with $J\cap\mathscr{C}_{c}(G)$ not dense in $J$ has an isotropy fibre $J_{\epsilon}\neq\{0\}$ with $J_{\epsilon}\cap\mathbb{C}[\Gamma]=\{0\}$.
+It can be shown using the calculation of isotropy fibres of the singular ideal in [19, Theorem 5.5] that the first example in [25] of a non-Hausdorff groupoid $G$ with $J\cap\mathscr{C}_{c}(G)$ not dense in $J$ has an isotropy fibre $J_{\epsilon}\neq\{0\}$ with $J_{\epsilon}\cap\mathbb{C}[\Gamma]=\{0\}$.
 Therefore, this remark may be applied to this example to provide an alternate proof of non-density.
@@
 Let $\Gamma$ be an amenable discrete group.
-Let $\text{Sub}(\Gamma)$ denote the set of all subgroups in $\Gamma$.
+Let $\text{Sub}(\Gamma)$ denote the set of all subgroups of $\Gamma$.
 Consider the map
@@
 $X$ $\longmapsto\mathbbm{1}_{X}$
-where $\mathbbm{1}_{X}\colon\Gamma\to\{0,1\}$ is the characteristic function, and equip $\{0,1\}^{\Gamma}$ with the product topology.
+where $\mathbbm{1}_{X}\colon\Gamma\to\{0,1\}$ is the indicator function, and equip $\{0,1\}^{\Gamma}$ with the product topology.
 We equip $\text{Sub}(\Gamma)$ with the topology induced by this map, known as the Chabauty topology, making $\text{Sub}(\Gamma)$ a Stone space (this topology agrees with the Fell topology of [14]).
@@
 Let $\lambda_{\Gamma/X}\colon\Gamma\to B\big(\ell^{2}(\Gamma/X)\big)$ denote the associated quasi-regular representation as described in [4].
-Concretely, we have
-$\lambda_{\Gamma/X}(g)\delta_{hX}\coloneq\delta_{ghX}$ for $g\in\Gamma$ and $hX\in\Gamma/X$, where $\delta_{hX}\in\ell^{2}(\Gamma/X)$ denotes a standard basis vector.
+Concretely, we have $\lambda_{\Gamma/X}(g)\delta_{hX}\coloneq\delta_{ghX}$ for $g\in\Gamma$ and $hX\in\Gamma/X$, where $\delta_{hX}\in\ell^{2}(\Gamma/X)$ denotes a standard basis vector.
 We also write $\lambda_{\Gamma/X}$ for the associated -representation on $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$.
-Let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ be closed in the Chabauty topology and invariant under conjugation by elements of $\Gamma$.
-Following [16], we let the ideal $J_{\Gamma,\mathcal{X}}$ denote the $C^{*}$-kernel of $\lambda_{\Gamma/\mathcal{X}}\coloneq\oplus_{X\in\mathcal{X}}\lambda_{\Gamma/X}$ inside $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$.
-Concretely, we have $J_{\Gamma,\mathcal{X}}=\bigcap_{X\in\mathcal{X}}\text{ker}(\lambda_{\Gamma/X})$.
+Let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$.
+Following [19], we let the ideal $J_{\Gamma,\mathcal{X}}$ denote the intersection $\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/X})$ inside $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$.
+We will primarily be interested in the case where $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ is closed in the Chabauty topology and invariant under conjugation by elements of $\Gamma$.
 Definition 3.3.
@@
 Let $G$ be an amenable second-countable étale groupoid.
-For $x\in G^{(0)}$ write $\mathcal{X}(x)\coloneq\pi^{-1}(x)\cap\tilde{G}^{(0)}_{{\rm{ess}}}$, where $\pi\colon\tilde{G}^{(0)}\to G^{(0)}$ is the canonical surjection and $\tilde{G}^{(0)}_{\rm{ess}}$ is the unit space of the essential Hausdorff cover (see Subsection 1.2).
-The isotropy groups of $G$ are amenable, and in this case it was shown in [16, Definition 5.1 & Theorem 5.5] that $J_{x}=J_{G^{x}_{x},\mathcal{X}(x)}$ for every $x\in G^{(0)}$.
+For $x\in G^{(0)}$, recall from Definition 1.3 that $\mathcal{X}(x)\coloneq\pi^{-1}\big(\{x\}\big)\cap\tilde{G}^{(0)}_{{\rm{ess}}}$ is a closed conjugation-invariant collection of subgroups of the isotropy group $G^{x}_{x}$.
+The isotropy groups of $G$ are amenable, and in this case [19, Definition 5.1 & Theorem 5.5] showed that $J_{x}=J_{G^{x}_{x},\mathcal{X}(x)}$ for every $x\in G^{(0)}$.
 The following theorem is then immediate.
@@
 Remark 3.5.
-Since $J_{\Gamma,\mathcal{X}}=\{0\}$ whenever $\{e\}\in\mathcal{X}$, in the above characterisation one only needs to check whether $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ whenever $\{x\}\notin\mathcal{X}(x)$.
-These units $x$ are shown in [16, Corollary 5.6] to be precisely the extremely dangerous points described in [24, Proposition 1.12].
-Note also that the set of units $x\in G^{(0)}$ for which $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ is a $G$-invariant set.
+In Theorem 3.4 it suffices to check whether $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ for extremely dangerous points $x\in G^{(0)}$ (as described in [30, Proposition 1.12]).
+This is because $J_{\Gamma,\mathcal{X}}=\{0\}$ whenever the trivial group $\{e\}$ belongs to $\mathcal{X}$, and a unit $x\in G^{(0)}$ satisfies $\{x\}\notin\mathcal{X}(x)$ if and only if it is extremely dangerous (see [19, Corollary 5.6]).
 It is clear that finite groups satisfy the Density Property, and hence $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for any amenable second-countable étale groupoid with finite isotropy groups.
-Question 2.11.
+Question 3.6.
 Which discrete amenable groups satisfy the Density Property?
-Assume that there exists a (discrete) countable amenable group $\Gamma$ which fails to have the Density Property, and let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ be a closed conjugation-invariant set of subgroups for which $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is not dense in $J_{\Gamma,\mathcal{X}}$, i.e.
-$\mathcal{X}\notin\mathcal{D}_{\Gamma}$.
-Note that the trivial subgroup, $\{e\}$, is not in $\mathcal{X}$ (amenability of $\Gamma$ ensures that the quasi-regular representation $\lambda_{\Gamma/\{e\}}$ has trivial kernel).
-By [16, Section 6], there exists an amenable second-countable étale groupoid $G$ and a unit element $x_{0}\in G^{(0)}$ for which $G^{x_{0}}_{x_{0}}=\Gamma$ and $\mathcal{X}(x_{0})=\mathcal{X}$.
+Assume that there exists a (discrete) countable amenable group $\Gamma$ that fails to have the Density Property, and let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ be a closed conjugation-invariant set of subgroups for which $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is not dense in $J_{\Gamma,\mathcal{X}}$ (i.e.
+$\mathcal{X}\notin\mathcal{D}_{\Gamma}$).
+In particular, the trivial subgroup $\{e\}$ does not belong to $\mathcal{X}$.
+By [19, Section 6], there exists an amenable and second-countable étale groupoid $G$ and a unit element $x_{0}\in G^{(0)}$ for which $G^{x_{0}}_{x_{0}}=\Gamma$ and $\mathcal{X}(x_{0})=\mathcal{X}$.
 It follows from Theorem 3.4 that $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$.
@@
 Corollary 3.7.
-The intersection $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for all amenable second-countable étale groupoids if and only if all countable amenable groups satisfy the Density Property.
+The intersection $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for all second-countable amenable étale groupoids if and only if all (discrete) countable amenable groups satisfy the Density Property.
 4. Groups with the Density Property
 In this section, we study the set $\mathcal{D}_{\Gamma}$ for amenable discrete groups $\Gamma$, and show that certain classes of groups satisfy the Density Property (see Definition 3.3).
-Using Theorem 2.9, we are then able to deduce that $J\cap\mathscr{C}_{c}(G)$ is dense in the singular ideal $J$ for certain classes of groupoids.
-Throughout this section we work with the reduced group $C^{*}$-algebra $C^{*}_{r}(\Gamma)$ since this is canonically isomorphic to the full $C^{*}$-algebra.
-The set of subgroups $\text{Sub}(\Gamma)$ will always be equipped with the Chabauty topology, and given a closed and conjugation-invariant subset $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ we write $J_{\Gamma,\mathcal{X}}\coloneq\bigcap_{X\in\mathcal{X}}\text{ker}\lambda_{\Gamma/X}\trianglelefteq C^{*}_{r}(\Gamma)$ (see Subsection 2.1).
-Given a subgroup $\Lambda\subseteq\Gamma$, we write $E_{\Lambda}\colon C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Lambda)$ for the canonical conditional expectation and $I_{\Lambda}\colon C^{*}_{r}(\Lambda)\to C^{*}_{r}(\Gamma)$ for the canonical inclusion.
-If $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ is closed and conjugation-invariant, then $\Lambda\cap\mathcal{X}\coloneq\{\Lambda\cap X:X\in\mathcal{X}\}$ is a closed and conjugation-invariant set of subgroups of $\Lambda$ by [16, Proposition 7.4].
-We isolate the following observations from [16, Propositions 7.4 & 7.7].
+We do this in three parts.
+In Subsection 4.1 we simplify the conditions one needs to check for finding elements of the set $\mathcal{D}_{\Gamma}$.
+In Subsection 4.2 we prove some permanence results for the Density Property.
+We then use these observations in Subsection 4.3 to prove Theorem 4.16, which in particular states that all countable finite-by-nilpotent groups satisfy the Density Property.
+Combining with Theorem 3.4, we are able to deduce that $J\cap\mathscr{C}_{c}(G)$ is dense in the singular ideal $J$ for certain classes of groupoids.
+Throughout the section we work with the reduced group $C^{*}$-algebra $C^{*}_{r}(\Gamma)$ since this is canonically isomorphic to the full $C^{*}$-algebra.
+The set of subgroups $\text{Sub}(\Gamma)$ will always be equipped with the Chabauty topology, and given a closed and conjugation-invariant subset $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ we write $J_{\Gamma,\mathcal{X}}\coloneq\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/X})\trianglelefteq C^{*}_{r}(\Gamma)$ (see Section 3).
+Let us record three preliminary lemmas which will be useful throughout the subsections.
+For a subgroup $\Lambda\subseteq\Gamma$, we write $I_{\Lambda}\colon C^{*}_{r}(\Lambda)\to C^{*}_{r}(\Gamma)$ for the canonical inclusion and $E_{\Lambda}\colon C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Lambda)$ for the canonical conditional expectation.
+By [19, Proposition 7.4], $\Lambda\cap\mathcal{X}\coloneq\{\Lambda\cap X\colon X\in\mathcal{X}\}$ is a closed and conjugation-invariant set of subgroups of $\Lambda$.
+In the first lemma we isolate the following observations from [19, Propositions 7.4 & 7.7] which describe the images of $J_{\Lambda,\Lambda\cap\mathcal{X}}$ and $J_{\Gamma,\mathcal{X}}$ under the inclusion and conditional expectation maps.
 Lemma 4.1.
@@
   2. (ii)
-    Let $N$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. Then,
-    $E_{N}(J_{\Gamma,\mathcal{X}})=J_{N,\mathcal{X}}\quad\text{and}\quad E_{N}(J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma])=J_{N,\mathcal{X}}\cap\mathbb{C}[N].$
+    Let $N\subseteq\Gamma$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. Then,
+    $E_{N}\big(J_{\Gamma,\mathcal{X}}\big)=J_{N,\mathcal{X}}\quad\text{and}\quad E_{N}\big(J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]\big)=J_{N,\mathcal{X}}\cap\mathbb{C}[N].$
 Proof.
-For a proof of (i) see [16, Proposition 7.4]. For a proof of (ii), see the proof of [16, Proposition 7.7] (where $\Phi$ denotes $E_{N}$). ∎
-We begin our study into the Density Property by showing it is preserved under countable increasing unions.
+For a proof of (i) see [19, Proposition 7.4]. For a proof of (ii), see the proof of [19, Proposition 7.7] (where $\Phi$ denotes $E_{N}$). ∎
+The next two lemmas concern ideals and dense -subalgebras inside general $C^{*}$-algebras.
+Lemma 4.2.
+Let A be a $C^{*}$-algebra, $\mathcal{A}\subseteq A$ a dense -subalgebra, and $I\trianglelefteq A$ an ideal.
+Then $I\cap\mathcal{A}$ is dense in $I$ if and only if $I\cap\mathcal{A}$ contains a bounded net $(u_{\beta})$ such that $\lim_{\beta}\|a-au_{\beta}\|=0$ for all $a\in I$.
+Moreover, the $u_{\beta}$ can be chosen to be positive contractions.
+We will need the following consequence of this lemma: For any closed and conjugation-invariant set $\mathcal{X}\subseteq\text{Sub}(\Gamma)$, we have $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ contains a net of positive contractions converging strongly to the identity in $J_{\Gamma,\mathcal{X}}$.
+Proof.
+Assume that $I\cap\mathcal{A}$ is dense in $I$.
+Since $I$ is a $C^{*}$-algebra, it contains an approximate unit $(u_{\beta})$.
+Then $I\cap\mathcal{A}$ is a dense -subalgebra, so each $u_{\beta}$ can be approximated by a sequence of positive elements in the unit ball of $I\cap\mathcal{A}$.
+By a diagonalisation argument, $I\cap\mathcal{A}$ contains a net of positive contractions in $I\cap\mathcal{A}$ converging strongly to the identity in $I$.
+Conversely, assume that $I\cap\mathcal{A}$ contains a net of elements of norm at most $M>0$ converging strongly to the identity in $I$.
+Let $a\in I$ and $\varepsilon>0$.
+There exist $u\in I\cap\mathcal{A}$ and $b\in\mathcal{A}$ such that $\left\|u\right\|\leq M$, $\left\|a-au\right\|<\frac{\varepsilon}{2}$ and $\left\|a-b\right\|<\frac{\varepsilon}{2M}$.
+Then $bu\in I\cap\mathcal{A}$ and
+$\left\|a-bu\right\|\leq\left\|a-au\right\|+\left\|(a-b)u\right\|<\varepsilon.$
+Lemma 4.3.
+Let $A$ be a $C^{*}$-algebra, $\mathcal{A}\subseteq A$ a -subalgebra and $I\trianglelefteq A$ an ideal.
+Let $Q\colon A\to A/I$ denote the quotient map.
+If $Q(\mathcal{A})$ is dense in $Q(A)$ and $I\cap\mathcal{A}$ is dense in $I$, then $\mathcal{A}$ is dense in $A$.
+Proof.
+Let $a\in A$ and $\varepsilon>0$.
+Since $Q(\mathcal{A})$ is dense in $Q(A)$, we may choose $b\in\mathcal{A}$ such that $\|Q(a)-Q(b)\|<\frac{\varepsilon}{2}$.
+By definition of the quotient norm, there is $\tilde{c}\in I$ such that $\|a-b-\tilde{c}\|<\frac{\varepsilon}{2}$.
+Since $I\cap\mathcal{A}$ is dense in $I$, there is $c\in I\cap\mathcal{A}$ with $\|c-\tilde{c}\|<\frac{\varepsilon}{2}$.
+Hence, $d\coloneq b+c\in\mathcal{A}$ satisfies $\|a-d\|<\varepsilon$.
+∎
+$\bullet$
+4.1. Reduction Properties
+In this subsection we show that in order to determine whether $J_{\Gamma,\mathcal{X}}$ belongs to $\mathcal{D}_{\Gamma}$, it suffices to assume that $\Gamma$ is generated by the subgroups $X\in\mathcal{X}$ (Theorem 4.5), and that $\mathcal{X}$ is minimal in a certain sense (Theorem 4.7).
+The following lemma is well-known to experts in group theory – it allows us to approximate an arbitrary element $a\in C^{*}_{r}(\Gamma)$ by its restriction to finitely many cosets of a normal subgroup $N$.
+If $\Gamma$ is an amenable discrete group and $N$ a normal subgroup, define $a_{gN}\coloneq\delta_{g}\cdot E_{N}(\delta_{g^{-1}}\cdot a)\in C^{*}_{r}(\Gamma)$ for $a\in C^{*}_{r}(\Gamma)$ and $gN\in\Gamma/N$.
+Note that this is independent of the choice of representative $g$.
+Whenever $a_{gN}=0$ for all but finitely many cosets $gN$, the (finite) sum $\sum_{gN}a_{gN}=a$.
+Lemma 4.4.
+Let $\Gamma$ be an amenable discrete group and $N$ a normal subgroup.
+For any $a\in C^{*}_{r}(\Gamma)$, there is a net $(a_{i})$ of finite linear combinations of $a_{gN}$ (for $gN\in\Gamma/N$) that converges to $a$.
+Proof.
+By amenability of $\Gamma/N$ and [6, Theorem 2.6.8], there is a net $(\tilde{\varphi}_{i})$ of finitely supported positive definite functions $\Gamma/N\to\mathbb{C}$ converging pointwise to $1$.
+Let $q_{N}\colon\Gamma\to\Gamma/N$ denote the quotient map, and define the functions $\varphi_{i}\coloneq\tilde{\varphi}_{i}\circ q_{N}$.
+The $\varphi_{i}$ are positive definite, and converge pointwise to $1$ on $\Gamma$.
+Without loss of generality we may assume $\varphi_{i}(e)=1$ for all $i$.
+Then, by [6, Theorem 2.5.11(4)], the multiplier maps $m_{i}\colon C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Gamma)$ defined for $a=\sum_{g\in\Gamma}a_{g}\delta_{g}\in\mathbb{C}[\Gamma]$ as $m_{i}(a)=\sum_{g\in\Gamma}\varphi_{i}(g)a_{g}\delta_{g}$ are unital and completely positive, and are therefore contractions.
+For $a\in C^{*}_{r}(\Gamma)$, observe that $\big(m_{i}(a)\big)_{gN}=\tilde{\varphi}_{i}(gN)a_{gN}$ for each $gN\in\Gamma/N$, and hence $m_{i}(a)$ is a finite linear combination of $a_{gN}$.
+Since $(\varphi_{i})$ converges pointwise to $1$, $m_{i}(a)$ converges to $a$ whenever $a\in\mathbb{C}[\Gamma]$.
+Since the $m_{i}$ are contractions, $m_{i}(a)$ converges to $a$ for all $a\in C^{*}_{r}(\Gamma)$.
+Defining $a_{i}\coloneq m_{i}(a)$, the net $(a_{i})$ satisfies the conclusion of the lemma.
+∎
+We now prove our first reduction property.
+Theorem 4.5.
+Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups.
+Let $\Lambda\subseteq\Gamma$ be a subgroup such that $X\subseteq\Lambda$ for all $X\in\mathcal{X}$.
+Then, $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $\mathcal{X}\in\mathcal{D}_{\Lambda}$.
+Proof.
+Let $N\subseteq\Gamma$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. It suffices to prove the theorem in the case $\Lambda=N$, since $N$ is independent of the ambient group.
+If $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$, then continuity of $E_{N}$ and Lemma 4.1 (ii) imply that $J_{N,\mathcal{X}}\cap\mathbb{C}[N]=E_{N}(J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma])$ is dense in $J_{N,\mathcal{X}}=E_{N}(J_{\Gamma,\mathcal{X}})$.
+This proves the “only if” direction.
+Conversely, assume that $J_{N,\mathcal{X}}\cap\mathbb{C}[N]$ is dense in $J_{N,\mathcal{X}}$.
+By Lemma 4.2 there exists a net of contractions $(u_{\beta})\subseteq J_{N,\mathcal{X}}\cap\mathbb{C}[N]$ converging strongly to the identity in $J_{N,\mathcal{X}}$.
+Since $J_{\Gamma,\mathcal{X}}\cap I_{N}\big(C^{*}_{r}(N)\big)=I_{N}\big(J_{N,\mathcal{X}}\big)$ by Lemma 4.1 (i), we can regard $(u_{\beta})$ as a net in $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$.
+Let $a\in J_{\Gamma,\mathcal{X}}$, and let $(a_{i})$ be as in Lemma 4.4.
+By Lemma 4.1, $a_{gN}\in J_{\Gamma,\mathcal{X}}$ for all $gN\in\Gamma/N$.
+Since $a_{i}$ is a finite linear combination of $a_{gN}$, it follows that $a_{i}\in J_{\Gamma,\mathcal{X}}$ for all $i$.
+Moreover, since $\delta_{g^{-1}}\cdot a_{gN}\in I_{N}\big(J_{N,\mathcal{X}}\big)$, we have $\lim_{\beta}\|a_{gN}\cdot u_{\beta}-a_{gN}\|=0$ for all $gN\in\Gamma/N$ and therefore $\lim_{\beta}\|a_{i}\cdot u_{\beta}-a_{i}\|=0$ for all $i$.
+Since $(a_{i})$ converges to $a$ and $\|u_{\beta}\|\leq 1$ for all $\beta$, it follows from the above that $\lim_{\beta}\|a\cdot u_{\beta}-a\|=0$.
+Lemma 4.2 then implies that $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$, completing the proof.
+∎
+The following corollary will be used in the proofs of Theorems 4.9 and 4.16.
+Corollary 4.6.
+If $\Gamma$ is an amenable discrete group and $N$ a normal subgroup, then $\{N\}\in\mathcal{D}_{\Gamma}$.
+Proof.
+By Theorem 4.5, we can assume that $N=\Gamma$.
+Note that the ideal $J_{\Gamma,\mathcal{X}}=J_{\Gamma,\{\Gamma\}}$ is equal to the kernel of the trivial representation $\epsilon\colon C^{*}_{r}(\Gamma)\to\mathbb{C}$.
+Let $(F_{i})$ be a Følner net for $\Gamma$ and set $v_{i}=\frac{1}{|F_{i}|}\sum_{g\in F_{i}}\delta_{g}$.
+For any $g\in\Gamma$, we have $\lim_{i}\|\delta_{g}\cdot v_{i}-v_{i}\|=0$, hence (by linearity and density) $\lim_{i}\|a\cdot v_{i}-\epsilon(a)v_{i}\|=0$ for all $a\in C^{*}_{r}(\Gamma)$.
+By Kesten’s criteria ([6, Theorem 2.6.8(8)]) we have $\|v_{i}\|=1$ for all $i$, so the above limit implies $\lim_{i}\|a\cdot v_{i}\|=|\epsilon(a)|$ for all $a\in C^{*}_{r}(\Gamma)$.
+Set $u_{i}=\delta_{e}-v_{i}$ for all $i$.
+Then $(u_{i})$ is a net of elements in $J_{\Gamma,\{\Gamma\}}\cap\mathbb{C}[\Gamma]$ of norm at most 2 satisfying $\lim_{i}\|a\cdot u_{i}-a\|=\lim_{i}\|a\cdot v_{i}\|=|\epsilon(a)|=0$ for any $a\in J_{\Gamma,\{\Gamma\}}$.
+It follows from Lemma 4.2 that $J_{\Gamma,\{\Gamma\}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\{\Gamma\}}$, as desired.
+∎
+In order to determine whether $\mathcal{X}$ belongs to $\mathcal{D}_{\Gamma}$, Theorem 4.5 allows us to reduce the size of the ambient group $\Gamma$.
+The next result allows us to reduce the size of $\mathcal{X}$, and assume that it is minimal in a certain sense.
+For any discrete group $\Gamma$ and closed conjugation-invariant set $\mathcal{X}\subseteq\text{Sub}(\Gamma)$, let $\mathcal{X}_{\text{min}}\subseteq\mathcal{X}$ be the collection of subgroups $X\in\mathcal{X}$ for which $X^{\prime}\subseteq X$ and $X^{\prime}\in\mathcal{X}$ implies $X=X^{\prime}$.
+Then $\mathcal{X}_{\text{min}}$ is conjugation-invariant but not necessarily closed.
+Theorem 4.7.
+Let $\Gamma$ be a discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups.
+For every $X\in\mathcal{X}$, there is $Y\in\mathcal{X}_{\mathrm{min}}$ such that $Y\subseteq X$.
+Consequently, for $\Gamma$ amenable, we have $J_{\Gamma,\mathcal{X}}=J_{\Gamma,\overline{\mathcal{X}}_{\mathrm{min}}}$, so $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $\overline{\mathcal{X}}_{\mathrm{min}}\in\mathcal{D}_{\Gamma}$.
+Proof.
+Fix $X\in\mathcal{X}$.
+Let $\mathcal{Z}_{X}=\{Y\in\mathcal{X}\colon Y\subseteq X\}$ and (partially) order $\mathcal{Z}_{X}$ by inclusion.
+Let $\mathcal{C}\subseteq\mathcal{Z}_{X}$ be a chain in $\mathcal{Z}_{X}$, and view $\mathcal{C}$ as a net ordered by inclusion.
+Define $Y_{*}\coloneq\bigcap_{Y\in\mathcal{C}}Y$.
+The net $\mathcal{C}$ converges to $Y_{*}$ by definition of the Chabauty topology, so $Y_{*}$ belongs to $\mathcal{X}$ and hence to $\mathcal{Z}_{X}$.
+Therefore, every chain in $\mathcal{Z}_{X}$ has a lower bound.
+Zorn’s lemma implies that $\mathcal{Z}_{X}$ contains a minimal element $Y\in\mathcal{\mathcal{X}}$.
+This proves the first part of the theorem.
+Now assume that $\Gamma$ is amenable.
+If $X_{1}\subseteq X_{2}$ are subgroups of $\Gamma$, then $\ker(\lambda_{\Gamma/X_{1}})\subseteq\ker(\lambda_{\Gamma/X_{2}})$ (see [3, Appendix E and F]).
+Therefore, $J_{\Gamma,\mathcal{X}}=\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/X})=\bigcap_{X\in\mathcal{X}_{\text{min}}}\ker(\lambda_{\Gamma/X})$, and by [4, Proposition 3.3] we have $\bigcap_{X\in\mathcal{X}_{\text{min}}}\ker(\lambda_{\Gamma/X})=\bigcap_{X\in\overline{\mathcal{X}}_{\text{min}}}\ker(\lambda_{\Gamma/X})$.
+This completes the proof.
+∎
+$\bullet$
+4.2. Permanence Properties
+In this subsection we show that the set $\mathcal{D}_{\Gamma}$ is preserved under quotients (Theorem 4.9) and increasing unions (Proposition 4.10) of the ambient group.
+As a consequence, the Density Property is preserved under taking quotients and inductive limits.
+We also show that $\mathcal{D}_{\Gamma}$ contains any closed collection of normal subgroups (Corollary 4.13).
+Firstly, we prove that $\mathcal{D}_{\Gamma}$ is closed under finite unions.
+Proposition 4.8.
+Let $\Gamma$ be an amenable discrete group and suppose $\mathcal{X}_{1},\dots,\mathcal{X}_{n}\subseteq\mathrm{Sub}(\Gamma)$ are closed and conjugation-invariant sets of subgroups belonging to $\mathcal{D}_{\Gamma}$.
+Then the union $\mathcal{X}\coloneq\bigcup_{i=1}^{n}\mathcal{X}_{i}$ also belongs to $\mathcal{D}_{\Gamma}$.
+Proof.
+By definition, we have $J_{\Gamma,\mathcal{X}}=\bigcap^{n}_{i=1}J_{\Gamma,\mathcal{X}_{i}}$.
+For each $i=1,\dots,n$ there exists, by Lemma 4.2, a net $(u_{i,\beta})$ of positive contractions in $J_{\Gamma,\mathcal{X}_{i}}\cap\mathbb{C}[\Gamma]$ converging strongly to the identity in $J_{\Gamma,\mathcal{X}_{i}}$, and hence in $J_{\Gamma,\mathcal{X}}\subseteq J_{\Gamma,\mathcal{X}_{i}}$.
+Therefore, by a diagonalisation argument, the net $(u_{\beta_{1},...,\beta_{n}})\coloneq(u_{1,\beta_{1}}\cdot...\cdot u_{n,\beta_{n}})$ (equipped with the lexicographic partial ordering on $(\beta_{1},...,\beta_{n})$) has a subnet $(u_{\gamma})$ converging strongly to the identity in $J_{\Gamma,\mathcal{X}}$.
+The net $(u_{\gamma})$ lies in $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$, so $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Lemma 4.2.
+∎
+For an amenable discrete group $\Gamma$ and a normal subgroup $N$, let $Q_{N}:C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Gamma/N)$ be the -homomorphism induced by the quotient map $q_{N}:\Gamma\to\Gamma/N$.
+The map
+$q^{*}_{N}\colon\text{Sub}(\Gamma/N)$ $\rightarrow\text{Sub}(\Gamma);\text{ }Y\mapsto q_{N}^{-1}(Y),$
+is a continuous embedding satisfying $\text{Conj}_{g}\circ q^{*}_{N}=q^{*}_{N}\circ\text{Conj}_{q_{N}(g)}$ for all $g\in\Gamma$, where $\text{Conj}_{g}$ denotes conjugation by $g$.
+Therefore, if $\mathcal{Y}$ is a closed and conjugation-invariant set of subgroups of $\Gamma/N$, then $q^{*}_{N}(\mathcal{Y})$ is a closed and conjugation-invariant set of subgroups of $\Gamma$.
+Theorem 4.9.
+Let $\Gamma$ be an amenable discrete group and $N$ a normal subgroup. If $\mathcal{Y}\subseteq\mathrm{Sub}(\Gamma/N)$ is a closed and conjugation-invariant set of subgroups of $\Gamma/N$, then
+(4.1) $J_{\Gamma,q_{N}^{*}(\mathcal{Y})}=Q_{N}^{-1}(J_{\Gamma/N,\mathcal{Y}})\quad\text{and}\quad Q_{N}\big(J_{\Gamma,q_{N}^{*}(\mathcal{Y})}\cap\mathbb{C}[\Gamma]\big)=J_{\Gamma/N,\mathcal{Y}}\cap\mathbb{C}[\Gamma/N].$
+Moreover, $q^{*}_{N}(\mathcal{Y})\in\mathcal{D}_{\Gamma}$ if and only if $\mathcal{Y}\in\mathcal{D}_{\Gamma/N}$. Consequently, if $\Gamma$ satisfies the Density Property, then so does $\Gamma/N$.
+Proof.
+For $Y\in\mathcal{Y}$, it is easy to see that $\lambda_{\Gamma/q_{N}^{-1}(Y)}=\lambda_{(\Gamma/N)/Y}\circ Q_{N}$ and therefore $\ker\big(\lambda_{\Gamma/q_{N}^{-1}(Y)}\big)=Q_{N}^{-1}\big(\ker(\lambda_{(\Gamma/N)/Y})\big)$.
+Hence,
+$J_{\Gamma,q^{*}_{N}(\mathcal{Y})}=\bigcap_{Y\in\mathcal{Y}}\ker\big(\lambda_{\Gamma/q_{N}^{-1}(Y)}\big)=Q^{-1}_{N}\big(\bigcap_{Y\in\mathcal{Y}}\ker(\lambda_{(\Gamma/N)/Y})\big)=Q^{-1}_{N}(J_{\Gamma/N,\mathcal{Y}}).$
+This implies $Q_{N}\big(J_{\Gamma,q_{N}^{*}(\mathcal{Y})}\cap\mathbb{C}[\Gamma]\big)=J_{\Gamma/N,\mathcal{Y}}\cap\mathbb{C}[\Gamma/N]$ since $J_{\Gamma/N,\mathcal{Y}}\cap\mathbb{C}[\Gamma/N]\subseteq\mathbb{C}[\Gamma/N]=Q_{N}\big(\mathbb{C}[\Gamma]\big)$, proving (4.1).
+Next, suppose $q_{N}^{*}(\mathcal{Y})\in\mathcal{D}_{\Gamma}$, so that $J_{\Gamma,q^{*}_{N}(\mathcal{Y})}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,q^{*}_{N}(\mathcal{Y})}$.
+It follows from (4.1) and continuity of $Q_{N}$ that $J_{\Gamma/N,\mathcal{Y}}\cap\mathbb{C}[\Gamma/N]$ is dense in $J_{\Gamma/N,\mathcal{Y}}$.
+Therefore $\mathcal{Y}\in\mathcal{D}_{\Gamma/N}$, as required.
+[…]
+We now show the Density Property is preserved under taking countable increasing unions.
 Proposition 4.10.
@@
 Hence, $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]=\bigcup_{n}J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Lambda_{n}]$ is dense in $\bigcup_{n}J_{\Gamma,\mathcal{X}}\cap I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$.
-The countable increasing union $\bigcup_{n=1}^{\infty}I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$ is dense in $C^{*}_{r}(\Gamma)$, so $\bigcup_{n}J_{\Gamma,\mathcal{X}}\cap I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$ is dense in $J_{\Gamma,\mathcal{X}}$ by [11, Lemma III.4.1], completing the proof.
+The countable increasing union $\bigcup_{n=1}^{\infty}I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$ is dense in $C^{*}_{r}(\Gamma)$, so $\bigcup_{n}J_{\Gamma,\mathcal{X}}\cap I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$ is dense in $J_{\Gamma,\mathcal{X}}$ by [12, Lemma III.4.1], completing the proof.
 ∎
-Next, we prove a general lemma about ideals and dense -subalgebras.
+As an application of Theorem 4.9 and Proposition 4.10, the Density Property is preserved under inductive limits of groups.
+Corollary 4.11.
+If $(\varphi_{n}:\Gamma_{n}\to\Gamma_{n+1})_{n}$ is a sequence of homomorphisms between groups with the Density Property, then the inductive limit $\Gamma=\lim_{n}(\varphi_{n},\Gamma_{n})$ has the Density Property.
+Proof.
+Let $(\mu_{n}:\Gamma_{n}\to\Gamma)_{n}$ be the limit homomorphisms.
+By Theorem 4.9, $\mu_{n}(\Gamma_{n})$ has the Density Property for every $n\in\mathbb{N}$.
+Since the subgroups of this form increase to $\Gamma$, Proposition 4.10 implies $\Gamma$ has the Density Property.
+∎
+Let $\mathcal{X}$ be a closed and conjugation-invariant set of subgroups of $\Gamma$.
+Proposition 4.10 shows that we can approximate the ideal $J_{\Gamma,\mathcal{X}}$ through approximations of $\Gamma$.
+We finish the subsection by proving that we can also approximate $J_{\Gamma,\mathcal{X}}$ through approximations of $\mathcal{X}$.
+Given a finite subset $F\subseteq\Gamma$ and $X\in\mathcal{X}$, let $X_{F}\subseteq X$ be the subgroup generated by the set $\bigcup_{g\in\Gamma}X\cap gFg^{-1}$.
+Define $\mathcal{X}_{F}\coloneq\{X_{F}\colon X\in\mathcal{X}\}$, which is conjugation-invariant, and let $J_{\Gamma,\mathcal{X}_{F}}=\bigcap_{X_{F}\in\mathcal{X}_{F}}\ker(\lambda_{\Gamma/X_{F}})$.
+Notice that $J_{\Gamma,\mathcal{X}_{F}}\subseteq J_{\Gamma,\mathcal{X}_{K}}$ whenever $F\subseteq K$ (see [3, Appendix E and F]) and $J_{\Gamma,\mathcal{X}_{F}}=J_{\Gamma,\overline{\mathcal{X}_{F}}}$ (by [4, Proposition 3.3]).
+Proposition 4.12.
+Let $\Gamma$ be a countable discrete amenable group and $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups.
+Let $F_{n}\subseteq\Gamma$ be a sequence of finite sets that increase to $\Gamma$.
+Then, $J_{\Gamma,\mathcal{X}}=\overline{\bigcup_{n}J_{\Gamma,\mathcal{X}_{F_{n}}}}$.
+Proof.
+For each $n\in\mathbb{N}$, define $\mathcal{X}_{n}=\overline{\bigcup_{m\geq n}\mathcal{X}_{F_{m}}}$.
+We have $\mathcal{X}_{n+1}\subseteq\mathcal{X}_{n}$ for all $n\in\mathbb{N}$, so that $J_{\Gamma,\mathcal{X}_{n}}\subseteq J_{\Gamma,\mathcal{X}_{n+1}}$.
+Set $\mathcal{X}_{\infty}=\bigcap_{n}\mathcal{X}_{n}$.
+We first show $J_{\Gamma,\mathcal{X}_{\infty}}=\overline{\bigcup_{n}J_{\Gamma,\mathcal{X}_{n}}}$ using that $\mathcal{X}_{\infty}$ and $\mathcal{X}_{1},\mathcal{X}_{2},\dots$ are closed conjugation-invariant collections of subgroups.
+For $n\in\mathbb{N}\cup\{\infty\}$, let $\Gamma\cdot\mathcal{X}_{n}$ be the étale groupoid of cosets of $\mathcal{X}_{n}$ (as described in [19, Section 5.1]) and denote by $q_{n}:C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Gamma\cdot\mathcal{X}_{n})$ the canonical -homomorphism $\delta_{g}\mapsto\delta_{g\mathcal{X}_{n}}$, $g\in\Gamma$.
+As noted in [19], we have $\ker(q_{n})=J_{\Gamma,\mathcal{X}_{n}}.$ For $n_{1},n_{2}\in\mathbb{N}\cup\{\infty\}$ such that $n_{1}\leq n_{2}$, $\mathcal{X}_{n_{2}}$ is a closed $\Gamma\cdot\mathcal{X}_{n_{1}}$ invariant subset of $(\Gamma\cdot\mathcal{X}_{n_{1}})^{0}=\mathcal{X}_{n_{1}}$, so the restriction maps $q_{n_{2},n_{1}}:C_{r}^{*}(\Gamma\cdot\mathcal{X}_{n_{1}})\to C_{r}^{*}(\Gamma\cdot\mathcal{X}_{n_{2}})$ are -homomorphisms.
+Moreover, since $\mathcal{X}_{n}$ decrease to $\mathcal{X}_{\infty}$, the maps $(q_{\infty,n})_{n\in\mathbb{N}}$ form an inductive limit for the inductive sequence $(q_{n+1,n})_{n\in\mathbb{N}}$.
+Therefore, by [32, Proposition 6.2.4] and the fact that $q_{m}=q_{m,n}\circ q_{n}$ for all $m,n\in\mathbb{N}\cup\{\infty\}$ with $m\geq n$, we have $a\in J_{\Gamma,\mathcal{X}_{\infty}}$ if and only if $\lim_{n\to\infty}\|q_{n}(a)\|=0$.
+Hence, for every $\varepsilon>0$, we can choose $n\in\mathbb{N}$ such that $\|q_{n}(a)\|\leq\frac{\varepsilon}{2}$.
+Also, since $\ker(q_{n})=J_{\Gamma,\mathcal{X}_{n}}$ (and since $q_{n}$ is a quotient map) there exists $b\in J_{\Gamma,\mathcal{X}_{n}}$ such that $\big|\|a-b\|-\|q_{n}(a)\|\big|\leq\frac{\varepsilon}{2}$.
+Therefore, $\|a-b\|\leq\varepsilon$.
+This proves $J_{\Gamma,\mathcal{X}_{\infty}}=\overline{\bigcup_{n}J_{\Gamma,\mathcal{X}_{n}}}$.
+Next, let us show that $J_{\Gamma,\mathcal{X}_{n}}=J_{\Gamma,\mathcal{X}_{F_{n}}}$ for all $n\in\mathbb{N}$.
+By [4, Proposition 3.3], we have $J_{\Gamma,\mathcal{X}_{n}}=J_{\Gamma,\bigcup_{m\geq n}\mathcal{X}_{F_{m}}}$.
+Since $X_{F_{n}}\subseteq X_{F_{m}}$ for all $X\in\mathcal{X}$ and $m\geq n$, it follows from [3, Appendix E and F] that $J_{\Gamma,\bigcup_{m\geq n}\mathcal{X}_{F_{m}}}=J_{\Gamma,\mathcal{X}_{F_{n}}}$.
+Finally, we prove that $\mathcal{X}_{\infty}=\mathcal{X}$.
+For fixed $X\in\mathcal{X}$, the sequence $(X_{F_{n}})$ converges to $X$ in the Chabauty topology, so that $\mathcal{X}\subseteq\mathcal{X}_{\infty}$.
+For the other inclusion, let $Y\in\mathcal{X}_{\infty}$.
+Since $\mathcal{X}_{\infty}=\bigcap_{n}\overline{\bigcup_{m\geq n}\mathcal{X}_{F_{m}}}$, there are sequences $X_{k}\in\mathcal{X}$ and $n_{k}\to\infty$ such that the sequence $((X_{k})_{F_{n_{k}}})_{k}$ converges to $Y$.
+By passing to a subsequence if necessary, we may assume that $(X_{k})_{k}$ converges to some $X\in\mathcal{X}$.
+Since $(X_{k})_{F_{n_{k}}}\subseteq X_{k}$ for all $k$, it is clear that $Y\subseteq X$.
+On the other hand, for every $x\in X$ we have $x\in X_{k}$ eventually.
+Since the $F_{n_{k}}$ are increasing to $\Gamma$, we have $x\in F_{n_{k}}$ eventually.
+Combining these observations gives that $x\in(X_{k})_{F_{n_{k}}}$ eventually, so $x\in Y$ and hence $X\subseteq Y$.
+This completes the proof of the proposition.
+∎
+We have the following immediate corollary.
+Corollary 4.13.
+Let $\Gamma$ be a countable discrete amenable group and $\mathcal{X}$ a closed collection of normal subgroups. Then, $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
+Proof.
+Let $F$ be a finite subset of $\Gamma$.
+Since $X\in\mathcal{X}$ is normal, we have $\bigcup_{g\in\Gamma}X\cap gFg^{-1}=\bigcup_{g\in\Gamma}g(X\cap F)g^{-1}$, so that there are only finitely many distinct $X_{F}$, each being normal in $\Gamma$.
+By Corollary 4.6 and Proposition 4.8, we have $\mathcal{X}_{F}\in\mathcal{D}_{\Gamma}$.
+Now, choosing a sequence ($F_{n}$) of finite subsets increasing to $\Gamma$, we have $J_{\Gamma,\mathcal{X}}=\overline{\bigcup_{n}J_{\Gamma,\mathcal{X}_{F_{n}}}}$ by Proposition 4.12.
+Since $\mathcal{X}_{F_{n}}\in\mathcal{D}_{\Gamma}$ for all $n\in\mathbb{N}$, it follows that $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
+∎
+$\bullet$
+4.3. Applications
+We will now show that finite-by-nilpotent groups satisfy the Density Property.
+We also show that any closed conjugation-invariant collection $\mathcal{X}$ of finite (or co-finite) subgroups of an amenable group $\Gamma$ belongs to $\mathcal{D}_{\Gamma}$.
+Let us establish a key lemma that unifies the cases.
 Lemma 4.14.
-Let A be a $C^{*}$-algebra, $\mathcal{A}\subseteq A$ a dense -subalgebra, and $I\trianglelefteq A$ an ideal.
-Then $I\cap\mathcal{A}$ is dense in $I$ if and only if $I\cap\mathcal{A}$ contains a net $(u_{\beta})$ of contractions such that $\lim_{\beta}\|a-au_{\beta}\|=0$ for all $a\in I$.
-Moreover, $u_{\beta}$ can be chosen to be positive.
-We will need the following consequence of this lemma: For any closed and conjugation-invariant set $\mathcal{X}\subseteq\text{Sub}(\Gamma)$, we have $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ contains a net of positive contractions converging strongly to the identity in $J_{\Gamma,\mathcal{X}}$.
-Proof.
-Assume that $I\cap\mathcal{A}$ is dense in $I$.
-Since $I$ is a $C^{*}$-algebra, it contains an approximate unit $(u_{\beta})$.
-Then $I\cap\mathcal{A}$ is a dense -subalgebra, so each $u_{\beta}$ can be approximated by a sequence of positive elements in the unit ball of $I\cap\mathcal{A}$.
-By a diagonalisation argument, $I\cap\mathcal{A}$ contains a net of positive elements $I$ converging strongly to the identity in $I$.
-Conversely, assume that $I\cap\mathcal{A}$ contains a net of contractions converging strongly to the identity in $I$.
-Let $a\in I$ and $\varepsilon>0$.
-There exist $u\in I\cap\mathcal{A}$ and $b\in\mathcal{A}$ such that $\left\|u\right\|\leq 1$, $\left\|a-au\right\|<\frac{\varepsilon}{2}$ and $\left\|a-b\right\|<\frac{\varepsilon}{2}$.
-Then $bu\in I\cap\mathcal{A}$ and
-$\left\|a-bu\right\|\leq\left\|a-au\right\|+\left\|(a-b)u\right\|<\varepsilon.$
-We now prove our first application of this lemma.
-Proposition 3.4.
-Let $\Gamma$ be an amenable discrete group and suppose $\mathcal{X}_{1},\dots,\mathcal{X}_{n}\subseteq\mathrm{Sub}(\Gamma)$ are closed and conjugation-invariant sets of subgroups satisfying $\mathcal{X}_{i}\in\mathcal{D}_{\Gamma}$ for all $i=1,\dots,n$.
-Then, $\mathcal{X}\coloneq\bigcup_{i=1}^{n}\mathcal{X}_{i}$ satisfies $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
-Proof.
-By definition, we have $J_{\Gamma,\mathcal{X}}=\bigcap^{n}_{i=1}J_{\Gamma,\mathcal{X}_{i}}$.
-For each $i=1,\dots,n$ there exists, by Lemma 3.3, a net $(u_{i,\beta})$ of positive contractions in $J_{\Gamma,\mathcal{X}_{i}}\cap\mathbb{C}[\Gamma]$ converging strongly to the identity in $J_{\Gamma,\mathcal{X}_{i}}$ and hence in $J_{\Gamma,\mathcal{X}}\subseteq J_{\Gamma,\mathcal{X}_{i}}$.
-Therefore, by a diagonalisation argument, the net $(u_{\beta_{1},\beta_{2}..,\beta_{n}}):=(u_{1,\beta_{1}}\cdot u_{2,\beta_{n}}\cdot...\cdot u_{n,\beta_{n}})$ (equipped with the lexicographic partial ordering on $(\beta_{1},...,\beta_{n})$) has a subnet $(u_{\gamma})$ converging strongly to the identity in $J_{\Gamma,\mathcal{X}}$.
-The net $(u_{\gamma})$ lies in $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$, so $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Lemma 3.3.
-∎
-Next, we show the Density Property is preserved under taking quotients of groups by normal subgroups.
-For an amenable discrete group $\Gamma$ and a normal subgroup $N$, let $Q_{N}:C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Gamma/N)$ be the *-homomorphism induced by the quotient map $q_{N}:\Gamma\to\Gamma/N$.
-The map
-$q^{*}_{N}:\text{Sub}(\Gamma/N)$ $\rightarrow\text{Sub}(\Gamma),\text{ }X\mapsto q_{N}^{-1}(X),$
-is a continuous injection satisfying $\text{Conj}_{g}\circ q^{*}_{N}=q^{*}_{N}\circ\text{Conj}_{q_{N}(g)}$ for all $g\in\Gamma$, where $\text{Conj}_{g}$ denotes conjugation by $g$.
-Therefore, if $\mathcal{X}$ is a closed and conjugation-invariant set of subgroups of $\Gamma/N$, then $q^{*}_{N}(\mathcal{X})$ is a closed and conjugation-invariant set of subgroups of $\Gamma$.
-Theorem 3.5.
-Suppose $\Gamma$ is an amenable discrete group and $N$ a normal subgroup. If $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma/N)$ is a closed and conjugation-invariant set of subgroups of $\Gamma/N$, then
-$J_{\Gamma,q_{N}^{*}(\mathcal{X})}=Q_{N}^{-1}(J_{\Gamma/N,\mathcal{X}}).$
-Moreover, if $q^{*}_{N}(\mathcal{X})\in\mathcal{D}_{\Gamma}$, then $\mathcal{X}\in\mathcal{D}_{\Gamma/N}$. Consequently, if $\Gamma$ satisfies the Density Property, then so does $\Gamma/N$.
-Proof.
-For $X\in\mathcal{X}$, it is easy to see that $\lambda_{\Gamma/q_{N}^{-1}(X)}=\lambda_{(\Gamma/N)/\mathcal{X}}\circ Q_{N}$ and therefore $\ker(\lambda_{\Gamma/q_{N}^{-1}(X)})=Q_{N}^{-1}(\ker(\lambda_{(\Gamma/N)/\mathcal{X}}))$.
-Hence,
-$J_{\Gamma,q^{*}_{N}(\mathcal{X})}=\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/q_{N}^{-1}(X)})=Q^{-1}_{N}\big(\bigcap_{X\in\mathcal{X}}\ker(\lambda_{(\Gamma/N)/X})\big)=Q^{-1}_{N}(J_{\Gamma/N,\mathcal{X}}).$
-Now, suppose $q_{N}^{*}(\mathcal{X})\in\mathcal{D}_{\Gamma}$.
-By Lemma 3.3, we may choose a net of positive contractions $(u_{\lambda})\subseteq J_{\Gamma,q^{*}_{N}(\mathcal{X})}\cap\mathbb{C}[\Gamma]$ converging strongly to the identity in $J_{\Gamma,q^{*}_{N}(\mathcal{X})}$.
-Since $J_{\Gamma,q_{N}^{*}(\mathcal{X})}=Q_{N}^{-1}(J_{\Gamma/N,\mathcal{X}})$, $Q_{N}$ surjects $J_{\Gamma,q_{N}^{*}(\mathcal{X})}$ onto $J_{\Gamma/N,\mathcal{X}}$ and therefore $(Q_{N}(u_{\lambda}))\subseteq J_{\Gamma/N,\mathcal{X}}\cap\mathbb{C}[\Gamma/N]$ is a net of positive contractions converging strongly to the identity in $J_{\Gamma/N,\mathcal{X}}$.
-By Lemma 3.3, it follows that $\mathcal{X}\in\mathcal{D}_{\Gamma/N}$.
-∎
-We will now begin to prove the following result.
-Theorem 3.6.
-Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups.
-Let $\Lambda\subseteq\Gamma$ be a subgroup such that $X\subseteq\Lambda$ for all $X\in\mathcal{X}$.
-Then, $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $\mathcal{X}\in\mathcal{D}_{\Lambda}$.
-The following lemma is probably well-known to experts in group theory - it allows us to approximate an element $a\in C^{*}_{r}(\Gamma)$ by its restriction to finitely many cosets of a normal subgroup $N$.
-If $\Gamma$ is an amenable discrete group and $N$ a normal subgroup, for $gN\in\Gamma/N$, and $a\in C^{*}_{r}(\Gamma)$, define $a_{gN}\coloneq\delta_{g}\cdot E_{N}(\delta_{g^{-1}}\cdot a)\in C^{*}_{r}(\Gamma)$.
-Whenever $a_{gN}=0$ for all but finitely many cosets $gN$, the (finite) sum $\sum_{gN}a_{gN}=a$.
-Lemma 3.7.
-Let $\Gamma$ be an amenable discrete group and $N$ a normal subgroup.
-Choose a Følner net $(F_{i})$ in $\Gamma/N$ and define $\varphi_{i}:\Gamma/N\to\mathbb{C}$ via $gN\mapsto\frac{|gF_{i}\cap F_{i}|}{|F_{i}|}$.
-For any $a\in C^{*}_{r}(\Gamma)$ define
-$a_{i}\coloneq\sum_{gN\in\Gamma/N}\varphi_{i}(gN)a_{gN}.$
-Then, $\|a_{i}\|\leq\|a\|$ for all $i$ and the net $(a_{i})$ converges to $a$.
-Proof.
-[…]
-The fact that $(F_{i})$ is a Følner net implies $(\tilde{\varphi}_{i}(g))$ converges to $1$ for all $g\in\Gamma$.
-Therefore, $(m_{i}(a))$ converges to $a$ whenever $a\in\mathbb{C}[\Gamma]$, and hence for all $a\in C^{*}_{r}(\Gamma)$ (by boundedness of $m_{i})$.
-∎
-Proof of Theorem 3.6.
-Let $N$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. It suffices to prove the theorem in the case $\Lambda=N$, since $N$ is independent of the ambient group.
-If $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$, then continuity of $E_{N}$ and Lemma 3.1 (ii) imply that $J_{N,\mathcal{X}}\cap\mathbb{C}[N]=E_{N}(J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma])$ is dense in $J_{N,\mathcal{X}}=E_{N}(J_{\Gamma,\mathcal{X}})$.
-This proves the “only if” direction.
-Conversely, assume that $J_{N,\mathcal{X}}\cap\mathbb{C}[N]$ is dense in $J_{N,\mathcal{X}}$.
-By Lemma 3.3 there exists a net of contractions $(u_{beta})\subseteq J_{N,\mathcal{X}}\cap\mathbb{C}[N]$ converging strongly to the identity in $J_{N,\mathcal{X}}$.
-Since $J_{\Gamma,\mathcal{X}}\cap I_{N}\big(C^{*}_{r}(N)\big)=I_{N}\big(J_{N,\mathcal{X}}\big)$ by Lemma 3.1 (i), we can regard $(u_{\beta})$ as a net in $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$.
-Let $a\in J_{\Gamma,\mathcal{X}}$ and fix a Følner net $(F_{i})$ for $\Gamma/N$, with $(a_{i})$ as in Lemma 3.7.
-By Lemma 3.1, $a_{gN}\in J_{\Gamma,\mathcal{X}}$ for all $gN\in\Gamma/N$.
-Since $a_{i}$ is a finite linear combination of the $a_{gN}$, it follows that $a_{i}\in J_{\Gamma,\mathcal{X}}$ for all $i$.
-Moreover, since $\delta_{g^{-1}}\cdot a_{gN}\in J_{N,\mathcal{X}}$, we have $\lim_{\beta}\|a_{gN}\cdot u_{\beta}-a_{gN}\|=0$ for all $gN\in\Gamma/N$ and therefore $\lim_{\beta}\|a_{i}\cdot u_{\beta}-a_{i}\|=0$ for all $i$.
-Since $(a_{i})$ converges to $a$, $\|a_{i}\|\leq\|a\|$ and $\|u_{\beta}\|\leq 1$ for all $i,\beta$, it follows from the above that $\lim_{\beta}\|a\cdot u_{\beta}-a\|=0$.
-Lemma 3.3 then implies $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$, completing the proof.
-∎
-Corollary 3.8.
-If $\Gamma$ is an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ is a finite collection of normal subgroups, then $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
-Proof.
-By Proposition 3.4, we only need to establish the case where $\mathcal{X}=\{N\}$.
-By Theorem 3.6, it then suffices to establish the case where $\Gamma=N$.
-In this case, $J_{\Gamma,\{\Gamma\}}$ is the kernel of the trivial representation $\epsilon:C^{*}_{r}(\Gamma)\to\mathbb{C}$.
-Let $(F_{i})$ be a Følner net for $\Gamma$ and set $v_{i}=\frac{1}{|F_{i}|}\sum_{g\in F_{i}}\delta_{g}$.
-For any $g\in\Gamma$, we have $\lim_{i}\|\delta_{g}\cdot v_{i}-v_{i}\|=0$, hence (by linearity and density) $\lim_{i}\|a\cdot v_{i}-\epsilon(a)v_{i}\|=0$ for all $a\in C^{*}_{r}(\Gamma)$.
-By Kesten’s criteria ([5, Theorem 2.6.8(8)]) we have $\|v_{i}\|=1$ for all $i$, so the above limit implies $\lim_{i}\|a\cdot v_{i}\|=|\epsilon(a)|$ for all $a\in C^{*}_{r}(\Gamma)$.
-Set $u_{i}=\frac{1}{2}(\delta_{e}-v_{i})$ for all $i$.
-Then $(u_{i})$ is a net of contractions in $J_{\Gamma,\{\Gamma\}}\cap\mathbb{C}[\Gamma]$, and $\lim_{i}\|a\cdot u_{i}-a\|=\lim_{i}\frac{1}{2}\|a\cdot v_{i}\|=\frac{1}{2}|\epsilon(a)|=0$ for any $a\in J_{\Gamma,\{\Gamma\}}$.
-It follows from Lemma 3.3 that $J_{\Gamma,\{\Gamma\}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\{\Gamma\}}$, as desired.
-∎
-Theorem 3.6 shows that in order to determine whether a closed and conjugation-invariant set $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ belongs to $\mathcal{D}_{\Gamma}$, it suffices to assume that $\Gamma$ is generated by the subgroups $X\in\mathcal{X}$.
-The next result shows that we can also assume $\mathcal{X}$ is minimal in a certain sense.
-For any discrete group $\Gamma$ and $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ closed conjugation-invariant, let $\mathcal{X}_{\text{min}}$ be the collection of $X\in\mathcal{X}$ such that $X^{\prime}\subseteq X$ and $X^{\prime}\in\mathcal{X}$ implies $X=X^{\prime}$.
-Then, $\mathcal{X}_{\text{min}}$ is conjugation-invariant but not necessarily closed.
-Theorem 3.9.
 Let $\Gamma$ be a discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups.
-Then, for every $X\in\mathcal{X}$, there is $Y\in\mathcal{X}_{\mathrm{min}}$ such that $Y\subseteq X$.
-Consequently, for $\Gamma$ amenable, we have $J_{\Gamma,\mathcal{X}}=J_{\Gamma,\overline{\mathcal{X}_{\mathrm{min}}}}$ and $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $\overline{\mathcal{X}_{\mathrm{min}}}\in\mathcal{D}_{\Gamma}$.
-Proof.
-Fix $X\in\mathcal{X}$.
-Let $\mathcal{Z}_{X}=\{Y\in\mathcal{X}:Y\subseteq X\}$ and (partial) order $\mathcal{Z}_{X}$ by inclusion.
-Let $\mathcal{C}\subseteq\mathcal{Z}_{X}$ be a chain in $\mathcal{Z}_{X}$.
-Then, $\mathcal{C}$ is directed, so we can view it as a net, ordered by inclusion.
-By compactness of $\mathcal{X}$ the net $\mathcal{C}$ has a convergent subnet $(Y_{\beta})$, and let $Y_{*}$ be its limit point.
-As $(Y_{\beta})$ is a subnet, for every $Y^{\prime}\in\mathcal{C}$, there is $\beta_{0}$ such that $Y_{\beta_{0}}\subseteq Y^{\prime}$.
-Therefore, for all $\beta\geq\beta_{0}$, we have $Y_{\beta}\subseteq Y^{\prime}$ and hence $Y_{*}\subseteq Y^{\prime}$.
-Hence, $Y_{*}\subseteq Y^{\prime}\subseteq X$ for all $Y^{\prime}\in\mathcal{C}$.
-Therefore, every chain in $\mathcal{Z}_{X}$ has a minimal element.
-Zorn’s lemma implies $\mathcal{Z}_{X}$ contains a minimal element $Y\in\mathcal{\mathcal{X}}$.
-If $X_{1}\subseteq X_{2}$ are subgroups of $\Gamma$, then $\ker(\lambda_{\Gamma/X_{1}})\subseteq\ker(\lambda_{\Gamma/X_{2}})$ (see [2, Appendix E and F]).
-Therefore, $J_{\Gamma,\mathcal{X}}=\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/X})=\bigcap_{X\in\mathcal{X}_{\text{min}}}\ker(\lambda_{\Gamma/X})$.
-By [3, Proposition 3.3], we have
-$\bigcap_{X\in\mathcal{X}_{\text{min}}}\ker(\lambda_{\Gamma/X})=\bigcap_{X\in\overline{\mathcal{X}_{\text{min}}}}\ker(\lambda_{\Gamma/X})$, which completes the proof. ∎
-We will now show that abelian groups satisfy the Density Property, and that collections of finite subgroups always belong to $\mathcal{D}_{\Gamma}$.
-First, we establish a key lemma that unifies the two cases.
-Lemma 3.10.
-Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups.
 If every $X\in\mathcal{X}_{\mathrm{min}}$ is finitely generated, then $\mathcal{X}_{\mathrm{min}}$ is finite.
 Proof.
-For each $X\in\mathcal{X}_{\text{min}}$, the set $\mathcal{Z}(X)=\{X^{\prime}\in\mathcal{X}:X\subseteq X^{\prime}\}=\bigcap^{n}_{i=1}\{X^{\prime}\in\mathcal{X}:x_{i}\in X^{\prime}\}$, where $\{x_{1},...,x_{n}\}$ is a finite generating set for $X$.
-Therefore, $\mathcal{Z}(X)$ is open in $\mathcal{X}$ with respect to the Chabauty topology.
-Since $\mathcal{Z}(X)\cap\mathcal{X}_{\text{min}}=\{X\}$ for all $X\in\mathcal{X}_{\text{min}}$, the set $\mathcal{X}_{\text{min}}$ is discrete in the compact space $\mathcal{X}$, and is therefore finite.
+For each $X\in\mathcal{X}_{\text{min}}$, note that $\mathcal{Z}(X)\coloneq\{X^{\prime}\in\mathcal{X}\colon X\subseteq X^{\prime}\}=\bigcap^{n}_{i=1}\{X^{\prime}\in\mathcal{X}\colon x_{i}\in X^{\prime}\}$, where $\{x_{1},...,x_{n}\}$ is a finite generating set for $X$.
+Therefore, $\mathcal{Z}(X)$ is open in $\mathcal{X}$ with respect to the Chabauty topology, and hence $\{\mathcal{Z}(X)\colon X\in\mathcal{X}_{\text{min}}\}$ forms an open cover of $\mathcal{X}$ by Theorem 4.7.
+Since $\mathcal{Z}(X)\cap\mathcal{X}_{\text{min}}=\{X\}$ for each $X\in\mathcal{X}_{\text{min}}$, the cover $\{\mathcal{Z}(X)\colon X\in\mathcal{X}_{\text{min}}\}$ admits no proper subcover, so compactness of $\mathcal{X}$ ensures that $\mathcal{X}_{\text{min}}$ is finite.
 ∎
-Now, we see our result for abelian groups and collections of finite subgroups as a simple consequence of Theorems 3.6, Theorem 3.9, Corollary 3.8 and Lemma 3.10.
+Recall that a group $\Gamma$ is nilpotent if there is a finite sequence $\{e\}=\Lambda_{0}\trianglelefteq\Lambda_{1}\trianglelefteq\cdots\trianglelefteq\Lambda_{n}=\Gamma$ of subgroups that are normal in $\Gamma$ such that $\Lambda_{i+1}/\Lambda_{i}$ is contained in the centre $Z(\Gamma/\Lambda_{i})$ of $\Gamma/\Lambda_{i}$ for all $i=0,\dots,n-1$.
+We say $\Gamma$ is finite-by-nilpotent if there is a finite normal subgroup $N\trianglelefteq\Gamma$ such that $\Gamma/N$ is nilpotent.
+A subgroup $X\subseteq\Gamma$ is called subnormal if there is a sequence $X=\Lambda_{0}\trianglelefteq\Lambda_{1}\trianglelefteq\cdots\trianglelefteq\Lambda_{k}=\Gamma$ such that $\Lambda_{i}$ is normal in $\Lambda_{i+1}$ for all $i<k$.
+Such a sequence is called a subnormal series.
+In a nilpotent group, every subgroup is subnormal.
+Therefore, in a finite-by-nilpotent group, every subgroup $X$ is contained as a finite index subgroup inside some subnormal subgroup $\Lambda$ ([24, Section 6.3]).
+We record a known lemma about finite-by-nilpotent groups that immediately justifies the consideration of Lemma 4.14.
+Lemma 4.15.
+Let $\Gamma$ be a finitely generated finite-by-nilpotent group. Then every subgroup of $\Gamma$ is finitely generated.
+Proof.
+Every finitely generated finite-by-nilpotent group is finite-by-polycyclic, and the property in the lemma holds for such groups ([17]). ∎
+We now see our result for finite-by-nilpotent groups and collections of finite (or co-finite) subgroups as a consequence of all our reduction and permanence properties (Theorem 4.5, Theorem 4.7, Theorem 4.9, Corollary 4.6, Proposition 4.8) together with the two previous lemmas.
 Theorem 4.16.
@@
   1. (I)
-    $\Gamma$ is countable and abelian, or
+    every $X\in\mathcal{X}$ is finite,
   2. (II)
-    every $X\in\mathcal{X}$ is finite,
+    every $X\in\mathcal{X}$ is co-finite, or
+  3. (III)
+    $\Gamma$ is countable and finite-by-nilpotent.
 then $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
 Proof.
-Let $\Gamma$ be a countable abelian group.
-By Proposition 3.2, it suffices to assume that $\Gamma$ is finitely generated.
-Then, every subgroup of $\Gamma$ is also finitely generated, so Lemma 3.10 implies that $\mathcal{X}_{\text{min}}$ is finite.
-Since $\Gamma$ is abelian, each $X\in\mathcal{X}_{\text{min}}$ is normal, and hence Corollary 3.8 implies $\mathcal{X}_{\text{min}}\in\mathcal{D}_{\Gamma}$.
-Theorem 3.9 then implies $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
-In the case where every $X\in\mathcal{X}$ is finite, we have that $\mathcal{X}_{\text{min}}$ is finite (by Lemma 3.10) and each $X\in\mathcal{X}_{\text{min}}$ is finite.
-It follows from [16, Lemma 7.9] (see also the proof of [4, Theorem 4.7]) that the normal subgroup $N$ generated by all $X\in\mathcal{X}_{\text{min}}$ is finite.
-Since $C^{*}_{r}(N)=\mathbb{C}[N]$, it holds trivially that $\mathcal{X}_{\text{min}}\in\mathcal{D}_{N}$, so Theorems 3.6 and 3.9 imply $\mathcal{X}\in\mathcal{D}_{\Gamma}.$
-The following Corollary is immediate by Theorems 2.9 and 3.11.
+(I) Assume that every $X\in\mathcal{X}$ is finite.
+By Lemma 4.14 the set $\mathcal{X}_{\text{min}}$ is finite.
+Since each subgroup $X\in\mathcal{X}_{\text{min}}$ is finite, it follows from [19, Lemma 7.9] (see also the proof of [5, Theorem 4.7]) that the normal subgroup $N$ generated by all $X\in\mathcal{X}_{\text{min}}$ is finite.
+Then $C^{*}_{r}(N)=\mathbb{C}[N]$, so it holds trivially that $\mathcal{X}_{\text{min}}\in\mathcal{D}_{N}$, and Theorems 4.5 and 4.7 imply $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
+(II) Assume every $X\in\mathcal{X}$ is co-finite.
+Note that if $\Lambda\subseteq\Gamma$ is a subgroup, then $\Lambda\cap X$ is co-finite in $\Lambda$, for every $\Lambda\cap X\in\Lambda\cap\mathcal{X}$.
+Therefore, Proposition 4.10 allows us to reduce to the special case where $\Gamma$ is finitely generated.
+Since every co-finite subgroup of a finitely generated group is finitely generated, Lemma 4.14 implies $\mathcal{X}_{\text{min}}$ is finite.
+Hence, the normal subgroup $N\coloneq\bigcap_{X\in\mathcal{X}_{\text{min}}}X$ is co-finite.
+Write $\mathcal{Y}=\{X/N\colon X\in\mathcal{X}_{\text{min}}\}$.
+Then, $\mathcal{Y}$ is a conjugation-invariant set of subgroups in the finite group $\Gamma/N$ with $q^{*}_{N}(\mathcal{Y})=\mathcal{X}_{\text{min}}$.
+Since $\mathcal{Y}\in\mathcal{D}_{\Gamma/N}$ (automatically due to finiteness of $\Gamma/N$), we have $\mathcal{X}_{\text{min}}=q^{*}_{N}(\mathcal{Y})\in\mathcal{D}_{\Gamma}$ by Theorem 4.9, and hence $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Theorem 4.7.
+(III) Let $\Gamma$ be a countable and finite-by-nilpotent group.
+Our goal is to show that $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
+Every subgroup of $\Gamma$ is finite-by-nilpotent so, by Proposition 4.10, we can assume that $\Gamma$ is finitely generated.
+Then, by Lemma 4.15, every subgroup of $\Gamma$ is finitely generated, so Lemma 4.14 implies that $\mathcal{X}_{\text{min}}$ is finite.
+By Theorem 4.7 it therefore suffices to assume that $\mathcal{X}$ is finite.
+We show that $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by induction on $|\mathcal{X}|$.
+The base case $|\mathcal{X}|=1$ is covered by Corollary 4.6.
+Suppose that $|\mathcal{X}|\geq 2$, and assume that whenever a conjugation-invariant collection of subgroups in a finite-by-nilpotent group $\Lambda$ has cardinality less than $|\mathcal{X}|$, then it belongs to $\mathcal{D}_{\Lambda}$.
+Let $N\subseteq\Gamma$ be the subgroup generated by all elements of $\mathcal{X}$.
+If $N$ contains every element of $\mathcal{X}$ as a finite index subgroup, then $\mathcal{X}\in\mathcal{D}_{N}$ by (II) above, and hence $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Theorem 4.5.
+Suppose otherwise, and choose $X_{0}\in\mathcal{X}$ with infinite index in $N$.
+Since $\Gamma$ is finite-by-nilpotent, there exists a subnormal subgroup $\Lambda\subseteq\Gamma$ containing $X_{0}$ as a finite-index subgroup.
+Note that $N\nsubseteq\Lambda$, by assumption.
+Let $\Gamma=\Lambda_{0}\trianglerighteq\Lambda_{1}\trianglerighteq\dots\trianglerighteq\Lambda_{n}=\Lambda$ be a subnormal series, and let $k\in\{1,\dots,n\}$ be minimal such that $N\nsubseteq\Lambda_{k}$.
+Define non-empty subsets $\mathcal{X}_{1}\coloneq\{X\in\mathcal{X}\colon X\subseteq\Lambda_{k}\}$ and $\mathcal{X}_{2}=\mathcal{X}\setminus\mathcal{X}_{1}$.
+Since $\Lambda_{k}$ is normal in $\Lambda_{k-1}$, both $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ are conjugation-invariant collections of subgroups in the finite-by-nilpotent group $\Lambda_{k-1}$ with cardinality strictly less than $|\mathcal{X}|$.
+By the inductive hypothesis, both $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ belong to $\mathcal{D}_{\Lambda_{k-1}}$.
+By Proposition 4.8 we have $\mathcal{X}=\mathcal{X}_{1}\cup\mathcal{X}_{2}\in\mathcal{D}_{\Lambda_{k-1}}$, and hence $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Theorem 4.5.
+This completes the proof of the theorem.
+∎
+The following corollary is immediate by Theorems 3.4, Theorem 4.16 and Corollary 4.13, where $\mathcal{X}(x)$ is as defined in Definition 1.3.
 Corollary 4.17.
-Let $G$ be an amenable and second-countable étale groupoid, and let $\tilde{G}$ denote the Hausdorff cover groupoid. Assume for every $x\in G^{(0)}$ one of the following holds.
+Let $G$ be an amenable and second-countable étale groupoid. For each $x\in G^{(0)}$ assume that one of the following holds.
   1. (I)
-    The isotropy group $G_{x}^{x}$ is abelian.
+    The isotropy group $G_{x}^{x}$ is finite-by-nilpotent.
   2. (II)
-    The subgroups $X\in\mathcal{X}(x)$ are finite.
+    The subgroups $X\in\mathcal{X}(x)$ are all finite.
+  3. (III)
+    The subgroups $X\in\mathcal{X}(x)$ are all co-finite in $G^{x}_{x}$.
+  4. (IV)
+    The subgroups $X\in\mathcal{X}(x)$ are all normal in $G^{x}_{x}$.
 Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
 Note that condition (II) holds whenever all elements of $\tilde{G}^{(0)}_{\rm{ess}}$ are finite.
-The class of groupoids described in Corollary 3.12 (II) contains all amenable and second-countable étale groupoids with finite isotropy groups, and all groupoids of contracting self-similar groups (see [4, Corollary 7.13]).
-More generally, condition (II) is satisfied whenever the unit space of $G$ has finite source and range fibres.
-Therefore, within the class of amenable and second-countable étale groupoids, the class described in Corollary 3.12 (II) is larger than the class described in [4, Theorem 4.7].
-4. Ample Groupoids
-The main result of this section is that the algebraic singular ideal $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ for any ample groupoid.
-This reduces the question of whether $J_{\mathbb{C}}$ is dense in $J$ to the corresponding question for $J\cap\mathscr{C}_{c}(G)$.
-In order to prove this result we will need two preliminary lemmas.
+Therefore, within the class of amenable and second-countable étale groupoids, the class described in Corollary 4.17 (II) is larger than the class described in [5, Theorem 4.7].
+The final result of this section considers étale group bundles – that is, étale groupoids whose range and source maps are equal.
+Corollary 4.18.
+Let $G$ be an amenable and second-countable étale group bundle. Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
+Proof.
+We show that for every $x\in G^{(0)}$, the subgroups $X\in\mathcal{X}(x)$ are all normal in $G^{x}_{x}$. The result then follows by Corollary 4.17 (IV).
+Let $x\in G^{(0)}$, $X\in\mathcal{X}(x)$ and choose a net $(x_{\alpha})\subset G^{(0)}$ whose set of limit points is equal to $X$.
+Take $g\in G^{x}_{x}$ and let $U\subseteq G$ be an open bisection containing $g$.
+The $x_{\alpha}$ eventually belong to $s(U)$, so there exist $g_{\alpha}\in U$ with $s(g_{\alpha})=x_{\alpha}$.
+Since $s|_{U}\colon U\to s(U)$ is a homoeomorphism, the net $(g_{\alpha})$ converges to $g$.
+Then, since $g_{\alpha}=g_{\alpha}x_{\alpha}$ for all $\alpha$, continuity of multiplication on the groupoid ensures that $(g_{\alpha})$ has set of limit points equal to $gX$.
+However, $r(g_{\alpha})=s(g_{\alpha})$, so we have $g_{\alpha}=x_{\alpha}g_{\alpha}$ for all $\alpha$, and hence continuity of multiplication also ensures that $(g_{\alpha})$ has set of limit points equal to $Xg$.
+Therefore $gX=Xg$.
+The element $g\in G^{x}_{x}$ was arbitrary, so $X$ is a normal subgroup of $G^{x}_{x}$, as required.
+∎
+5. Structure of $J\cap\mathscr{C}_{c}(G)$
+In this section we provide explicit instructions for building elements of $J\cap\mathscr{C}_{c}(G)$, and show that functions constructed in this way have dense linear span in $J\cap\mathscr{C}_{c}(G)$.
+Therefore, when $J\cap\mathscr{C}_{c}(G)$ is dense in the singular ideal $J$, we are able to describe an explicit family of functions in $\mathscr{C}_{c}(G)$ whose linear span is dense in $J$.
+Let us describe the approach.
+One starts with a unit $x\in G^{(0)}$ and a finitely supported function $b\colon G^{x}_{x}\to\mathbb{C}$ that satisfies the linear equations (0.1).
+Given $h\in G_{x}$, translate the support of $b$ by left-multiplication, and choose open bisections $V_{1},\dots,V_{n}$ containing these translates.
+Lastly, choose an open neighbourhood $W\subseteq\bigcap_{i=1}^{n}s(V_{i})$ of $x$ such that the open bisections $U_{i}\coloneq V_{i}\cap s^{-1}(W)$ satisfy the conditions of Lemma 5.2.
+Then, for any $\psi\in C_{c}(W)$, the function $f^{\psi}$ (defined in (5.2)) belongs to $J\cap\mathscr{C}_{c}(G)$ (see Proposition 5.3).
+Overall, the function $f^{\psi}$ is determined by data in the tuple $(x,h,b,U_{1},\dots,U_{n},\psi)$.
+By ranging over all valid tuples, one obtains a family of functions whose linear span is dense in $J\cap\mathscr{C}_{c}(G)$ (see Theorem 5.5).
+We remark that since $\mathcal{X}(x)$ is invariant under conjugation, an element $b\in\mathbb{C}[G^{x}_{x}]$ satisfies (0.1) if and only if, for every subgroup $X\in\mathcal{X}(x)$, $b$ belongs to the kernel of the associated quasi-regular representation $\lambda_{G^{x}_{x}/X}$.
+5.1. Preliminary lemmas
+We will need three preliminary results.
 Lemma 5.1.
-Let $G$ be an étale groupoid, and let $f\in\mathscr{C}_{c}(G)$. Let $x\in G^{(0)}$, and assume that $f(g)=0$ for all $g\in G_{x}$. Then $f$ is continuous at each $g\in G_{x}$.
+Let $G$ be an étale groupoid. Fix $f\in\mathscr{C}_{c}(G)$, $x\in G^{(0)}$, and assume that $f(g)=0$ for all $g\in G_{x}$. Then $f$ is continuous at each $g\in G_{x}$.
 Proof.
@@
 ∎
-Observations similar to the following lemma have appeared in both [4, Theorem 4.7] and [16, Proposition 5.19]. Our proof uses the same techniques as those developed in the aforementioned papers.
+The following results isolate observations from [19, Proposition 5.19].
+We follow the same proofs (see also [5, Theorem 4.7]).
+Given $x\in G^{(0)}$, recall that a subgroup $X\subseteq G^{x}_{x}$ belongs to $\mathcal{X}(x)$ if and only if there exists a net $(x_{\alpha})\subset C$ of (Hausdorff) units whose set of limit points is precisely $X$, and such that every subnet of $(x_{\alpha})$ has its limit points contained in $X$ (see Definition 1.3).
 Lemma 5.2.
-Let $G$ be an étale groupoid.
-Fix $f\in J\cap\mathscr{C}_{c}(G)$ and $x_{0}\in G^{(0)}$.
-Let $g_{1},\dotsc,g_{n}\in G_{x_{0}}$ be all points in $G_{x_{0}}$ on which $f$ is non-zero, and let $U_{1},\dotsc,U_{n}$ be open bisections such that $U_{i}$ contains $g_{i}$ for each $1\leq i\leq n$.
-Set $V\coloneq\bigcap_{i=1}^{n}s(U_{i})$.
-Then there exists an open subset $W\subseteq V$ containing $x_{0}$ such that
-$f^{\psi}\coloneq\sum_{i=1}^{n}f(g_{i})(\psi\circ s|_{U_{i}})$ (4.1)
-satisfies $f^{\psi}\in J$ for any $\psi\in C_{c}(W)$.
+Let $G$ be an étale groupoid that is covered by countably many open bisections. Fix $x\in G^{(0)}$, and let $g_{1},\dots,g_{n}\in G_{x}$ be distinct points.
+There exist open bisections $U_{1},\dots,U_{n}\subseteq G$ satisfying the following conditions:
+  1. (i)
+    $g_{i}\in U_{i}$ for all $i\in\{1,\dots,n\}$;
+  2. (ii)
+    $s(U_{i})=s(U_{j})$ for all $i,j\in\{1,\dots,n\}$;
+  3. (iii)
+    $W_{I}=\emptyset$ whenever $I\subseteq\{1,\dotsc,n\}$ is not of the form
+    (5.1) $\{g_{i}\colon i\in I\}=kX\cap\{g_{1},\dots,g_{n}\}$
+    for some $k\in G_{x}$ and $X\in\mathcal{X}(x)$.
+Here $W_{I}\coloneq s\big(\bigcap_{i\in I}U_{i}\setminus\bigcup_{j\notin I}U_{j}\big)\cap C$.
+Note that for any small open set $U\subseteq G^{(0)}$ containing $x$, the cutdowns $U_{i}^{\prime}\coloneq U_{i}\cap s^{-1}(U)$ also satisfy (i) – (iii).
 Proof.
-Without loss of generality, we assume $s(U_{i})=V$ for all $1\leq i\leq n$.
-For each open set $W\subseteq V$, let $S^{W}_{i}\coloneq U_{i}\cap s^{-1}(W)$ for $1\leq i\leq n$.
-For $I\subseteq\{1,\dotsc,n\}$, set $\check{S}^{W}_{I}\coloneq(\bigcap_{i\in I}S^{W}_{i})\setminus(\bigcup_{j\notin I}S^{W}_{j})$ as well as $W_{I}\coloneq\text{int}\Big(\overline{s(\check{S}^{W}_{I})}^{G^{(0)}}\Big)$, where $\overline{(\,\cdot\,)}^{G^{(0)}}$ denotes closure in $G^{(0)}$.
-Our first claim is that there exists an open set $W\subseteq V$ containing $x_{0}$ such that
-$W_{I}\neq\emptyset\text{ implies that }x_{0}\in\overline{W_{I}}^{G^{(0)}}\text{ for all }I\subseteq\{1,\dotsc,n\}.$ (4.2)
-If $W=V$ satisfies (4.2), then we are done.
-Otherwise, choose a regular open subset $W$ of $G^{(0)}$ (i.e., $W=\text{int}\big(\overline{W}^{G^{(0)}}\big)$) such that $x_{0}\in W\subseteq V$ and $W\cap\overline{V_{I}}^{G^{(0)}}=\emptyset$ for all $I\subseteq\{1,\dotsc,n\}$ with $V_{I}\neq\emptyset$ and $x_{0}\notin\overline{V_{I}}^{G^{(0)}}$.
-Note that $\check{S}^{W}_{I}=\check{S}^{V}_{I}\cap s^{-1}(W)$, and therefore $s(\check{S}^{W}_{I})=s(\check{S}^{V}_{I})\cap W$.
-Since $W$ is open, we have $W_{I}=V_{I}\cap W$, and therefore $\overline{W_{I}}^{G^{(0)}}\cap W=\overline{V_{I}}^{G^{(0)}}\cap W$.
-Also $W_{I}=\text{int}\Big(\overline{s(\check{S}^{W}_{I})}^{G^{(0)}}\Big)\subseteq\text{int}\Big(\overline{s(\check{S}^{V}_{I})}^{G^{(0)}}\cap\overline{W}^{G^{(0)}}\Big)\subseteq V_{I}\cap\text{int}\big(\overline{W}^{G^{(0)}}\big)=V_{I}\cap W=\emptyset$ whenever $x_{0}\notin\overline{V_{I}}^{G^{(0)}}$.
-This completes the proof of the claim.
-Define $\mathcal{I}\coloneq\{I\subseteq\{1,\dotsc,n\}\colon W_{I}\neq\emptyset\}$. Our second claim is that
-$\sum_{i\in I}f(g_{i})=0\text{ for all }I\in\mathcal{I}.$ (4.3)
-By [4, Lemma 4.1(ii)], $\overline{s({\rm supp}^{\circ}(f))}^{G^{(0)}}$ has empty interior.
-Let $I\in\mathcal{I}$, and set $\check{W}_{I}\coloneq W_{I}\setminus\overline{s({\rm supp}^{\circ}(f))}^{G^{(0)}}$.
-Then, $\check{W}_{I}$ is open and dense in $W_{I}$.
-Therefore $x_{0}$ lies in the closure of $\check{W}_{I}$ in $G^{(0)}$ by (4.2).
-As $s(\check{S}_{I})$ is dense in $W_{I}$ and because $\check{W}_{I}$ is open, we conclude that there is a net $(x_{\beta})$ in $\check{W}_{I}\cap s(\check{S}_{I})$ converging to $x_{0}$ in $G^{(0)}$.
-Let $g_{\beta}$ be the unique element of $\check{S}_{I}$ such that $s(g_{\beta})=x_{\beta}$.
-By passing to a subnet if necessary, we may assume that $(g_{\beta})$ converges to some $\bm{g}$ in $\tilde{G}$ (i.e., in the Fell topology).
-By construction, we have that $\bm{g}\cap\{g_{1},\dotsc,g_{n}\}=\{g_{i}\colon i\in I\}$.
-Therefore,
-$\lim_{\beta}f(g_{\beta})=\lim_{\beta}\mathfrak{i}(f)(\iota(g_{\beta}))=\mathfrak{i}(f)(\bm{g})=\sum_{i\in I}f(g_{i}),$
-where the third equality used (1.2).
-Now $s(g_{\beta})\in\check{W}_{I}$ implies that $g_{\beta}\notin{\rm supp}^{\circ}(f)$, so that $f(g_{\beta})=0$.
-We conclude that $\sum_{i\in I}f(g_{i})=0$, proving (4.3).
-Take arbitrary $\psi\in C_{c}(W)$, and let $f^{\psi}$ be as in (4.1).
-We show that $f^{\psi}\in J$.
-For $I\in\mathcal{I}$ and $g\in\check{S}^{W}_{I}$ we have $f^{\psi}(g)=\sum_{i\in I}f(g_{i})\psi(s(g))=0$ using (4.3).
-It follows that ${\rm supp}^{\circ}(f^{\psi})\subseteq\bigcup_{I\,\in\,\mathcal{P}(\{1,\dotsc,n\})\setminus\mathcal{I}}\check{S}^{W}_{I}$ and therefore $s({\rm supp}^{\circ}(f^{\psi}))\subseteq\bigcup_{I\,\in\,\mathcal{P}(\{1,\dotsc,n\})\setminus\mathcal{I}}s(\check{S}^{W}_{I})$, where $\mathcal{P}$ denotes power set.
-Hence $s({\rm supp}^{\circ}(f^{\psi}))$ has empty interior because it is contained in a finite union of nowhere dense sets.
-This shows that $f^{\psi}\in J$, as desired.
+It is clear that there exist open bisections $V_{1},\dots,V_{n}\subseteq G$ satisfying the first two conditions.
+Write $V\coloneq s(V_{1})$ and $V_{I}\coloneq s\big(\bigcap_{i\in I}U_{i}\setminus\bigcup_{j\notin I}U_{j}\big)\cap C$ for subsets $I\subseteq\{1,\dots,n\}$.
+Choose an open subset $W\subseteq G^{(0)}$ such that $x\in W\subseteq V$ and $W\cap\overline{V_{I}}^{G^{(0)}}=\emptyset$ for all $I\subseteq\{1,\dotsc,n\}$ with $x\notin\overline{V_{I}}^{G^{(0)}}$.
+Define $U_{i}\coloneq V_{i}\cap s^{-1}(W)$ for $i\in\{1,\dots,n\}$.
+It is clear that the open bisections $U_{1},\dots,U_{n}\subseteq G$ satisfy conditions (i) and (ii).
+We claim that they also satisfy condition (iii).
+Let $I\subseteq\{1,\dots,n\}$ be such that $W_{I}\neq\emptyset$.
+We claim that $I$ is of the form in (5.1).
+It is clear that $W_{I}=W\cap V_{I}$ and hence $x\in\overline{W_{I}}^{G^{(0)}}$ by the construction of $W$.
+Let $(x_{\alpha})$ be a net in $W_{I}$ converging to $x$.
+By passing to a subnet if necessary we can assume that the net $\iota(x_{\alpha})$ converges in the Hausdorff cover groupoid $\tilde{G}$.
+Let $X\in\tilde{G}^{(0)}_{\rm{ess}}$ denote the limit.
+Note that $X\in\mathcal{X}(x)$ by definition.
+Also, by definition of the Fell topology, the set of limit points of the net $(x_{\alpha})$ is precisely $X$.
+Now, let $g_{\alpha}\in\bigcap_{i\in I}U_{i}\setminus\bigcup_{j\notin I}U_{j}$ be the unique elements satisfying $s(g_{\alpha})=x_{\alpha}$.
+Continuity of multiplication on the groupoid implies that the net $(g_{\alpha})$ has set of limit points $g_{i_{0}}X$, where $i_{0}$ is any element of $I$.
+The unit $x$ belongs to $X$, and hence $\{g_{i}\colon i\in I\}\subseteq g_{i_{0}}X$.
+On the other hand, if $j\notin I$ then $g_{\alpha}\notin U_{j}$ and hence the net $(g_{\alpha})$ does not converge to $g_{j}$.
+Therefore $g_{j}\notin g_{i_{0}}X$, and it follows that $\{g_{i}\colon i\in I\}=g_{i_{0}}X\cap\{g_{1},\dots,g_{n}\}$ for any $i_{0}\in I$.
+This completes the proof.
 ∎
-Remark 4.3.
-Let $f\in J$, $x_{0}\in G^{(0)}$ and assume that there are only finitely many points $g_{1},\dotsc,g_{n}$ in the isotropy group $G^{x_{0}}_{x_{0}}$ on which $f$ non-zero.
-Let $U_{1},\dotsc,U_{n},V$ be as in the statement of Lemma 4.2.
-If in addition we assume that the isotropy group $G^{x_{0}}_{x_{0}}$ is amenable, then the conclusion of Lemma 4.2 still holds.
-That is, there exists an open set $W\subseteq V$ containing $x_{0}$ such that $f^{\psi}\in J$ for any $\psi\in C_{c}(W)$, where $f^{\psi}$ is defined as in (4.1).
-If the isotropy group $G^{x_{0}}_{x_{0}}$ is non-amenable, then there may exist no such $W$.
-We now prove the main result of this section.
-Recall that an ample groupoid is an étale groupoid with a basis of compact open bisections.
-If $G$ is an ample groupoid, the Steinberg algebra $\mathbb{C}G$ over the complex numbers is defined in [26] as
-$\mathbb{C}G\coloneq\text{span}\left\{\mathbbm{1}_{U}\colon U\subseteq G\text{ a compact open bisection}\right\}$
-The algebraic singular ideal is the intersection $J_{\mathbb{C}}\coloneq J\cap\mathbb{C}G$.
-Since $\mathbb{C}G\subseteq\mathscr{C}_{c}(G)$ we have $J_{\mathbb{C}}\subseteq J\cap\mathscr{C}_{c}(G)$.
+The next proposition is a slight generalisation of the construction of singular elements found in [19, Proposition 5.19]. The proof is exactly the same, but we record it for completeness.
+Proposition 5.3.
+Let $G$ be an étale groupoid that is covered by countably many open bisections.
+Fix $x\in G^{(0)}$, and suppose $b\in\mathbb{C}[G^{x}_{x}]$ belongs to $\ker(\lambda_{G^{x}_{x}/X})$ for all $X\in\mathcal{X}(x)$.
+Let $\{g_{1},\dots,g_{n}\}\subseteq G^{x}_{x}$ denote the support of $b$, let $h\in G_{x}$, and let $U_{1},\dots,U_{n}$ be open bisections containing $hg_{1},\dots,hg_{n}$ and satisfying the conditions of Lemma 5.2.
+Then
+(5.2) $f^{\psi}\coloneq\sum_{i=1}^{n}b(g_{i})(\psi\circ s|_{U_{i}})$
+belongs to $J\cap\mathscr{C}_{c}(G)$ for any $\psi\in C_{c}\big(s(U_{i})\big)$ (note that this does not depend on $i$).
+Proof.
+For $I\subseteq\{1,\dots,n\}$ set $\check{S}_{I}=(\bigcap_{i\in I}U_{i})\setminus(\bigcup_{j\notin I}U_{j})$ and $W_{I}=s(\check{S}_{I})\cap C$, as in Lemma 5.2.
+Take arbitrary $\psi\in C_{c}\big(s(U_{i})\big)$, and let $f^{\psi}$ be as in (5.2).
+If $W_{I}\neq\emptyset$, Lemma 5.2(iii) ensures that $\{hg_{i}\colon i\in I\}=kX\cap\{hg_{1},\dots,hg_{n}\}$ for some $k\in G_{x}$ and $X\in\mathcal{X}(x)$.
+Clearly $h^{-1}k\in G^{x}_{x}$.
+Therefore, for $g\in\check{S}_{I}$ we have
+$f^{\psi}(g)=\sum_{i\in I}b(g_{i})\psi(s(g))=\sum_{t\in h^{-1}kX}b(t)\psi(s(g))=0$
+since $b\in\ker(\lambda_{G^{x}_{x}/X})$.
+It follows that any $g\in{\rm supp}^{\circ}(f^{\psi})$ belongs to $\check{S}_{I}$ for some $I\subseteq\{1,\dots,n\}$ with $W_{I}=\emptyset$, and hence $s\big({\rm supp}^{\circ}(f^{\psi})\big)\subseteq G^{(0)}\setminus C$.
+This implies that $f^{\psi}\in J$ by [22, Proposition 7.18], as desired.
+∎
+$\bullet$
+5.2. Building blocks for $J\cap\mathscr{C}_{c}(G)$
+Let $G$ be an étale groupoid that is covered by countably many open bisections. Each building block is assembled from the data in a tuple $\Xi\coloneq(x,h,b,U_{1},\dots,U_{n},\psi)$ where
+  • $x\in G^{(0)}$;
+  • $h\in G_{x}$;
+  • $b\in\mathbb{C}[G^{x}_{x}]$ belongs to $\ker(\lambda_{G^{x}_{x}/X})$ for all $X\in\mathcal{X}(x)$, and has support $\{g_{1},\dots,g_{n}\}$;
+  • $U_{1},\dots,U_{n}\subseteq G$ are open bisections satisfying the conditions of Lemma 5.2.
+That is, they have equal source, contain the elements $hg_{1},\dots,hg_{n}\in G_{x}$, and are such that $W_{I}=\emptyset$ whenever $I\subseteq\{1,\dotsc,n\}$ is not of the form
+    $\{hg_{i}\colon i\in I\}=kX\cap\{hg_{1},\dots,hg_{n}\}\quad\quad\text{for some }k\in G_{x},X\in\mathcal{X}(x).$
+    Here $W_{I}\coloneq s\big(\bigcap_{i\in I}U_{i}\setminus\bigcup_{j\notin I}U_{j}\big)\cap C$.
+  • $\psi\in C_{c}\big(s(U_{i})\big)$ (note that this does not depend on $i$).
+Given such a tuple $\Xi$ we define
+(5.3) $f^{\Xi}\coloneq\sum_{i=1}^{n}b(g_{i})(\psi\circ s|_{U_{i}}).$
+By Proposition 5.3 we have $f^{\Xi}\in J\cap\mathscr{C}_{c}(G)$.
+Let $\mathcal{E}\subseteq J\cap\mathscr{C}_{c}(G)$ denote the linear span of elements of the form $f^{\Xi}$ from (5.3), where we let $\Xi$ range over all valid tuples.
+We will see that $\mathcal{E}$ is dense in $J\cap\mathscr{C}_{c}(G)$.
+Lemma 5.4.
+Let $G$ be an étale groupoid that is covered by countably many open bisections.
+Fix $f\in J\cap\mathscr{C}_{c}(G)$, $x\in G^{(0)}$, and let $g_{1},\dotsc,g_{n}\in G_{x}$ be all elements of $G_{x}$ on which $f$ is non-zero.
+There exist open bisections $U_{1},\dots,U_{n}$ with equal source, containing $g_{1},\dots,g_{n}$ and such that
+$f^{\psi}\coloneq\sum_{i=1}^{n}f(g_{i})(\psi\circ s|_{U_{i}})$
+belongs to $\mathcal{E}$ for any $\psi\in C_{c}\big(s(U_{i})\big)$ (note that this does not depend on $i$).
+Proof.
+Let $U_{1},\dots,U_{n}$ be open bisections satisfying the conditions of Lemma 5.2 and such that $U_{i_{1}}$ and $U_{i_{2}}$ are disjoint whenever $r(g_{i_{1}})\neq r(g_{i_{2}})$.
+Partition the set $\{g_{1},\dots,g_{n}\}$ according to the images under the range map $r$ and write $\{g_{1},\dots,g_{n}\}=\bigsqcup_{j=1}^{m}\{g^{j}_{1},\dots,g^{j}_{k_{j}}\}$, $\{U_{1},\dots,U_{n}\}=\bigsqcup_{j=1}^{m}\{U^{j}_{1},\dots,U^{j}_{k_{j}}\}$, where $g_{i}^{j}\in U_{i}^{j}$.
+Fix $\psi\in C_{c}\big(s(U_{i})\big)$.
+For each $j\in\{1,\dots,m\}$ define
+$f^{j}\coloneq\sum_{i=1}^{k_{j}}f\big(g^{j}_{i}\big)\big(\psi\circ s|_{U^{j}_{i}}\big)$
+so that $f^{\psi}=\sum_{j=1}^{m}f^{j}$. For each $j$, we show that $f^{j}=f^{\Xi_{j}}$ for some valid tuple $\Xi_{j}$, where $f^{\Xi_{j}}$ is as in (5.3). The lemma then follows.
+Let $V_{j}$ be an open bisection containing ${(g^{j}_{1})}^{-1}$, and let $\varphi\in C_{c}(V_{j})$ be such that $\varphi\big(({g^{j}_{1}}\big)^{-1})=1$.
+Define $b_{j}\coloneq\eta_{x}(\varphi*f)$, where $\eta_{x}\colon\mathscr{C}_{c}(G)\to\mathbb{C}[G^{x}_{x}]$ is the restriction map.
+Since $\varphi*f\in J\cap\mathscr{C}_{c}(G)$, it follows from [19, Proposition 5.19] that $b_{j}$ belongs to $\ker(\lambda_{G^{x}_{x}/X})$ for all $X\in\mathcal{X}(x)$.
+Then, since $f(g_{i}^{j})=b_{j}\big({(g^{j}_{1})}^{-1}g^{j}_{i}\big)$ for all $i=1,\dots,k_{j}$, we have $f^{j}=f^{\Xi_{j}}$ where $\Xi_{j}\coloneq\big(x,g_{1}^{j},b_{j},U^{j}_{1},\dots,U^{j}_{k_{j}},\psi\big)$.
+The initial assumptions on the bisections $U_{1},\dots,U_{n}$ ensure that $\Xi_{j}$ is a valid tuple.
+∎
 For $f\in\mathscr{C}_{c}(G)$, define
-$\left\|f\right\|_{I}\coloneq\max\left\{\sup_{x\in G^{(0)}}\sum_{g\in G_{x}}\left|f(g)\right|,\sup_{x\in G^{(0)}}\sum_{g\in G^{x}}\left|f(g)\right|\right\}.$
+(5.4) $\left\|f\right\|_{I}\coloneq\max\left\{\sup_{x\in G^{(0)}}\sum_{g\in G_{x}}\left|f(g)\right|,\sup_{x\in G^{(0)}}\sum_{g\in G^{x}}\left|f(g)\right|\right\},$
+following [16].
 Theorem 5.5.
-Let $G$ be an ample groupoid. Then $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to the norm $\left\|\cdot\right\|_{I}$.
+Let $G$ be an étale groupoid that is covered by countably many open bisections. Then $\mathcal{E}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to the norm $\left\|\cdot\right\|_{I}$.
 The norm $\left\|\cdot\right\|_{I}$ dominates the full norm $\left\|\cdot\right\|_{C^{*}(G)}$ and hence any $C^{*}$-norm.
-Therefore, $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
+Therefore, $\mathcal{E}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
 Proof.
@@
 Whenever $\dot{f}\in\mathscr{C}_{c}(G)$ satisfies ${\rm supp}^{\circ}(\dot{f})\subseteq K$ we have $\left\|\dot{f}\right\|_{I}\leq C_{K}\left\|\dot{f}\right\|_{\infty}$ where $\left\|\cdot\right\|_{\infty}$ denotes the supremum norm, and $C_{K}$ is a constant depending only on $K$.
-Therefore, it suffices to find, for each $\varepsilon>0$, a function $f^{\prime}\in J_{\mathbb{C}}$ satisfying ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$.
+Therefore, it suffices to find, for each $\varepsilon>0$, a function $f^{\prime}\in\mathcal{E}$ satisfying ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$.
 Fix $\varepsilon>0$.
-Take $x\in s(K)$, and let $g_{1},\dotsc,g_{m}\in G_{x}$ be all points in $G_{x}$ on which $f$ is non-zero.
-By Lemma 4.2, there exists a compact open neighbourhood $U_{x}\subseteq G^{(0)}$ of $x$, and compact open bisections $U_{1},\dotsc,U_{m}\subseteq K$ satisfying the following: $g_{i}\in U_{i}$ and $s(U_{i})=U_{x}$ for all $i=1,\dotsc,m$, and the function $f^{x}$ defined by
-$f^{x}\coloneq\sum_{i=1}^{m}f(g_{i})\mathbbm{1}_{U_{i}}$
-lies in $J_{\mathbb{C}}$.
-Clearly ${\rm supp}^{\circ}(f^{x})\subseteq K$.
-We have $f|_{G_{x}}=f^{x}|_{G_{x}}$ and ${\rm supp}^{\circ}(f-f^{x})\subseteq K$, a compact set.
-Therefore, by Lemma 4.1, there exists a compact open neighbourhood $W_{x}\subseteq U_{x}$ of $x$ such that $\left|f(g)-f^{x}(g)\right|\leq\varepsilon$ whenever $s(g)\in W_{x}$.
-The sets $\{W_{x}\}_{x\in s(K)}$ form an open cover for the compact set $s(K)$, so select a finite subcover $W_{x_{1}},\dotsc,W_{x_{n}}$.
-Removing intersections if necessary, we may assume that the $W_{x_{j}}$ are pairwise disjoint (this is possible since the $W_{x_{i}}$ are contained in the Hausdorff space $G^{(0)}$, and are thus clopen in $G^{(0)}$).
+Take $x\in s(K)$, and let $g_{1},\dotsc,g_{n}\in G_{x}$ be all elements of $G_{x}$ on which $f$ is non-zero.
+Let $U_{1},\dots,U_{n}\subseteq K$ be open bisections as in Lemma 5.4, and define $U_{x}\coloneq s(U_{i})$ (note that this does not depend on $i$).
+In particular $x\in U_{x}$.
+Then
+$f^{\psi}\coloneq\sum_{i=1}^{n}f(g_{i})(\psi\circ s|_{U_{i}})$
+belongs to $\mathcal{E}$ for any $\psi\in C_{c}(U_{x})$.
+Fix some $\psi_{x}\in C_{c}(U_{x})$ with $\psi_{x}(x)=1$.
+Clearly ${\rm supp}^{\circ}(f^{\psi_{x}})\subseteq K$.
+We have $f|_{G_{x}}=f^{\psi_{x}}|_{G_{x}}$ and ${\rm supp}^{\circ}(f-f^{\psi_{x}})\subseteq K$, so by Lemma 5.1 there exists an open neighbourhood $W_{x}\subseteq U_{x}$ of $x$ such that $\left|f(g)-f^{\psi_{x}}(g)\right|\leq\varepsilon$ whenever $s(g)\in W_{x}$.
+The sets $\{W_{x}\}_{x\in s(K)}$ form an open cover for the compact set $s(K)$, so select a finite subcover $W_{x_{1}},\dotsc,W_{x_{m}}$.
+Let $\varphi_{1},\dots,\varphi_{m}\in C_{c}\big(G^{(0)}\big)$ be a partition of unity for $s(K)$ subordinate to this cover, and define
+$f^{\prime}\coloneq\sum_{j=1}^{m}(\varphi_{j}\circ s)f^{\psi_{x_{j}}}=\sum_{j=1}^{m}f^{\varphi_{j}.\psi_{x_{j}}}.$
+Then $f^{\prime}\in\mathcal{E}$ satisfies ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$, as required. ∎
+Remark 5.6.
+The construction in this subsection generalises to groupoids $G$ that are not covered by countably many open bisections.
+For an open subgroupoid $H\subseteq G$ that is covered by countably many open bisections, define $\mathcal{E}_{H}$ to be the linear span of functions of the form $f^{\Xi}$ in (5.3) for tuples $\Xi$ in $H$.
+Then, define $\mathcal{E}\subseteq J\cap\mathscr{C}_{c}(G)$ to be the subspace generated by all the $\mathcal{E}_{H}$.
+With this modification, the conclusion of Theorem 5.5 holds for any étale groupoid.
+6. Ample Groupoids
+In this section we describe an explicit spanning set for the algebraic singular ideal $J_{\mathbb{C}}$ of an ample groupoid.
+The construction is similar to that in Subsection 5.2.
+In Subsection 6.2 we prove that the algebraic singular ideal $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ for any ample groupoid.
+Recall that an ample groupoid is an étale groupoid with a basis of compact open bisections.
+If $G$ is an ample groupoid, the Steinberg algebra $\mathbb{C}G$ over the complex numbers is defined in [33] as
+$\mathbb{C}G\coloneq\text{span}\left\{\mathbbm{1}_{U}\colon U\subseteq G\text{ a compact open bisection}\right\}.$
+The algebraic singular ideal is the intersection $J_{\mathbb{C}}\coloneq J\cap\mathbb{C}G$. Since $\mathbb{C}G\subseteq\mathscr{C}_{c}(G)$, we have $J_{\mathbb{C}}\subseteq J\cap\mathscr{C}_{c}(G)$.
+$\bullet$
+6.1. Building blocks for $J_{\mathbb{C}}$
+Let $G$ be an ample groupoid that is covered by countably many open bisections. Each building block is assembled from the data in a tuple $\Theta\coloneq(x,h,b,U_{1},\dots,U_{n})$ where
+  • $x\in G^{(0)}$;
+  • $h\in G_{x}$;
+  • $b\in\mathbb{C}[G^{x}_{x}]$ belongs to $\ker(\lambda_{G^{x}_{x}/X})$ for all $X\in\mathcal{X}(x)$, and has support $\{g_{1},\dots,g_{n}\}$;
+  • $U_{1},\dots,U_{n}\subseteq G$ are compact open bisections satisfying the conditions of Lemma 5.2.
+That is, they have equal source, contain the elements $hg_{1},\dots,hg_{n}\in G_{x}$, and are such that $W_{I}=\emptyset$ whenever $I\subseteq\{1,\dotsc,n\}$ is not of the form
+    $\{hg_{i}\colon i\in I\}=kX\cap\{hg_{1},\dots,hg_{n}\}\quad\quad\text{for some }k\in G_{x},X\in\mathcal{X}(x).$
+    Here $W_{I}\coloneq s\big(\bigcap_{i\in I}U_{i}\setminus\bigcup_{j\notin I}U_{j}\big)\cap C$.
+Given such a tuple $\Theta$ we define
+(6.1) $f^{\Theta}\coloneq\sum_{i=1}^{n}b(g_{i})\mathbbm{1}_{U_{i}}.$
+Then $f^{\Theta}\in J_{\mathbb{C}}$ by Proposition 5.3.
+Theorem 6.1.
+Let $G$ be an ample groupoid that is covered by countably many open bisections. The algebraic singular ideal $J_{\mathbb{C}}$ is linearly spanned by elements of the form $f^{\Theta}$ from (6.1).
+Proof.
+Let $f\in J_{\mathbb{C}}$.
+Take a compact set $K\subseteq G$ with ${\rm supp}^{\circ}(f)\subseteq K$.
+Fix $x\in s(K)$, and let $g_{1},\dots,g_{n}$ be all elements of $G_{x}$ on which $f$ is non-zero.
+An argument identical to that in Lemma 5.4 yields compact open bisections $U_{1},\dots,U_{n}$ containing $g_{1},\dots,g_{n}$, with equal source, and such that
+$f^{x}\coloneq\sum_{i=1}^{n}f(g_{i})\mathbbm{1}_{U_{i}}$
+is a finite sum of functions of the form $f^{\Theta}$ from (6.1).
+Let $U_{x}\coloneq s(U_{i})$ (note that this does not depend on $i$).
+In particular $x\in U_{x}$.
+We have $f|_{G_{x}}=f^{x}|_{G_{x}}$, so, by Lemma 5.1, $f-f^{x}$ is continuous at each $g\in G_{x}$.
+However, both $f$ and $f^{x}$ belong to $\mathbb{C}G$, and hence only take finitely many values.
+Therefore, by shrinking the $U_{i}$ if necessary, we can assume that $f|_{U_{x}}=f^{x}|_{U_{x}}$ (one can always shrink the compact open bisections in a tuple $\Theta$).
+The sets $\{U_{x}\}_{x\in s(K)}$ form an open cover for the compact set $s(K)$, so select a finite subcover $U_{x_{1}},\dotsc,U_{x_{m}}$.
+Removing intersections if necessary, we may assume that the $U_{x_{j}}$ are pairwise disjoint (this is possible since the $U_{x_{j}}$ are contained in the Hausdorff space $G^{(0)}$, and are thus clopen in $G^{(0)}$).
 Define
-$f^{\prime}\coloneq\sum_{j=1}^{n}(\mathbbm{1}_{W_{x_{j}}}\circ s)f^{x_{j}}.$
-Then $f^{\prime}\in J_{\mathbb{C}}$ satisfies ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$, as required. ∎
+$f^{\prime}\coloneq\sum_{j=1}^{m}f^{x_{j}}.$
+It is clear that $f^{\prime}$ is a finite sum of functions of the form $f^{\Theta}$ from (6.1). This completes the proof, since $f^{\prime}=f$ by construction. ∎
+Remark 6.2.
+The construction in this subsection generalises to ample groupoids that are not covered by countably many open bisections.
+After a modification analogous to that in Remark 5.6, one obtains, for any ample groupoid, and explicit spanning set for $J_{\mathbb{C}}$.
+$\bullet$
+6.2. Density of $J_{\mathbb{C}}$ in $J\cap\mathscr{C}_{c}(G)$
+We prove that the algebraic singular ideal $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ for any ample groupoid – no countability assumptions are needed.
+In Corollary 6.4 this observation is applied to the class of groupoids arising from contracting self-similar groups.
+Theorem 6.3.
+Let $G$ be any ample groupoid. Then $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to the norm $\left\|\cdot\right\|_{I}$.
+As stated earlier (see (5.4)), the norm $\left\|\cdot\right\|_{I}$ dominates the full norm $\left\|\cdot\right\|_{C^{*}(G)}$ and hence any $C^{*}$-norm.
+Therefore, $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
+Proof.
+Given $f\in J\cap\mathscr{C}_{c}(G)$, write $f=\sum_{j=1}^{m}f_{j}$ for some $f_{j}\in C_{c}(V_{j})$ and open bisections $V_{j}$, and let $H\subseteq G$ be the subgroupoid generated by the open bisections $V_{1},\dots,V_{m}$.
+Then $H$ is an open subgroupoid that is covered by countably many open bisections.
+By working in $H$ instead, we can assume that the groupoid $G$ is covered by countably many open bisections.
+Then, by Theorem 5.5, it suffices to assume that $f=f^{\Xi}$ for some tuple $\Xi=(x,h,b,U_{1},\dots,U_{n},\psi)$ as in Subsection 5.2, and where
+$f^{\Xi}=\sum_{i=1}^{n}b(g_{i})(\psi\circ s|_{U_{i}}),$
+as in (5.3).
+Take a compact set $K\subseteq G$ containing each $U_{i}$.
+As in the proof of Theorem 5.5 it suffices to find, for each $\varepsilon>0$, a function $f^{\prime}\in J_{\mathbb{C}}$ satisfying ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f^{\Xi}-f^{\prime}\right\|_{\infty}\leq\varepsilon$.
+Fix $\varepsilon>0$.
+Define $U\coloneq s(U_{i})$ (note that this does not depend on $i$).
+Since $U$ is totally disconnected, there exists a locally constant function $\varphi\in C_{c}(U)$ satisfying $\left\|\psi-\varphi\right\|_{\infty}\leq\frac{\varepsilon}{n\max_{i}(\left|b(g_{i})\right|)}$.
+Consider the function
+$f^{\prime}\coloneq\sum_{i=1}^{n}b(g_{i})(\varphi\circ s|_{U_{i}}).$
+It is clear that $f^{\prime}$ belongs to the Steinberg algebra $\mathbb{C}G$, and satisfies the conditions ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f^{\Xi}-f^{\prime}\right\|_{\infty}\leq\varepsilon$.
+Moreover, by Proposition 5.3, $f^{\prime}$ belongs to the singular ideal $J$, and hence also to $J_{\mathbb{C}}$.
+This completes the proof.
+∎
 We apply the previous theorem to the class of groupoids arising from contracting self-similar groups.
 Corollary 6.4.
-Let $G$ be the groupoid arising from the action of a contracting self-similar group on a finite alphabet. Then $J_{\mathbb{C}}$ is dense in $J$.
+Let $G$ be the groupoid arising from a contracting self-similar group action. Then $J_{\mathbb{C}}$ is dense in $J$.
 Proof.
-By Theorem 4.4 it suffices to check that $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
-By [23] and [15] the groupoid $G$ is second-countable and amenable.
-Moreover, [4, Corollary 7.13] implies that every element of $\tilde{G}^{(0)}$ is finite.
+By Theorem 6.3 it suffices to show that $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
+By [26, Corollary 2.19] (see also [29] and [15]) the groupoid $G$ is second-countable and amenable.
+Moreover, [5, Corollary 7.13] implies that every element of $\tilde{G}^{(0)}$ is finite.
 The result follows by Corollary 4.17 (II).
 ∎
+Remark 6.5.
+We would like to thank B.
+Kwaśniewski for pointing out to us that the previous result generalises to groupoids $G$ arising from contracting self-similar groupoid actions on finite graphs without sources.
+Such a groupoid is amenable by [26, Corollary 2.19], and it follows from either [23, Corollary 6.23] or [1, Proposition 9.1] that elements of $\tilde{G}^{(0)}$ are finite.